### v1

# references.bib

Abstract
Compositional generalization (the ability to respond correctly to novel combinations of familiar components) is thought to be a cornerstone of intelligent behavior. Compositionally structured (e.g. disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear. To address this gap, we present a general theory of compositional generalization in kernel models with fixed, potentially nonlinear representations (which also applies to neural networks in the “lazy regime”). We prove that these models are functionally limited to adding up values assigned to conjunctions/combinations of components that have been seen during training (“conjunction-wise additivity”), and identify novel compositionality failure modes that arise from the data and model structure, even for disentangled inputs. For models in the representation learning (or “rich”) regime, we show that networks can generalize on an important non-additive task (associative inference), and give a mechanistic explanation for why. Finally, we validate our theory empirically, showing that it captures the behavior of deep neural networks trained on a set of compositional tasks. In sum, our theory characterizes the principles giving rise to compositional generalization in kernel models and shows how representation learning can overcome their limitations. We further provide a formally grounded, novel generalization class for compositional tasks that highlights fundamental differences in the required learning mechanisms (conjunction-wise additivity).
1 Introduction
Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant. Stitching together concepts in this way allows humans to generalize far beyond our prior experience, preparing us to cope with unfamiliar situations and imagine things that do not yet exist [lake_building_2017, frankland_concepts_2020]. Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem [fodor_connectionism_1988, battaglia_relational_2018, lake_generalization_2018, hupkes_compositionality_2020, keysers_measuring_2020]. Accordingly, a broad range of studies have investigated the conditions under which compositionally structured (i.e. “disentangled”) representations can be learned [hinton_transforming_2011, bengio_representation_2013, higgins_beta-vae_2017, locatello_challenging_2019, esmaeili_structured_2019, trauble_disentangled_2021, whittington_disentanglement_2023]. However, it remains unclear whether learning these representations is actually useful: while some work suggests that disentangled representations improve compositional generalization [esmaeili_structured_2019, tokmakov_learning_2019, van_steenkiste_are_2019, whittington_constellation_2021], other studies challenge this view [locatello_challenging_2019, montero_role_2021, schott_visual_2022, xu_compositional_2022]. In particular, it is often unclear how insights on a certain compositional task generalize to others, as the relationship among them remains unclear [hupkes_compositionality_2020].
To clarify when compositionally structured representations yield compositional generalization, we define a family of compositional tasks, theoretically analyse the compositional generalization of kernel models, and then validate our theory in several relevant deep neural network architectures. Kernel models are an important class of statistical models in their own right (describing e.g. linear readout fine-tuning), provide a simplified approximation to neural network learning under certain conditions [jacot_neural_2018, chizat_lazy_2019], and more broadly have provided fundamental insight into generalization behavior in humans and machine learning models [jakel_generalization_2008, jakel_does_2009, canatar_spectral_2021, canatar_out--distribution_2021].
Despite the broad importance and relative simplicity of kernel models, it remains largely unclear how they generalize on compositional tasks [[, though see ]; see Section 2]abbe_generalization_2023,lippl_mathematical_2024. To address this gap, we present a theory of compositional generalization in kernel models. We define a new generalization class for compositional tasks and show that characterizing kernel models on such tasks helps us understand realistic neural network architectures. Our specific contributions are as follows:
  • We prove fundamental limits on the generalization behavior of kernel models with disentangled inputs, showing that they are constrained to summing up values implicitly assigned to each component or combination of components seen during training. We define the class of compositional tasks with this form as “conjunction-wise additive” (Section 4.1).
  • We define a metric to quantify how salient different component “conjunctions” (entangled features unique to component combinations) are in the representation (Section 4.2).
  • For conjunction-wise additive tasks, we then highlight two important failure modes impacting generalization (memorization leak and shortcut bias) (Section 4.3), which describe how biases in the training data can limit compositional generalization even for kernel models with disentangled inputs.
  • We validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks (Section 4.4).
  • Finally, we show how feature learning can overcome certain limitations of kernel models and learn non-additive tasks (Section 4.5).
2 Related work
Compositional generalization. Compositionality is an important theme across human and machine reasoning problems, including visual reasoning [lake_human-level_2015, johnson_clevr_2017, marino_ok-vqa_2019, schwartenbeck_generative_2023, zhou_compositional_2024], language production [hupkes_compositionality_2020], rule learning [ito_compositional_2022, abdool_continual_2023], image generation [okawa_compositional_2023], and robotics [zhou_policy_2022]. Recent breakthroughs (especially in language and computer vision) have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly [srivastava_beyond_2023, lewis_does_2023, west_advances_2023, ma_crepe_2023, chen_skills--context_2023]. Attempts to improve compositional generalization often leverage meta-learning [lake_human_2019, mitchell_challenges_2021, kumar_meta-learning_2021, wu_adaptive_2023, lake_human-like_2023], or modular architectures, in the hopes that different modules will specialize for different components [andreas_neural_2017]. End-to-end training of these architectures often does not result in the desired modular specialization [bahdanau_systematic_2019, mittal_is_2022, jarvis_specialization_2023]; notably, [schug_discovering_2024] theoretically clarify the necessary conditions for correct modular specialization and compositional generalization in hypernetworks.
Here we highlight that kernel models add up values assigned to each combination of features (or “conjunction”) seen during training. Notably, constraining a network to be additive can be sufficient for compositional generalization [lachapelle_additive_2023, wiedemer_compositional_2023]. Even without such constraints, language models encode many semantic concepts in an additive manner [mikolov_distributed_2013, naito_revisiting_2021]. Conjunctive codes (representations specific to a particular feature combination), on the other hand, have a long history in neuroscience [alvarado_properties_1992, baker_impact_2002], and are theoretically linked to forming highly specific, episodic memories.
Kernel and rich regime. Prior work has revealed two key strategies by which deep neural networks learn [chizat_lazy_2019, woodworth_kernel_2020]. In the kernel (or “lazy”) regime (which can be induced with large initial weights or wide networks), the networks’ learning is well approximated by gradient descent on a model with a fixed representation [jacot_neural_2018]. In the rich regime (brought forth, for example, by small initial weights, small width, low-rank connectivity, or loss functions like cross-entropy), the networks learn structured (i.e. abstract and sparse) representations over the course of learning [savarese_how_2019, woodworth_kernel_2020, chizat_implicit_2020, lyu_gradient_2020, saxe_neural_2022, liu_how_2023, domine_exact_2023]. This is often suggested to give rise to better generalization and more human-like behavior in neural networks [fort_deep_2020, vyas_limitations_2022, flesch_orthogonal_2022].
When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(x)=w^{T}r(x)$, this model depends on its representation $r(x)$ only through its induced kernel $K(x,x^{\prime})=\langle r(x),r(x^{\prime})\rangle$ [scholkopf_kernel_2000]. Specifically, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, it can be written in its “dual form” as
$f_{a}(x)=\sum_{i=1}^{n}a_{i}K(x,x_{i}).$ (1)
Gradient descent and ridge regression learn the readout weights that describe the training data with minimal $\ell_{2}$-norm; this is also true of neural networks in the kernel regime [soudry_implicit_2018, gunasekar_characterizing_2018, ji_gradient_2020]. Norm minimization is a standard theoretical framework for analyzing how representational geometry influences generalization [canatar_spectral_2021, canatar_spectral_2023] and we here apply it to the compositional task setting. This is similar to [lippl_mathematical_2024] who characterize model behavior on a specific compositional task (transitive ordering) and [abbe_generalization_2023] who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components. Compared to [abbe_generalization_2023, lippl_mathematical_2024], we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
[…]
Shortcut learning refers to models exploiting spurious correlations between certain features and the target to learn the training data and has been essential for understanding why the performance of deep neural networks can drop substantially when tested out-of-distribution [geirhos_shortcut_2020]. Models are more likely to use shortcut features that are easily extracted from the input and highly predictive of the output [hermann_what_2020, tartaglini_developmentally-inspired_2022, hermann_foundations_2023]. Shortcut learning has also been connected to the implicit bias of gradient descent [shah_pitfalls_2020, nagarajan_understanding_2021]. Our findings highlight a particular compositional shortcut: models tend to use a more salient, partially predictive component conjunction (memorizing the remaining data) over the fully predictive but less salient ground-truth conjunction.
3 Model and task setup
3.1 Task space
We assume that our model extracts a set of components $z=\sum_{c=1}^{C}z_{c}$ from its input $x$. Each $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components. For example, $z$ could specify the color and shape of an object, and the background color of an image (Fig. 1a). Alternately, $z$ may express different words in an input sentence (“the red sphere is in front of the blue background”). We assume that each component $z_{c}$ is represented in an orthogonal subspace (e.g. by concatenating representations of individual components, Fig. 1a). Further, because we consider categorical components, we can assume that for each component $c$, all representations $z_{c}\in Z_{c}$ have equal magnitude $m_{c}$ and are represented with equal similarity $\rho_{c}$ to alternative component values. The target, $y\in\mathbb{R}$, is some function of $z$ (see Section 3.3). After training models on certain component combinations $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations, $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$ to understand how biases in the training data affect compositional generalization.
3.2 Models
Linear readout models. We first consider kernel models which take in a disentangled representation $z$, apply a nonlinear transform $\phi(z)\in\mathbb{R}^{h}$ and learn a linear readout using gradient descent or ridge regression, $f_{w}(x):=w^{T}\phi(z)$ (Fig. 1a). For $\phi$, we consider a neural network with random weights:
Definition 3.1.
Given an input $z\in\mathbb{R}^{d}$, a network depth $L\geq 2$, and a set of widths $H_{1},\dotsc,H_{L}\in\mathbb{N}$, we define a neural network by recursively defining the operation $\phi^{(l)}$ of the $l^{\text{th}}$ layer as
$a^{(0)}:=z,\quad a^{(l)}:=\phi^{(l)}(a^{(l-1)}):=\sigma\left(\tfrac{1}{\sqrt{H_{l}}}\left(W^{(l)}a^{(l-1)}+b^{(l)}\right)\right),\quad W^{(l)}\in\mathbb{R}^{H_{l}\times H_{l-1}}$ (2)
where $W^{(l)}$ and $b^{(l)}$ are i.i.d. sampled from a random distribution and $\sigma:\mathbb{R}\to\mathbb{R}$ is a nonlinearity. The complete transform is then given by $\phi:=\phi^{(L)}\circ\dotsb\circ\phi^{(1)}$.
[…]
Deep neural network models. Additionally, we consider different deep neural network architectures used for vision (convolutional networks, ResNets, and Vision Transformers, Section 4.4) and ReLU networks trained through backpropagation (Section 4.5).
Figure 1: a, Theoretical setup: we consider a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$. b, Training sets for symbolic addition. The grid represents nine components with associated values $-4,-3,\dotsc,4$. Each tile represents a data point in the training set. c, Context dependence. In context 1, feat. 1 determines the category; in context 2, feat. 2 determines the category. The training sets leave out different subsets of the lower right orthant. d, Transitive equivalence: six items are split up into two arbitrary equivalence classes (e.g. A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes.
3.3 Example tasks
In addition to deriving a general theory, we consider a set of example tasks that are important building blocks for compositional reasoning in machine learning and cognitive science (Appendix D discusses additional tasks including logical operations and partial exposure):
Symbolic addition. Many tasks involve inferring a magnitude associated with different underlying components (e.g. handwritten digits) [lorenzi_numerosities_2021, sheahan_neural_2021]. We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(I_{1})$ and $v_{2}(I_{2})$. The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$. After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items [wiedemer_compositional_2023]. In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values. This can require interpolation, extrapolation, or both (Fig. 1b).
Context dependence. The relevance of different stimuli often depends on the context we are in. Taking into account this context is a crucial aspect of cognition in humans and animals [bouton_stimulus_1999, taylor_context-dependent_2013, parker_cognitive_2014, ito_compositional_2022]. We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$. The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response. Both features have six possible values which are split up in two categories. If the model has learned this context dependence, it should be able to generalize to novel feature combinations. We evaluate on the orthant for which $z_{f1}$ indicates Cat. 2 and $z_{f2}$ indicates Cat. 1. In the most extreme generalization test (CD-3), we leave out the entire orthant, in easier versions we leave out conjunctions of two or one of those features (CD-2 and CD-1) (Fig. 1c).
Transitive equivalence. Relational reasoning is an important instance of compositional generalization and often involves extending the learned relations to new item combinations [halford_relational_2010, battaglia_relational_2018]. Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent. The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig. 1d). This is an important instance of relational cognition (often studied as “associative inference” in cognitive science [schlichting_memory_2015, spalding_ventromedial_2018]). Note that prior work has found that kernel models often successfully generalize a transitive ordering relation [lippl_mathematical_2024].
4 Results
4.1 Kernel models with compositional structure are conjunction-wise additive
Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs: specifically, they assign a value to each combination of components seen during training and generalize to test inputs by adding up the relevant combinations’ values. We call this motif “conjunction-wise additivity.” Below we formally state our finding and then explain its implications.
We first note that the kernel of the disentangled representation, $K(z,z^{\prime})=z^{T}z^{\prime}$, only depends on the components for which the two inputs are overlapping, $O(z,z^{\prime}):=\left\{c|z_{c}=z_{c}^{\prime}\right\}$ where $c\in\{1,\dotsc,C\}$. This is necessitated by the fact that separate components are represented in orthogonal subspaces. We call any representation satisfying this criterion “compositionally structured.” Notably, the hidden representation $\phi(z)$ is also compositionally structured (in the infinite-width limit). This is because the kernel of the neural network’s representation, $K_{\phi}(z,z^{\prime}):=\phi(z)^{T}\phi(z)$, only depends on the inputs $z,z^{\prime}$ through their kernel $K(z,z^{\prime})$ and thus conserves this property [[]; see Appendix B]cho_kernel_2009,tsuchida_invariance_2018,tsuchida_richer_2019,han_fast_2022.
Intriguingly, we find that any kernel model with a compositionally structured representation is constrained to be conjunction-wise additive. To state this finding, we define, for each $z\in Z$, the set of conjunctions for which $z$ overlaps with some element in the training set $z^{\text{tr}}\in Z^{\text{train}}$,
$O(z|Z^{\text{train}}):=\left\{S\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}\text{ s.t. }S\subseteq O(z,z^{\text{tr}})\right\}.$ (3)
We then prove:
Theorem 4.1.
For $z\in Z$, any kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$ can be expressed as a sum of functions $f_{S}$ over all conjunctions $S\in O(z|Z^{\text{train}})$, where each $f_{S}$ associates a value with each unique combination of features $z_{c}$ for $c\in S$:
$f(z)=\sum_{S\in\mathcal{O}(z|Z^{\text{train}})}f_{S}(z_{S}),\quad z_{S}:=(z_{c})_{c\in S}.$ (4)
We prove the theorem in Appendix A. Intuitively, it holds because the model associates a distinct weight vector with each conjunction/combination of components. The weight vectors associated with conjunctions not seen during training remain at their initial value and can therefore not be used.
To understand the theorem’s implications, we first consider a direct linear readout, $f(z)=w^{T}z$. This constrains the model to adding up a value for each component (“component-wise additivity”):
$f(z)=w^{T}z=w^{T}\sum_{c=1}^{C}z_{c}=\sum_{c=1}^{C}f_{c}(z_{c}),\quad f_{c}(z_{c}):=w^{T}z_{c}.$ (5)
Component-wise additivity is a strict subset of conjunction-wise additivity (each $f_{S}$ depending on more than one component is set to zero). It lets the model generalize perfectly on component-wise additive tasks like symbolic addition. However, it also prevents the model from learning any training set that is not component-wise additive, including context dependence and transitive equivalence.
The nonlinear transform $\phi$ can overcome this constraint [hornik_multilayer_1989, cybenko_approximation_1989, leshno_multilayer_1993, rigotti_importance_2013]. Indeed, conjunction-wise additive functions more broadly can learn arbitrary training data, as the full conjunction $S=\{1,\dotsc,C\}$ is in $O(z|Z^{\text{train}})$ for each $z\in Z^{\text{train}}$ and $f_{S}(z)$ can take on a distinct value for each $z$. However, for $z\in Z^{\text{test}}$ the full conjunction is not in $O(z|Z^{\text{train}})$. Thus, while conjunction-wise additivity (unlike component-wise additivity) does not impose a constraint on the kinds of training data the model can learn, it does impose constraints on the kinds of generalizations it can implement.
In particular, for inputs with two components, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ for $z\in Z^{\text{train}}$. However, for $z\in Z^{\text{test}}$, $f_{12}(z)$ falls away and kernel models are constrained to an additive computation: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$. Thus, they cannot generalize on any non-additive task, including transitive equivalence: while they learn the training data (using $f_{12}(z)$), they cannot generalize to the test data.
For inputs with more than two components, a conjunction-wise additive computation does not just encode each component and their full conjunction, but also partial conjunctions between components (e.g. 12, 13, and 23 for three components). For test inputs, the model can still not rely on the full conjunction (123), but it can rely on any partial conjunction it may have seen before.
To see whether such a model can, in principle, generalize on a given task, we must 1) identify the overlaps with the training set, $\mathcal{O}(z|Z^{\text{train}})$, and 2) determine whether the target can be written as a conjunction-wise sum. For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions (i.e. 12 and 13), but not the feature-feature conjunction (i.e. 23). As a result, model behavior on the test set can be written as $f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2})$. While the task is not component-wise additive, we can use $f_{12}$ and $f_{13}$ to encode the task: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the first feature $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
Hence, conjunction-wise additivity not only tells us whether kernel models with compositionally structured representations can solve a certain task, but also how they solve it. Further, it tells us to make a task non-additive: in Section D.2.3, we describe a modification that makes context dependence non-additive by requiring generalizations to novel context-feature conjunctions.
Importantly, conjunction-wise additivity only determines whether certain compositional generalizations are, in principle, feasible. Model generalization also depends on whether the model identifies the correct conjunction-wise function. We discuss this in the next two sections.
4.2 Overlap salience characterizes compositional representational geometry
Model generalization is determined by how prominently different component conjunctions are represented. This is measured by the similarity $K(S)$ between inputs with overlap $S\subseteq\{1,\dotsc,C\}$. $K(S)$ depends on responses to that particular conjunction $S$ as well as all of its sub-conjunctions $S^{\prime}\subsetneq S$. We thus define the scalar salience of an overlap $S$ as its unique contribution. We then normalize this salience, so all saliences add up to one:
Definition 4.2.
For a representation with kernel $K$ and a conjunction $S\subseteq\{1,\dotsc,C\}$, we define
$\overline{\text{Sal}}(\emptyset):=K(\emptyset),\quad\overline{\text{Sal}}(S):=K(S)-\sum_{S^{\prime}\subsetneq S}\overline{\text{Sal}}(S^{\prime}),\quad{\text{Sal}}(S)=\tfrac{\overline{\text{Sal}}(S)}{\sum_{\emptyset\neq S\subseteq\{1,\dotsc,C\}}\overline{\text{Sal}}(S)}.$ (6)
Figure 2: Normalized salience (i.e. $\text{Sal}_{C}(k)$) for neural networks with random weights and different nonlinearities for inputs with three components.
We now analyze how the representational salience evolves over different layers of a random network, considering an uncorrelated disentangled correlation with equal magnitudes (i.e. $\rho_{c}=0$ and $m_{c}=m_{c^{\prime}}$) (see Section B.2 for analysis beyond these assumptions). In that case, the salience only depends on the overlap size $k:=|S|$ and we denote it by $\text{Sal}_{C}(k)$.
Notably, conjunctions are only encoded and can only be used if they have non-zero salience. In particular, the disentangled representation has zero salience for all conjunctions of more than one component. As the network gets deeper, the salience of these conjunctions increases. In fact, for (leaky) ReLU networks, the salience of the full conjunction converges to one:
Proposition 4.3.
For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $\text{Sal}_{C}(k)\to 0$ for $k<C$ and $\text{Sal}_{C}(C)\to 1$.
We prove this statement in Section B.1. Empirically, we find that $\text{Sal}_{C}(C)$ indeed increases with depth, whereas $\text{Sal}_{C}(1)$ decreases (Fig. 2). The salience of intermediate conjunctions (e.g. $\text{Sal}_{3}(2)$) first increases and then decreases, its trajectory depending on the nonlinearity.
4.3 Kernel models suffer from two failure modes: memorization leak and shortcut bias
In Section 4.1, we found that kernel models sum up values assigned to different component conjunctions. This leaves ambiguous, however, whether they generalize correctly, as they can learn the training data using different combinations of conjunctions. Notably, kernel models trained with gradient descent or ridge regression generally learn the readout weights with minimal $\ell_{2}$-norm [[, Section 2;]]soudry_implicit_2018,gunasekar_characterizing_2018,ji_gradient_2020. This inductive bias gives rise to weights that are distributed across different predictive features. For in-distribution generalization, this is often useful, as it allows us to integrate different kinds of evidence. However, for compositional generalization, we often need to rely on a small group of conjunctions and the tendency towards distributed weights causes generalization failures.
Here we highlight two particular failure modes arising from this. First, the full conjunctions always serve as predictive features (“memorization leak”). Second, the model can often use statistical shortcuts to predict parts of the training set, using the full conjunction to learn the rest (“shortcut bias”). Below we explain how these failure modes impact symbolic addition and context dependence.
Symbolic addition suffers from a memorization leak. The functional form $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ implies that using the full conjunction $f_{12}(z)$ to learn the training set will necessarily distort the inferred values $f_{1}(z_{1}),f_{2}(z_{2})$. This, in turn, negatively impacts generalization on the test set. To analyze the impact of such a memorization leak in detail, we characterize model behavior on symbolic addition analytically for training sets that are balanced around zero:
Proposition 4.4.
Consider input elements in $\{[v]\}_{v\in\mathcal{V}},\mathcal{V}\subset\mathbb{R}$ with associated values $v$. We assume that the training set contains all pairs such that at least one component is $z_{c}\in\{[w]\}_{w\in\mathcal{W}},\mathcal{W}\subset\mathbb{R}$ and that the average value in both $\mathcal{V}$ and $\mathcal{W}$ is zero. Then, model behavior on the test set is given by
$f([i],[j])=m(i+j),\quad m:=\tfrac{p\cdot\text{Sal}_{2}(1)}{1+(p-2)\text{Sal}_{2}(1)},\quad p:=|\mathcal{W}|$ (7)
We prove the proposition in Section D.1.1. It implies that for any representation that encodes the conjunction (i.e. has $\text{Sal}_{2}(2)>0$ and therefore $\text{Sal}_{2}(1)<\tfrac{1}{2}$), the model underestimates the inferred values by a constant factor $m<1$. Further, $m$ decreases for smaller $\text{Sal}_{2}(1)$ and increases for larger training set size (Fig. 3a,b). Perhaps surprisingly, these are the only two factors influencing $m$. In particular, while we often assume that interpolation is easier than extrapolation, these changes in training set do not impact model behavior here. We found qualitatively similar behavior for randomly generated, dispersed training sets (Section D.1.2, Fig. 7).
Figure 3: Kernel models’ behavior on (a,b) symbolic addition and (c,d) context dependence. a, Factor $m$ distorting the inferred values as a function of $\text{Sal}_{2}(1)$ and the training set size $p=|\mathcal{W}|$. b, Inferred values on extrapolation task. c, Variants of context dependence for which different representational saliences yield successful generalization. Trajectories of networks with different nonlinearities are highlighted (color scale see Fig. 2). d, Inferred coefficients of the different conjunction types for two example representations: a network with three layers and a rectified linear or quadratic nonlinearity.
Notably, because the full conjunction can always be used to learn the training set, the memorization leak impacts a broad range of compositional tasks, including partial exposure (Section D.4).
Context dependence suffers from a shortcut bias. Next, we empirically analyzed model generalization on context dependence across different representational geometries and training sets. We found that on a given task, each model either generalizes with 100% or 0% accuracy. For CD-3, the model only generalizes when $\text{Sal}_{3}(2)$ is high relative to $\text{Sal}_{3}(1)$ (Fig. 3c). As a result, whether the network generalizes is highly sensitive to the nonlinearity and depth of the network. In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
To better understand this, we analyzed the total magnitude of model weights associated with the different conjunctions. Notably, there are many possible conjunctions the model could rely on (see Section D.2.4). In particular, context $1$ is more likely to have the target $y=1$ and context $2$ is more likely to have target $y=-1$. The model may thus rely on this statistical shortcut, using the full conjunction to memorize the remaining training data. Indeed, we found that representations that generalized unsuccessfully on CD-3 consistently had a high weight associated with the context component and the full conjunction (Fig. 3d). Notably, for CD-3, the context shortcut yields an accuracy of $\tfrac{2}{3}$. This explains why when $\text{Sal}_{3}(1)$ and $\text{Sal}_{3}(3)$ are high relative to $\text{Sal}_{3}(2)$, the model uses the context-driven shortcut together with the full conjunction, resulting in a failure to generalize. For CD-2, the context-driven shortcut only results in an accuracy of $\tfrac{9}{16}$, making it much less useful. This explains why only very low $\text{Sal}_{3}(2)$ yields failure to generalize on CD-2 or CD-1.
Our analysis illustrates how conjunction-wise additivity lets us identify component conjunctions that yield a shortcut. The representational salience of these conjunctions and the extent to which they can explain the training set then helps us understand when this shortcut impacts generalization.
4.4 Conjunction-wise additivity can describe generalization behavior in deep neural networks
So far, we have analyzed kernel models, for their analytical tractability. We conjecture that large-scale neural network architectures with more complex, interrelated inputs might also be prone to implementing a conjunction-wise additive computation. In that case, our analyis of the kernel models would shed light on when the statistics of the training data enable or prevent compositional generalization in deep neural networks as well. To test our conjecture, we trained convolutional networks (ConvNets, [lecun_backpropagation_1989]) on a version of the tasks with concatenated MNIST digits [lecun_gradient-based_1998] and residual networks (ResNets, [he_deep_2016]) and Vision Transformers (ViTs, [dosovitskiy_image_2020]) on a version of the tasks with concatenated CIFAR-10 images [krizhevsky_learning_2009]. We considered the same tasks except that each component was now given by images of a particular category, rather than being a single instance. Notably, different image categories are not necessarily equally correlated with each other. To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment. To compare the network behaviors to our theoretical predictions, we fit a conjunction-wise additive function to the network output using linear regression (“additivity analysis”; see Section C.3).
Symbolic addition. We trained the ConvNets on a total of 20,000 randomly generated MNIST samples for 100 epochs and trained the ResNets and ViTs on a total of 40,000 CIFAR-10 samples for 100 and 200 epochs, respectively. Remarkably, we found that the kernel theory matches network behavior almost perfectly: the networks’ average predictions (across the different possible permutations of the digit categories) were well predicted by an additive structure and their inferred values generally underestimated the ground truth (Fig. 4a, Figs. 9 and 10). This suggests that the kernel theory can capture the behavior of relevant neural network architectures trained on natural image inputs.
Next, to investigate the effect of increased conjunctivity, we varied the distance between the two MNIST digits. We hypothesized that the ConvNets’ local weight structure should produce a more conjunctive representation for digits that are closer together. To test this, we determined $\text{Sal}_{2}(1)$ in an intermediate layer of the network, averaging over different instances of all digits. We found that $\text{Sal}_{2}(1)$ was indeed smaller for lower distances (Fig. 4b). Further, the inferred values were more compressed for smaller distances, confirming that more conjunctive inputs exacerbate the memorization leak in ConvNets as well (Fig. 4c and Fig. 9).
Figure 4: Behavior of networks trained on MNIST and CIFAR versions of compositional tasks. Shaded regions indicate mean $\pm$ two standard errors. a, Inferred values on symbolic addition (extrapolation, distance of zero between digits). b, $\text{Sal}_{2}(1)$ in an intermediate layer of the ConvNet, for different distances between digits. c, Inferred values for the extrapolation task and different distances between digits. d, For the networks trained on variants of context dependence, the inferred coefficients for different conjunction types (distance of 20px).
Context dependence. We then trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples. Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but failing to generalize on CD-3. The additivity analysis revealed that this was due to a context-driven shortcut (Fig. 4d).
4.5 Feature learning can overcome the limitations of kernel models
Figure 5: Feature learning enables generalization on transitive equivalence. a, Learning curves in the lazy and rich regime. b, The weights of the network in the subspace corresponding to one underlying XOR-task. c, Weights for the same unit are plotted against each other and colored by whether they correspond to equivalent items (purple) or non-equivalent items (green). d, Accuracy of a ConvNet trained on an MNIST version of transitive equivalence.
We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence. To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input. By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime. Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs. 8 and 11). On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig. 5a).
To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm ([lyu_gradient_2020, chizat_implicit_2020]; cf. [vardi_implicit_2021]). In particular, a one-hidden-layer ReLU network tends to learn a sparse set of features [savarese_how_2019, chizat_implicit_2020]. Transitive equivalence consists of multiple overlapping equality relations (e.g. $A=B\neq D=E$). Notably, ReLU networks such an XOR-type problem by specializing one unit to each conjunction ([brutzkus_why_2019, saxe_neural_2022, xu_benign_2023]; Fig. 5b). Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g. $(A,B),(A,C)$, and $(B,C)$). This causes the unit to generalize to unseen item combinations (e.g. $(A,C)$), enabling the network to generalize. Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig. 5c).
Thus, rich networks’ capacity for abstraction gives rise to an additional compositional motif, allowing them to generalize on transitive equivalence. In particular, our findings highlight that transitive equivalence and transitive ordering are solved by fundamentally different network motifs.
To see whether large-scale neural networks can also benefit from this feature-learning mechanism, we trained ConvNets on an MNIST version of transitive equivalence. The networks were trained for 150 epochs on 20,000 samples. We found that if the digits were presented with a distance of zero, the network did not generalize compositionally at all. However, with increasing distance, the network started to improve its compositional generalization (Fig. 5e), demonstrating that a convolutional network can benefit from this rich compositional motif.
5 Discussion
Humans often generalize to new situations by stitching together concepts and knowledge from prior experience. Despite the broad importance of this ability (both for cognitive science and machine learning), it has remained unclear under what conditions compositionally structured representations give rise to compositional generalization. Here we have taken a step toward formalizing this relationship by clarifying the full range of compositional motifs implemented by kernel models with categorical compositionally structured representations (“conjunction-wise additivity”). For conjunction-wise additive tasks, we analyzed how representational geometry and training data impact successful generalization on important compositional building blocks, validating our analysis for deep neural networks trained on natural image data. For non-additive tasks, our results immediately imply that kernel models are unable to generalize. We then analyzed how feature-learning mechanisms can overcome these limitations and demonstrated that deep networks benefit from the same mechanisms. Taken together, this suggests that conjunction-wise additivity provides a useful generalization class for understanding compositional generalization in neural networks and perhaps even humans.
Limitations. We considered tasks with a fixed number of input components and fixed output dimensionality and assumed that all input components are represented orthogonally, excluding continuously varying components such as position. Further, our theoretical analysis focused on kernel models with fixed representations (though we considered deep neural networks through empirical simulations). Future work could extend the presented theory beyond these assumptions, and could further explore the range of compositional motifs in other learning models (e.g. rich neural networks and modular networks) more systematically. Nevertheless, we believe that this work is an important step towards a comprehensive theory of compositional generalization, covering a broad range of tasks and highlighting several phenomena that might help us understand more complex models as well.
Acknowledgments
[…]
Appendix A Proof of Theorem 4.1
See 4.1
Proof.
Because $K$ is compositionally structured, its similarity $K(I,I^{\prime})$ only depends on the overlap $O(I,I^{\prime})\subseteq\{1,\dotsc,C\}$. We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define the overlap in the training dataset as
$\mathcal{D}_{S}(I):=\left\{I^{\prime}\in\mathcal{D}^{(\text{train})}|\forall_{c\in S}I_{c}=I^{\prime}_{c}\right\}.$ (8)
The key idea is to decompose $\mathcal{D}^{(\text{train})}$ into these different overlaps in order to separate the sum into its components. However, by our definition, the datasets $\mathcal{D}_{S}(I)$ are not disjoint. Indeed, $S\subseteq S^{\prime}$ implies $\mathcal{D}_{S^{\prime}}(I)\subseteq\mathcal{D}_{S}(I)$ and in particular $\mathcal{D}_{\emptyset}(I)=\mathcal{D}^{(\text{train})}$. To adjust for this, we define $\delta_{S}$ as the similarity added by $\kappa_{S}$ to the similarity between conjunctions with one component fewer, recursively defining
$\delta_{\emptyset}=\kappa_{\emptyset},\quad\delta_{S}=\kappa_{S}-\sum_{S^{\prime}\subsetneq S}\delta_{S^{\prime}}.$ (9)
We then decompose
$f(I)=\sum_{I^{\prime}\in\mathcal{D}^{(\text{train})}}a_{I^{\prime}}K(I,I^{\prime})=\sum_{S\subseteq\{1,\dotsc,C\}}\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}}.$ (10)
This equality obtains because for each $I^{\prime}\in\mathcal{D}^{(\text{train})}$,
$\sum_{S:I^{\prime}\in\mathcal{D}_{S}(I)}\delta_{S}=\delta_{O(I,I^{\prime})}+\sum_{S\subsetneq O(I,I^{\prime})}\delta_{S^{\prime}}=\kappa_{O(I,I^{\prime})}=K(I,I^{\prime}),$ (11)
which is true by definition. We note that for $S\notin\mathcal{O}(I)$, $\mathcal{D}_{S}(I)=\emptyset$. Defining
$f_{S}(I):=\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}},$ (12)
proves the proposition. ∎
Appendix B Representational geometries across different deep neural network architectures
We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022].
B.1 Proof of Proposition 4.3
See 4.3
Proof.
Note that the proof is a minor extension of Lemma S1.3 in [lippl_mathematical_2024]. We present it here in a self-contained manner. We consider a nonlinearity
$\sigma(u):=A\min(u,0)+\max(u,0),\quad A\in[0,1).$ (13)
By prior work, [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022],
$\mathbb{E}_{w}\left[\phi({w}^{T}{h}^{(l)}(z))\phi({w}^{T}{h}^{(l)}(z^{\prime}))\right]=\sigma^{2}\|h^{(l)}(z)\|_{2}\|h^{(l)}(z^{\prime})\|_{2}k\left(\hat{h}^{(l)}(z)^{T}\hat{h}^{(l)}(z^{\prime})\right),$ (14)
where
$k(u)=\frac{(1-A)^{2}}{2\pi}\left(\sqrt{1-u^{2}}+(\pi-\cos^{-1}(u))u\right)+Au,$ (15)
$\sigma^{2}$ is the variance of the sampled weights, and $\hat{h}=h/\|\hat{h}\|_{2}$.
This means that for any two inputs that have a certain similarity $u$, their similarity in the $L$-th layer is given by $k^{(L)}(u)$, where $k^{(L)}$ denotes the $L$-times application of $k$. Let distinct trials in the input have a similarity of $\kappa_{d}$ and let identical trials have a similarity of $\kappa_{i}$. Any set of trials with overlapping components will have a similarity $\kappa$, $\kappa_{d}<\kappa<\kappa_{i}$. We denote their corresponding similarity in the $L$-th layer by $\kappa_{i}^{(L)},\kappa_{d}^{(L)},\kappa^{(L)}$. Our goal is now to show that
$\lim_{L\to\infty}s^{(L)}=0,\quad s^{(L)}:=\frac{\kappa^{(L)}-\kappa^{(L)}_{d}}{\kappa^{(L)}_{i}-\kappa^{(L)}_{d}}=0.$ (16)
This implies directly that the salience of all partial conjunctions converges to zero, which in turn implies that the salience of the full conjunction converges to one.
(14) implies that $\kappa_{i}^{(L+1)}=\sigma^{2}\kappa_{i}^{(L)}k(1)=\sigma^{2}\kappa_{i}^{(L)}\tfrac{1+A^{2}}{2}$. Notably, all inputs have the same magnitude and therefore have the same magnitude through all layers; this is given by $\sqrt{\kappa_{i}^{(L)}}$. We can therefore denote
$\kappa^{(L+1)}=\sigma^{2}\kappa_{i}^{(L)}k\left(\kappa^{(L)}/\kappa_{i}^{(L)}\right).$ (17)
Thus,
$s^{(L+1)}=\frac{k(\kappa^{(L)}/\kappa_{i}^{(L)})-k(\kappa_{d}^{(L)}/\kappa_{i}^{(L)})}{k(1)-k(\kappa_{d}^{(L)}/\kappa_{i}^{(L)})}$ (18)
We thus define new normalized variables $\hat{\kappa}^{(L)}:=\kappa^{(L)}/\kappa_{i}^{(L)},\hat{\kappa}_{d}^{(L)}:=\kappa_{d}^{(L)}/\kappa_{i}^{(L)}$, i.e.
$s^{(L)}=\frac{\hat{\kappa}^{(L)}-\hat{\kappa}^{(L)}_{d}}{1-\hat{\kappa}_{d}^{(L)}},$ (19)
and therefore
$\hat{\kappa}^{(L)}=(1-\hat{\kappa}_{d}^{(L)})s^{(L)}+\hat{\kappa}_{d}^{(L)}.$ (20)
Note that $\hat{\kappa}^{(L+1)}=k(\hat{\kappa}^{(L)})/k(1)$ and $\hat{\kappa}_{d}^{(L+1)}=k^{(}\hat{\kappa}^{(L)}_{d})/k(1)$. We thus define
$\tilde{k}(u):=\frac{k(u)}{k(1)}=u+\frac{\rho}{\pi}(\sqrt{1-u^{2}}-\cos^{-1}(u)u),\quad\rho:=\frac{(1-A)^{2}}{(1+A)^{2}}.$ (21)
Note that
$\tilde{k}^{\prime}(u)=1+\frac{\rho}{\pi}\left(-\frac{u}{\sqrt{1-u^{2}}}-\cos^{-1}(u)+\frac{u}{\sqrt{1-u^{2}}}\right)=1-\frac{\rho}{\pi}\cos^{-1}(u).$ (22)
Note that $k(1)=1$ and as for all $10\leq u<$ 1, $\tilde{k}^{\prime}(u)<1$, this is the only fixed point and $\hat{\kappa}^{(L)},\hat{\kappa}_{d}^{(L)}\to\infty$. We can therefore define
$s^{(L+1)}=\frac{\tilde{k}\left((1-\hat{\kappa}_{d}^{(L)})s^{(L)}+\hat{\kappa}_{d}^{(L)}\right)-\hat{\kappa}_{d}^{(L)}}{1-\kappa_{d}^{(L)}}.$ (23)
We now determine the fixed mapping to this mapping assuming that $\hat{\kappa}_{d}^{(L)}$ is fixed at some value $d$, i.e.:
$\begin{split}f(s,d)&=\frac{\tilde{k}((1-d)s+d)-\tilde{k}(d)}{1-\tilde{k}(d)}.\end{split}$ (24)
$s=0$ is a fixed point. Further,
$\frac{\partial f(s,d)}{\partial s}=\frac{(1-d)\tilde{k}^{\prime}((1-d)s+d)}{1-\tilde{k}(d)}=\frac{(1-d)(1-\tfrac{\rho}{\pi}\cos^{-1}((1-d)s+d)}{1-d-\tfrac{\rho}{\pi}(\sqrt{1-d^{2}}-\cos^{-1}(d)d)}.$ (25)
We now prove that this for $0<s\leq\tfrac{1}{2}$, this derivative is smaller than 1. Specifically,
$\begin{split}&(1-d)(1-\tfrac{\rho}{\pi}\cos^{-1}((1-d)s+d)=\\ &1-d-\tfrac{\rho}{\pi}(\sqrt{1-d^{2}}-\cos^{-1}(d)d)+\tfrac{\rho}{\pi}r(s,d),\end{split}$ (26)
where the residual is given by
$\begin{split}r(s,d):=(d-1)\cos^{-1}((1-d)s+d)+\sqrt{1-d^{2}}-d\cos^{-1}d.\end{split}$ (27)
We now need to prove that $r(s,d)<0$. Note that $r(s,d)$ is monotonically increasing in $s$ and therefore
$\begin{split}r(s,d)\leq r(\tfrac{1}{2},d)&=(d-1)\cos^{-1}(\tfrac{1}{2}+\tfrac{1}{2}d)+\sqrt{1-d^{2}}-d\cos^{-1}d<0,\end{split}$ (28)
where we infer the latter inequality by visual inspection of the plot of this function. ∎
B.2 Extended analysis of representational salience
Figure 6: Extended analysis of overlap salience in random neural networks. a, Salience of the first component (1), the second component (2), and their conjunction (12), where we vary the two components’ saliences in the input. b, Salience for inputs with four components.
To complement Section 4.2, we analyze the impact of different magnitudes for different components. In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig. 6a). As becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation. While the salience of the full conjunction still eventually converges to one for most nonlinearities, it takes longer to do so for a less balanced input representation. Further, for a Gaussian nonlinearity, an imbalanced representation actually decreases the limit salience the representation appears to be converging to for the full conjunction.
In Fig. 6, we plot the different saliences for an input with four components. Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig. 2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions. Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two.
Appendix C Detailed methods
C.1 Models
Kernel model.
We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn [pedregosa_scikit-learn_2011].
Rich and lazy ReLU networks.
All networks were trained with Pytorch and Pytorch Lightning [paszke_pytorch_2019]. We consider ReLU networks with one hidden layer and $H=1000$ units. We initialize by $\sigma\sqrt{2/H}$, considering $\sigma\in[10^{-6},1]$. In particular, when reporting results on rich networks (without further specification), we assume $\sigma=10^{-6}$. When reporting results on lazy network, we assume $\sigma=1$.
Convolutional neural networks.
We considered networks with four convolutional layers (kernel size is five, two layers have 32 filters, two have 64 filters) and two densely connected layers (with 512 and 1024 units). Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation. All weights are initialized with He initialization [he_delving_2015].
Residual neural networks.
We trained a residual neural network with eight blocks in total, two with 16, 32, 64, and 128 channels, respectively, using the Adam optimizer with a learning rate of $10^{-3}$ for 100 epochs.
Vision Transformers.
Finally, we trained a Vision Transformer (ViT) with six attention heads, 256 dimensions for both the attention layer and the MLP, and a depth of four, using Adam with a learning rate of $10^{-4}$ for 200 epochs.
Data augmentation.
We did not use data augmentation for MNIST. For CIFAR-10, we used a random flip and a random crop.
C.2 Reproducibility
Computational simulations.
We ran all computational simulations either on a CPU or on a single GPU. Each kernel experiment (total number: 53 experiments) ran within fifteen minutes, each rich network (total number: 450 experiments) and ConvNet experiment (total number: 600 experiments) took between one and two hours, and each ViT or ResNet experiment (total number: 140 experiments) took between five and ten hours. We used a computational cluster to run many of the experiments in parallel. As a rough estimate, the total experiments therefore required 900 hours of CPU computations and 2,600 hours of GPU computations.
Code repository.
The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization. We additionally uploaded the repository including all simulated data to Zenodo (https://doi.org/10.5281/zenodo.11308156).
C.3 Additivity analysis
To analyze how well a conjunction-wise additive computation can describe network behavior, we considered as the set of possible features a concatenation of one-hot vectors coding for each possible conjunction. We then removed all features that are constant at zero on the training dataset and used linear regression to try and predict network behavior on both training and test set for all remaining features. The resulting $R^{2}$ defines the “additivity” of the network behavior (i.e. $R^{2}=1$ indicates full conjunction-wise additivity). Furthermore, we can use the inferred values assigned to these different conjunctions to compare kernel models, rich and lazy networks, and convolutional networks. Note that for the convolutional networks, we first average the model predictions across all different images instantiating a given compositional input.
Appendix D Compositional tasks
D.1 Symbolic addition
D.1.1 Proof of Proposition 4.4
See 4.4
Proof.
We split up the training data into
$\mathcal{I}^{(1)}:=\{[w]|w\in\mathcal{W}\}^{2},$ (29)
and
$\mathcal{I}^{(2)}$ $:=\bigcup_{w\in\mathcal{W}}\mathcal{I}^{(1,w)}\cup\mathcal{I}^{(2,w)},$ (30)
$\mathcal{I}^{(1,w)}$ :=$:=\left\{([w],[v])|v\in\mathcal{V}\setminus\mathcal{W}\right\},\quad\mathcal{I}^{(2,w)}:=\left\{([v],[w])|v\in\mathcal{V}\setminus\mathcal{W}\right\}.$ (31)
We denote the dual coefficient associated with each training point $([i],[j])$ by $a_{ij}\in\mathbb{R}$. Note that the problem is symmetric and therefore we know that $a_{ij}=a_{ji}$. We define a few summed coefficients:
$b_{v}:=\sum_{w\in\mathcal{W}}a_{vw},\quad\overline{b}_{w}:=\sum_{v\in\mathcal{V}\setminus\mathcal{W}}a_{vw},\quad c_{w}:=\sum_{w^{\prime}\in\mathcal{W}}a_{ww^{\prime}},$ (32)
$b:=\sum_{v\in\mathcal{V}\setminus\mathcal{W}}b_{v}=\sum_{w\in\mathcal{W}}b_{w},\quad c:=\sum_{w\in\mathcal{W}}c_{w}.$ (33)
Note that the sum over all dual coefficients is given by $2b+c$. Let $p:=|\mathcal{W}|$ and $q:=|\mathcal{V}|-|\mathcal{W}|$. Then, setting $\delta_{2}:=\kappa_{2}-\kappa_{0}$ and $\delta_{1}:=\kappa_{1}-\kappa_{0}$, the set of dual equations is given by
$([w],[w^{\prime}])\in\mathcal{I}^{(1)}:\kappa_{0}(2b+c)+\delta_{1}(\overline{b}_{w}+\overline{b}_{w^{\prime}}+c_{w}+c_{w^{\prime}})+(\delta_{2}-2\delta_{1})a_{ww^{\prime}}$ $=w+w^{\prime},$ (34)
$([w],[v])\in\mathcal{I}^{(1,w)}:\kappa_{0}(2b+c)+\delta_{1}(b_{v}+\overline{b}_{w}+c_{w})+(\delta_{2}-2\delta_{1})a_{wv}$ $=w+v.$ (35)
(Note that the equation corresponding to $([v],[w])$ is equivalent due to the problem’s symmetry.)
The prediction is given by
$f([v_{1}],[v_{2}])=\kappa_{0}(2b+c)+\delta_{1}(b_{v_{1}}+b_{v_{2}}).$ (36)
We now sum (35) over $w$ (setting $\overline{w}:=\sum_{w\in\mathcal{W}}w$):
$\begin{split}\overline{w}+pv&=p\kappa_{0}(2b+c)+p\delta_{1}b_{v}+\delta_{1}(b+c)+(\delta_{2}-2\delta_{1})b_{v}\\ &=((p-2)\delta_{1}+\delta_{2})b_{v}+(2p\kappa_{0}+\delta_{1})b+(p\kappa_{0}+\delta_{1})c\end{split}$ (37)
Thus,
$\delta_{1}b_{v}=\frac{\delta_{1}p}{(p-2)\delta_{1}+\delta_{2}}v+\delta_{1}(\overline{w}-(2p\kappa_{0}+\delta_{1})b-(p\kappa_{0}+\delta_{1})c).$ (38)
Thus, setting
$m:=\frac{\delta_{1}p}{(p-2)\delta_{1}+\delta_{2}},\quad d:=2\delta_{1}(\overline{w}-(2p\kappa_{0}+\delta_{1})b-(p\kappa_{0}+\delta_{1})c)+\kappa_{0}(2b+c),$ (39)
we can write
$f([v_{1}],[v_{2}])=m(v_{1}+v_{2})+d.$ (40)
To simplify $d$, we sum (34) over all $w,w^{\prime}$:
$\begin{split}2p\overline{w}&=p^{2}\kappa_{0}(2b+c)+2p\delta_{1}(b+c)+(\delta_{2}-2\delta_{1})c\\ &=2p(p\kappa_{0}+\delta_{1})b+(p^{2}\kappa_{0}+2(p-1)\delta_{1}+\delta_{2})c.\end{split}$ (41)
We further sum (37) over all $v$, setting $\overline{v}:=\sum_{v\in\mathcal{V}\setminus\mathcal{W}}v$:
$((p+q-2)\delta_{1}+\delta_{2}+2pq\kappa_{0})b+q(p\kappa_{0}+\delta_{1})c=q\overline{w}+p\overline{v}.$ (42)
We can now compute $b$ and $c$ from this system of equations and plug it into $d$.
Finally, $\overline{w}=\overline{v}=0$ immediately implies that $b=c=0$ and therefore $d=0$. ∎
D.1.2 Behavior of kernel models on other tasks
Figure 7: Kernel models on dispersed training sets. a, Four example training sets containing two random instances containing each item. b, Generalization loss as a function of the representation’s overlap salience. c,d, Inferred values for c) four example tasks and d) the average across fifty randomly drawn training sets.
To test whether the observed behavior was specific to the kinds of training sets investigated in Proposition 4.4 and the main text, we additionally considered dispersed, randomly generated training sets with two randomly drawn trials containing each item (Fig. 7a). We found that the generalized loss (averaged across all cases) increased roughly linearly with salience (Fig. 7b). For individual training sets, the inferred values were distorted in a complex, irregular manner (Fig. 7c). Further, because the training sets were no longer symmetric, the items took on different values when presented as the first or second component. On average, however, the inferred values were roughly proportional to their actual value and the factor of proportionality still decreased with salience (Fig. 7d).
D.1.3 Behavior of rich networks
Figure 8: Behavior of rich networks on symbolic addition. a, Additivity as a function of initialization scale in networks with one hidden layer. b, Similarity between distinct, identical, and overlapping trials, before and after training in the lazy or rich regime. c, Inferred values in the lazy and rich regime for networks with one hidden layer. d, Zoomed-in inferred values for interpolation part of the task involving interpolation and extrapolation. e, Additivity as a function of the number of hidden layers. f, Inferred values for networks with different numbers of layers.
We first trained rich networks with one hidden layer on the different variants of symbolic addition. We found that they generally were highly additive ($R^{2}>0.992$, Fig. 8a). Fig. 8b confirms that the ReLU networks indeed change their representation in the rich regime: at initialization, the similarity between different trials is approximately clustered by whether those trials are distinct, overlapping, or identical. After lazy training, this remains unchanged, whereas after rich training, the similarities look entirely different. Finally, the inferred values look similar on extrapolation, but do not exhibit a memorization leak on interpolation (Fig. 8c). On the task involving both interpolation and extrapolation, the networks suffer from a memorization leak on the extrapolaton region but not the interpolation region (Fig. 8d). This indicates that rich networks tend to perform better at interpolation, unlike lazy networks which are not sensitive to this difference.
D.1.4 Behavior of deep networks
Next, we trained networks with one to four hidden layers, using $\sigma=0.1$. We found that the networks were generally highly additive ($R^{2}>0.985$, Fig. 8e) and inferred highly similar values. In particular, deeper networks also suffered from a memorization leak on extrapolation but not interpolation.
D.1.5 Behavior of vision models trained on MNIST and CIFAR-10
We now discuss in more detail the behavior of the different vision models.
First, we trained ConvNets on an MNIST version of symbolic addition for the different training sets. For all tasks, we found that compositional generalization was worse than in-distribution generalization (Fig. 9a) and become worse with smaller distance between the digits (i.e. higher conjunctivity). Further all networks were highly additive ($R^{2}>0.98$). Finally, the values inferred by the network were well consistent with the kernel theory: First, they were an affine function of the true item value for all training sets. Second, extrapolation was worse than interpolation and the task involving extrapolation and interpolation. Notably, interpolation and the task involving both yielded equally good values indicating that these networks may also not behave worse just based on extrapolation vs. interpolation. Further, smaller distance between digits (i.e. more conjunctive inputs) generally exacerbated the value misestimation. Finally, we also considered an asymmetric extrapolation task (“Asymmetric”), which trained on all trials containing an item $[-4]$. Consistent with our theory, we found that this resulted in an affine distortion in values.
Figure 9: Behavior of ConvNets trained on MNIST versions of symbolic addition. a, Mean squared error for different training sets and different evaluation sets. b, Additivity on the different tasks. c, Inferred values.
We then trained a ResNet and a ViT on a CIFAR-10 version of the task. We found that the error on the compositional test dataset was consistently higher than on in-distribution generalization (Fig. 10a). Further, similarly to the ConvNets trained on MNIST, it was particularly high on the asymmetric training set. Notably, the additivity of all of these models was reasonably high ($R^{2}>0.93$, Fig. 10b). Finally, we found that the models all inferred values that were distorted in a roughly affine manner (Fig. 10c).
Figure 10: Behavior of ResNets and ViTs on CIFAR versions of symbolic addition. a, Mean squared error over training. b, Additivity. c, Inferred values for the different tasks.
D.2 Context dependence
D.2.1 General task definition
We consider inputs with three components, $(z_{\text{co}},z_{f1},z_{f2})$. We assume that $z_{\text{co}}\in C_{1}\cup C_{2}$, where $C_{1}$ is the set of possible contexts under which $z_{f1}$ is relevant and $C_{2}$ is the set of possible contexts under which $z_{f2}$ is relevant. We further assume that there are decision functions $d_{1}(z_{f1}),d_{2}(z_{f2})\in\mathbb{R}$. (For example, in the example in the main text, these function map three features to the first category (i.e. $y=-1$) and three features to the second category (i.e. $y=1$).) The target is then given by
$y(z_{\text{co}},z_{f1},z_{f2})=\begin{cases}d_{1}(z_{f1})&\text{ if }z_{\text{co}}\in C_{1},\\ d_{2}(z_{f2})&\text{ if }z_{\text{co}}\in C_{2}.\end{cases}$ (43)
Note that in the main text, we consider $C_{1}=\{1\}$, $C_{2}=\{2\}$, and six possible values for $z_{f1},z_{f2}$, where the decision function maps three onto $1$ and three onto $-1$.
D.2.2 Novel stimulus compositions are conjunction-wise additive
If the test set consists in novel combinations of stimuli, this is a conjunction-wise additive computation. Namely, suppose that for all test inputs $(z_{\text{co}},z_{f1},z_{f2})$, the two features have never been observed in conjunction, but both $(z_{\text{co}},z_{f1})$ and $(z_{\text{co}},z_{f2})$ have been. (This includes the case considered in the main text.) In this case, we can define functions $f_{12}$ and $f_{13}$ to implement the appropriate mapping:
$f_{12}(z_{\text{co}},z_{f1})$ :=$:=\begin{cases}d_{1}(z_{f1})&\text{ if }z_{\text{co}}\in C_{1},\\ 0&\text{ if }z_{\text{co}}\in C_{2},\end{cases}\quad f_{13}(z_{\text{co}},z_{f2}):=\begin{cases}0&\text{ if }z_{\text{co}}\in C_{1},\\ d_{2}(z_{f2})&\text{ if }z_{\text{co}}\in C_{2},\end{cases}$ (44)
$f(z_{\text{co}},z_{f1},z_{f2})$ $=f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$ (45)
D.2.3 Novel rule compositions are not conjunction-wise additive
We could also imagine an alternative generalization rule in a task where there are multiple components indicating the same context: $C_{1}=\{\text{co}_{1},\text{co}_{2}\}$ and $C_{2}=\{\text{co}_{3},\text{co}_{4}\}$. We then leave out certain features with certain contexts. For example, suppose we had never seen two values for $z_{f1}$ and $z_{f2}$ in conjunction with $z_{\text{co}}\in\{\text{co}_{2},\text{co}_{4}\}$. In principle, if the model understood that $z_{\text{co}}=\text{co}_{1},\text{co}_{2}$ (and $z_{\text{co}}=\text{co}_{3},\text{co}4$ resp.) signify the same context (i.e. learned to abstract the context from the context cue), it could generalize successfully as it had observed these features in conjunction with $z_{\text{co}}=\text{co}_{1},\text{co}_{3}$. However, the conjunction-wise additive mapping depends on having observed each context in conjunction with each feature and this task is therefore non-additive.
D.2.4 Coefficient groups
In Fig. 3d, we grouped the inferred coefficients into categories. We here explain these categories:
  • Right conj.: This is the correct conjunction the model should use to solve the task, i.e. between $z_{\text{co}}=\text{co}_{1}$ and $z_{f1}$ and between $z_{\text{co}}=\text{co}_{2}$ and $z_{f2}$.
  • Wrong conj.: This is the incorrect conjunction between context and feature, i.e. between $z_{\text{co}}=\text{co}_{1}$ and $z_{f2}$ and between $z_{\text{co}}=\text{co}_{2}$ and $z_{f1}$.
  • Sensory feat.: This is any conjunction involving sensory features, i.e. $z_{f1},z_{f2},(z_{f1},z_{f2})$.
  • Context only: This is the component $z_{\text{co}}$ by itself.
  • Memorization: This is the full conjunction of all three components $(z_{\text{co}},z_{f1},z_{f2})$.
We then compute the average absolute magnitude within each of these groups in order to determine their overall relevance to model behavior.
D.2.5 Rich networks
We find that rich networks generalize consistently on CD-1 and CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 11).
Figure 11: Inferred values for different conjunctive groups on context dependence.
D.2.6 Deep neural networks trained on vision data
We found that convolutional neural networks trained on MNIST successfully generalized on CD-1 and CD-2, but not CD-3 (Fig. 12). For smaller distances between digits, the models tended to generalize worse on CD-1 and CD-2 and gradually reverted to chance accuracy (i.e. 0.5) for CD-3. Further, the networks were generally highly additive ($R^{2}>0.95$), but became worse for lower distance (Fig. 12b). Finally, across all distances, they had a high magnitude associated with the context cue, though this magnitude decreased for small distances — consistent with the accuracy of the network increasing from below chance to chance level (Fig. 12c).
Figure 12: Performance of convolutional neural networks trained on an MNIST version of context dependence. For different distances and training sets, we plot a, the accuracy on different splits, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
Finally, we considered ResNets and ViTs trained on a CIFAR version of the task. We found that they generally performed above chance for CD-2 and CD-1, but below chance for CD-3 (Fig. 13a). Further, their additivity was generally high for ResNets ($R^{2}>0.95$) and slightly lower for ViTs ($R^{2}>0.9$) (Fig. 13b) and both networks had a large magnitude associated with the context cue for CD-3, but not CD-2 or CD-1 (Fig. 13c).
Figure 13: Performance of ResNets and ViTs trained on a CIFAR version of context dependence. We plot a, their accuracy on different splits over training, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
D.3 Transitive equivalence
We additionally trained ReLU networks of various depth on the transitive equivalence (using $\sigma=0.1$ as initialization magnitude). We consistently found that they were able to generalize to the test set (Fig. 14).
Figure 14: Performance of networks with 100 hidden units per layer and varying depth on transitive equivalence. We used an initialization constant $\sigma=0.1$.
D.4 Invariance and partial exposure
We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut distortion. In both cases, the input consists of two components and the mapping only depends on the first (Fig. 15a). In the invariance case, we don’t see the second component vary at all, in the partial exposure case, we see one instance of the second component. Note that the partial exposure task has previously been studied in the context of network generalization [dasgupta_distinguishing_2022, chan_transformers_2022].
To understand the impact of different representational saliences, we consider the generalization margin $m:=y\hat{y}$ on the test set, where $y$ is the ground-truth label and $\hat{y}$ is the model’s estimate. Because we consider support vector machines, the margin on the training set is one; a smaller margin on the test set indicates worse performance. We determined a mathematical formula for the margins as a function of $\text{Sal}_{2}(1)$. Below we first describe its implications and then how we derived this.
Invariance suffers from a memorization leak.
On invariance, we find that the model’s test margin is expressed as $m=\tfrac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}$. This means that for a fully compositional representation ($\hat{\kappa}_{1}=0.5$), its training and test margins are both one. However, as $\hat{\kappa}_{1}$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig. 15b).
Shortcut distortion.
On the partial exposure task, if the model used item 1 to solve the task, it would get two out of three training examples correct and could memorize the last data point. This is a statistical shortcut and we find that norm minimization (just as for context dependence) ends up partially relying on this strategy as this decreases the $\ell_{2}$-norm of the readout weights. As a result, the test margin for the partial exposure task decreases even more strongly as a function of $\text{Sal}_{2}(1)$: $m=\tfrac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig. 15b).
Figure 15: a, Schema for invariance and partial exposure. b, Generalization margin for the two tasks.
Derivations.
We analytically compute the kernel models’ test set prediction on the invariance task. The training set is given by $\{(-1,-1),(-1,1)\}$ and its kernel is therefore
$K=\begin{pmatrix}\kappa_{2}&\kappa_{1}\\ \kappa_{1}&\kappa_{2}\end{pmatrix},$ (46)
where $\kappa_{2}$ is the similarity between identical trials and $\kappa_{1}$ is the similarity between overlapping trials. Hence, the dual coefficients are given by
$a=K^{-1}\begin{pmatrix}1\\ -1\end{pmatrix}=\frac{1}{\kappa_{2}^{2}-\kappa_{1}^{2}}\begin{pmatrix}\kappa_{2}&-\kappa_{1}\\ -\kappa_{1}&\kappa_{2}\end{pmatrix}\begin{pmatrix}1\\ -1\end{pmatrix}=\frac{1}{\kappa_{2}^{2}-\kappa_{1}^{2}}\begin{pmatrix}\kappa_{2}+\kappa_{1}\\ -(\kappa_{2}+\kappa_{1})\end{pmatrix}.$ (47)
The test set is given by $\{(1,-1),(1,1)\}$ and its kernel with respect to the training set is therefore
$\tilde{K}=\begin{pmatrix}\kappa_{1}&\kappa_{0}\\ \kappa_{0}&\kappa_{1}\end{pmatrix},$ (48)
where $\kappa_{0}$ is the similarity between distinct trials. Hence the test set predictions are given by
$\hat{y}=\tilde{K}a=\frac{1}{\kappa_{2}^{2}-\kappa_{1}^{2}}\begin{pmatrix}(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})\\ -(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0}).\end{pmatrix}$ (49)
As the ground truth labels are $y=\{1,-1\}$, the margin $m=y\hat{y}$ is identical for both test set points:
$m=\frac{(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})}{\kappa_{2}^{2}-\kappa_{1}^{2}}=\frac{\kappa_{1}-\kappa_{0}}{\kappa_{2}-\kappa_{1}}=\frac{(\kappa_{2}-\kappa_{0})\text{Sal}_{2}(1)}{(\kappa_{2}-\kappa_{0})-(\kappa_{1}-\kappa_{0})}=\frac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}.$ (50)
For partial exposure, the training set is given by $\{(-1,-1),(-1,1),(1,-1)\}$ and its kernel is therefore
$K=\begin{pmatrix}\kappa_{2}&\kappa_{1}&\kappa_{1}\\ \kappa_{1}&\kappa_{2}&\kappa_{0}\\ \kappa_{1}&\kappa_{0}&\kappa_{2}\end{pmatrix}.$ (51)
The test set is given by $\{(1,1)\}$ and the test set kernel is therefore
$\tilde{K}=\begin{pmatrix}\kappa_{0}&\kappa_{1}&\kappa_{1}\end{pmatrix}$ (52)
The margin is therefore given by
$m=y\hat{y}=-\hat{y}=-\tilde{K}K^{-1}\begin{pmatrix}1\\ -1\\ 1\end{pmatrix}.$ (53)
We solve this equation for the special case where $\kappa_{0}=0$ and $\kappa_{1}=1$ using Mathematica and find that
$m=\frac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}.$ (54)
D.5 Other mathematical operations
We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}[I_{1}]$ and $v_{2}[I_{2}]$ together with some composition function $C(v_{1}[I_{1}],v_{2}[I_{2}])$. This task will only be additive if the composition function is additive (e.g. if it is subtraction). If it is, e.g. multiplication, division, or exponentiation, the task will be non-additive.
D.6 Logical operations
[…]
D.7 Transitive ordering
Transitive ordering is a popular task in cognitive science (often called transitive inference, [mcgonigle_are_1977]). Here the subject is presented with two items $I_{1}$, $I_{2}$ drawn from an unobserved hierarchy $>$. It should then categorize whether $I_{1}>I_{2}$ or $I_{2}>I_{1}$. Crucially, this task can be solved by assigning a rank $r(I_{c})$ to each item and computing the response as $f(I)=r(I_{1})-r(I_{2})$ [lippl_mathematical_2024]. It is therefore additive, in contrast to transitive equivalence. This is also the case if we assume that there are multiple such hierarchies (e.g. $a_{1}>\dotsc,a_{5}$ and $b_{1}>\dotsc,b_{5}$). In this case, the model would generalize to comparisons between these different hierarchies as well.

### v2

@@
-references.bib
+
 Abstract
@@
 disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear.
-To address this gap, we present a general theory of compositional generalization in kernel models with fixed, potentially nonlinear representations (which also applies to neural networks in the “lazy regime”).
-We prove that these models are functionally limited to adding up values assigned to conjunctions/combinations of components that have been seen during training (“conjunction-wise additivity”), and identify novel compositionality failure modes that arise from the data and model structure, even for disentangled inputs.
-For models in the representation learning (or “rich”) regime, we show that networks can generalize on an important non-additive task (associative inference), and give a mechanistic explanation for why.
-Finally, we validate our theory empirically, showing that it captures the behavior of deep neural networks trained on a set of compositional tasks.
-In sum, our theory characterizes the principles giving rise to compositional generalization in kernel models and shows how representation learning can overcome their limitations.
-We further provide a formally grounded, novel generalization class for compositional tasks that highlights fundamental differences in the required learning mechanisms (conjunction-wise additivity).
+To address this gap, we present a general theory of compositional generalization in kernel models with fixed representations, a tractable framework for characterizing the impact of dataset statistics on generalization.
+We find that kernel models are constrained to adding up values assigned to each combination of components seen during training (“conjunction-wise additivity”).
+This imposes fundamental restrictions on the set of tasks these models can learn, in particular preventing them from transitively generalizing equivalence relations.
+Even for compositional tasks that kernel models can in principle learn, we identify novel failure modes in compositional generalization that arise from biases in the training data and affect important compositional building blocks such as symbolic addition and context dependence (memorization leak and shortcut bias).
+Finally, we empirically validate our theory, showing that it captures the behavior of deep neural networks (convolutional networks, residual networks, and Vision Transformers) trained on a set of compositional tasks with similarly structured data.
+Ultimately, this work provides a theoretical perspective on how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
 1 Introduction
 Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant.
-Stitching together concepts in this way allows humans to generalize far beyond our prior experience, preparing us to cope with unfamiliar situations and imagine things that do not yet exist [lake_building_2017, frankland_concepts_2020].
-Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem [fodor_connectionism_1988, battaglia_relational_2018, lake_generalization_2018, hupkes_compositionality_2020, keysers_measuring_2020].
-Accordingly, a broad range of studies have investigated the conditions under which compositionally structured (i.e.
-“disentangled”) representations can be learned [hinton_transforming_2011, bengio_representation_2013, higgins_beta-vae_2017, locatello_challenging_2019, esmaeili_structured_2019, trauble_disentangled_2021, whittington_disentanglement_2023].
-However, it remains unclear whether learning these representations is actually useful: while some work suggests that disentangled representations improve compositional generalization [esmaeili_structured_2019, tokmakov_learning_2019, van_steenkiste_are_2019, whittington_constellation_2021], other studies challenge this view [locatello_challenging_2019, montero_role_2021, schott_visual_2022, xu_compositional_2022].
-In particular, it is often unclear how insights on a certain compositional task generalize to others, as the relationship among them remains unclear [hupkes_compositionality_2020].
-To clarify when compositionally structured representations yield compositional generalization, we define a family of compositional tasks, theoretically analyse the compositional generalization of kernel models, and then validate our theory in several relevant deep neural network architectures.
-Kernel models are an important class of statistical models in their own right (describing e.g.
-linear readout fine-tuning), provide a simplified approximation to neural network learning under certain conditions [jacot_neural_2018, chizat_lazy_2019], and more broadly have provided fundamental insight into generalization behavior in humans and machine learning models [jakel_generalization_2008, jakel_does_2009, canatar_spectral_2021, canatar_out--distribution_2021].
-Despite the broad importance and relative simplicity of kernel models, it remains largely unclear how they generalize on compositional tasks [[, though see ]; see Section 2]abbe_generalization_2023,lippl_mathematical_2024.
+Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Lake et al., 2017; Frankland & Greene, 2020).
+Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018).
+While a wide range of studies have investigated the conditions under which compositionally structured (i.e.
+“disentangled”) representations can be learned (Hinton et al., 2011; Higgins et al., 2017; Träuble et al., 2021; Whittington et al., 2023; Ren et al., 2023), it remains unclear when learning these representations is actually useful to a downstream neural network.
+Some work suggests that disentangled representations improve compositional generalization (Esmaeili et al., 2019; van Steenkiste et al., 2019).
+However, others challenge this view (Locatello et al., 2019; Schott et al., 2022).
+In general, it remains unclear why insights from some compositional tasks fail to generalize to others, and a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
+To make progress on this question, we focus on standard statistical learning, a fundamental basis for compositional generalization that affects almost any machine learning model.
+Generalization in statistical learning depends on the similarities between different inputs.
+Notably, in compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square).
+As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the generalization behavior of the model.
+To understand when and how statistical learning leads to successful compositional behavior, we developed a theory for the behavior of kernel models on compositional tasks.
+Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008, 2009).
+For example, kernel models accurately describe the behavior of learning and fine-tuning in neural networks under certain conditions (Jacot et al., 2018; Malladi et al., 2023).
+More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a, b).
+Despite their broad relevance and relative simplicity, it has remained unclear how even kernel models generalize on compositional tasks (though see Abbe et al., 2023; Lippl et al., 2024, see Section 2).
 To address this gap, we present a theory of compositional generalization in kernel models.
-We define a new generalization class for compositional tasks and show that characterizing kernel models on such tasks helps us understand realistic neural network architectures.
+We define a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory), theoretically characterize the compositional behavior of kernel models trained on such tasks, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures.
 Our specific contributions are as follows:
-  • We prove fundamental limits on the generalization behavior of kernel models with disentangled inputs, showing that they are constrained to summing up values implicitly assigned to each component or combination of components seen during training.
-We define the class of compositional tasks with this form as “conjunction-wise additive” (Section 4.1).
-  • We define a metric to quantify how salient different component “conjunctions” (entangled features unique to component combinations) are in the representation (Section 4.2).
-  • For conjunction-wise additive tasks, we then highlight two important failure modes impacting generalization (memorization leak and shortcut bias) (Section 4.3), which describe how biases in the training data can limit compositional generalization even for kernel models with disentangled inputs.
-  • We validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks (Section 4.4).
-  • Finally, we show how feature learning can overcome certain limitations of kernel models and learn non-additive tasks (Section 4.5).
+  • In Section 4, we show that kernel models with disentangled inputs are constrained to summing up values implicitly assigned to each component or combination of components seen during training.
+We call such computations “conjunction-wise additive.”
+  • In Section 5, we then characterize how representational geometry determines whether kernel models will generalize successfully on conjunction-wise additive compositional tasks, highlighting two important failure modes (memorization leak and shortcut bias).
+This highlights that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
+  • Finally, in Section 6, we validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks.
+Overall, we take a step towards a general theory of compositional generalization by characterizing the impact of dataset statistics on compositional generalization for a broad range of tasks.
+Our theory systematically clarifies the compositional generalization behavior of deep neural networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
 2 Related work
 Compositional generalization.
-Compositionality is an important theme across human and machine reasoning problems, including visual reasoning [lake_human-level_2015, johnson_clevr_2017, marino_ok-vqa_2019, schwartenbeck_generative_2023, zhou_compositional_2024], language production [hupkes_compositionality_2020], rule learning [ito_compositional_2022, abdool_continual_2023], image generation [okawa_compositional_2023], and robotics [zhou_policy_2022].
-Recent breakthroughs (especially in language and computer vision) have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly [srivastava_beyond_2023, lewis_does_2023, west_advances_2023, ma_crepe_2023, chen_skills--context_2023].
-Attempts to improve compositional generalization often leverage meta-learning [lake_human_2019, mitchell_challenges_2021, kumar_meta-learning_2021, wu_adaptive_2023, lake_human-like_2023], or modular architectures, in the hopes that different modules will specialize for different components [andreas_neural_2017].
-End-to-end training of these architectures often does not result in the desired modular specialization [bahdanau_systematic_2019, mittal_is_2022, jarvis_specialization_2023]; notably, [schug_discovering_2024] theoretically clarify the necessary conditions for correct modular specialization and compositional generalization in hypernetworks.
-Here we highlight that kernel models add up values assigned to each combination of features (or “conjunction”) seen during training.
-Notably, constraining a network to be additive can be sufficient for compositional generalization [lachapelle_additive_2023, wiedemer_compositional_2023].
-Even without such constraints, language models encode many semantic concepts in an additive manner [mikolov_distributed_2013, naito_revisiting_2021].
-Conjunctive codes (representations specific to a particular feature combination), on the other hand, have a long history in neuroscience [alvarado_properties_1992, baker_impact_2002], and are theoretically linked to forming highly specific, episodic memories.
-Kernel and rich regime.
-Prior work has revealed two key strategies by which deep neural networks learn [chizat_lazy_2019, woodworth_kernel_2020].
-In the kernel (or “lazy”) regime (which can be induced with large initial weights or wide networks), the networks’ learning is well approximated by gradient descent on a model with a fixed representation [jacot_neural_2018].
-In the rich regime (brought forth, for example, by small initial weights, small width, low-rank connectivity, or loss functions like cross-entropy), the networks learn structured (i.e.
-abstract and sparse) representations over the course of learning [savarese_how_2019, woodworth_kernel_2020, chizat_implicit_2020, lyu_gradient_2020, saxe_neural_2022, liu_how_2023, domine_exact_2023].
-This is often suggested to give rise to better generalization and more human-like behavior in neural networks [fort_deep_2020, vyas_limitations_2022, flesch_orthogonal_2022].
-When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(x)=w^{T}r(x)$, this model depends on its representation $r(x)$ only through its induced kernel $K(x,x^{\prime})=\langle r(x),r(x^{\prime})\rangle$ [scholkopf_kernel_2000].
-Specifically, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, it can be written in its “dual form” as
-$f_{a}(x)=\sum_{i=1}^{n}a_{i}K(x,x_{i}).$ (1)
-Gradient descent and ridge regression learn the readout weights that describe the training data with minimal $\ell_{2}$-norm; this is also true of neural networks in the kernel regime [soudry_implicit_2018, gunasekar_characterizing_2018, ji_gradient_2020].
-Norm minimization is a standard theoretical framework for analyzing how representational geometry influences generalization [canatar_spectral_2021, canatar_spectral_2023] and we here apply it to the compositional task setting.
-This is similar to [lippl_mathematical_2024] who characterize model behavior on a specific compositional task (transitive ordering) and [abbe_generalization_2023] who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components.
-Compared to [abbe_generalization_2023, lippl_mathematical_2024], we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
-[…]
-Shortcut learning refers to models exploiting spurious correlations between certain features and the target to learn the training data and has been essential for understanding why the performance of deep neural networks can drop substantially when tested out-of-distribution [geirhos_shortcut_2020].
-Models are more likely to use shortcut features that are easily extracted from the input and highly predictive of the output [hermann_what_2020, tartaglini_developmentally-inspired_2022, hermann_foundations_2023].
-Shortcut learning has also been connected to the implicit bias of gradient descent [shah_pitfalls_2020, nagarajan_understanding_2021].
-Our findings highlight a particular compositional shortcut: models tend to use a more salient, partially predictive component conjunction (memorizing the remaining data) over the fully predictive but less salient ground-truth conjunction.
+Compositionality is an important theme in both human and machine reasoning problems, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023).
+Recent breakthroughs have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023).
+Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017), in the hopes that different modules will specialize for different components (Schug et al., 2024).
+Constraining a network to be linear (or additive) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a, b).
+However, these networks are extremely limited in the tasks they can learn (as we will show below) and end-to-end training of modular architectures without such constraints often does not result in the desired specialization (Bahdanau et al., 2019; Mittal et al., 2022; Jarvis et al., 2023).
+Conjunctions.
+Henceforth, we will call a combination of components a conjunction.
+Conjunctive codes have a long history in neuroscience (Alvarado & Rudy, 1992; Baker et al., 2002), and are theoretically linked to the formation of highly specific episodic memories.
+Kernel regime.
+When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(z)=w^{T}\phi(z)$, it can be shown that the resulting behavior of that model can be described in terms of the kernel $K(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ induced by the representation $\phi(z)$ (Schölkopf, 2000).
+We will refer to such models as kernel models.
+Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, a kernel model can be written in its “dual form” as
+$f_{a}(z)=\sum_{i=1}^{n}a_{i}K(z,z_{i}),\mbox{ for dual coefficients}a\in\mathbb{R}^{n}.$ (1)
+Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by gradient descent on a model with a fixed representation (a kernel model) (Jacot et al., 2018).
+This is called the “kernel regime,” in contrast to the feature-learning regime, which is brought forth, for example, by small initial weights or small width (Chizat et al., 2019).
+Norm minimization.
+Gradient descent, ridge regression, and neural networks in the kernel regime learn the readout weights that describe the training data with minimal $\ell_{2}$-norm (Soudry et al., 2018; Gunasekar et al., 2018; Ji et al., 2020).
+Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021b, 2023) and we here apply it to the compositional task setting.
+This is similar to Lippl et al.
+(2024) who characterize model behavior on a specific compositional task (transitive ordering) and Abbe et al.
+(2023) who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components.
+Compared to these prior works, we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
+Memorization leak.
+Machine learning models sometimes memorize (parts of) their training data instead of learning a generalizable rule (Zhang et al., 2020; Dasgupta et al., 2022).
+This may improve generalization on long-tailed data (Feldman, 2020), but often impairs out-of-distribution performance (Elangovan et al., 2021).
+We find that models partially memorize their training data even when they extract the correct rule.
+Jarvis et al.
+(2023) analyze a related phenomenon in deep linear networks.
+Shortcut learning.
+Shortcut learning refers to models exploiting spurious correlations between certain features of the input data and the target (Shah et al., 2020; Nagarajan et al., 2021).
+This substantially impacts their performance out of distribution (where those correlations may not hold) (Geirhos et al., 2020).
+We theoretically analyze how shortcut biases impact compositional generalization.
 3 Model and task setup
 3.1 Task space
-We assume that our model extracts a set of components $z=\sum_{c=1}^{C}z_{c}$ from its input $x$.
-Each $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components.
-For example, $z$ could specify the color and shape of an object, and the background color of an image (Fig.
+We assume that our model extracts from its input a set of disentangled components $(z_{c})_{c=1}^{C}$, e.g.
+color, shape, and background (Fig.
 1a).
-Alternately, $z$ may express different words in an input sentence (“the red sphere is in front of the blue background”).
-We assume that each component $z_{c}$ is represented in an orthogonal subspace (e.g.
-by concatenating representations of individual components, Fig.
+For instance, the model could have been pre-trained to extract these components from an image or an input sentence.
+Each component $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components represented by $d$-dimensional real vectors and because we consider categorical components, we assume that all components have unit magnitude and are orthogonal to each other (these assumptions can be relaxed, Section A.1).
+We then assume that the model represents these components by adding them together: $z:=\sum_{c=1}^{C}z_{c}$.
+A simple example of this is given by a concatenation of one-hot representations of all components (Fig.
 1a).
-Further, because we consider categorical components, we can assume that for each component $c$, all representations $z_{c}\in Z_{c}$ have equal magnitude $m_{c}$ and are represented with equal similarity $\rho_{c}$ to alternative component values.
-The target, $y\in\mathbb{R}$, is some function of $z$ (see Section 3.3).
-After training models on certain component combinations $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations, $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$ to understand how biases in the training data affect compositional generalization.
+The target, $y\in\mathbb{R}$, is given by an arbitrary function of $z$, ensuring that our framework is agnostic to the underlying compositional structure.
+After training models on certain combinations of components $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$.
+Our theory applies to any task in this family; that is, any task with a fixed number of categorical input components and a scalar output, regardless of the input-output mapping.
+Below we describe three example tasks within this general framework that represent important building blocks for compositional reasoning in machine learning and cognitive science.
+We will focus on these tasks in the main text, to illustrate our general theory and analyze its consequences in more detail.
+In Appendix D, we describe additional tasks captured by our theory.
+Symbolic addition.
+Many tasks involve inferring a magnitude associated with different underlying components (e.g.
+adding handwritten digits) (Lorenzi et al., 2021; Sheahan et al., 2021).
+We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$.
+The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$.
+After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items.
+In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values (given by the rows and columns in the training set in Fig.
+1b).
+We vary the size of this subset (i.e.
+the number of rows/columns) as well as whether the task requires only interpolation or also extrapolation.
+Context dependence.
+The relevance of different stimuli often depends on our current context.
+Taking this into account is a crucial aspect of cognition in humans and animals (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022).
+We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$.
+The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response.
+Both features have six possible values which are split up in two categories.
+If the model has learned this context dependence, it should be able to generalize to novel feature combinations.
+We evaluate on the subset of data for which $z_{f1}$ indicates Cat.
+2 and $z_{f2}$ indicates Cat.
+1.
+In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of two or one of those features (CD-2 and CD-1) (Fig.
+1c).
+Transitive equivalence.
+Relational reasoning is an important example of compositional generalization and often involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018).
+Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent.
+The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig.
+1d).
+This is an important instance of relational cognition (often studied as “associative inference” in cognitive science (Schlichting & Preston, 2015; Spalding et al., 2018)).
+Although prior work has found that kernel models often successfully generalize on a transitive ordered relation ($A>B$ and $B>C$ imply $A>C$, Lippl et al., 2024), the behavior of kernel models on equivalence relations has remained unclear.
 3.2 Models
-Linear readout models.
-We first consider kernel models which take in a disentangled representation $z$, apply a nonlinear transform $\phi(z)\in\mathbb{R}^{h}$ and learn a linear readout using gradient descent or ridge regression, $f_{w}(x):=w^{T}\phi(z)$ (Fig.
+Kernel models.
+Our theory characterizes kernel models with a disentangled input $z\in\mathbb{R}^{d}$ that apply a transform $\phi(z)\in\mathbb{R}^{h}$ and learn a linear readout, $f_{w}(x):=w^{T}\phi(z)$, using gradient descent or ridge regression (Fig.
 1a).
-For $\phi$, we consider a neural network with random weights:
-Definition 3.1.
-Given an input $z\in\mathbb{R}^{d}$, a network depth $L\geq 2$, and a set of widths $H_{1},\dotsc,H_{L}\in\mathbb{N}$, we define a neural network by recursively defining the operation $\phi^{(l)}$ of the $l^{\text{th}}$ layer as
-$a^{(0)}:=z,\quad a^{(l)}:=\phi^{(l)}(a^{(l-1)}):=\sigma\left(\tfrac{1}{\sqrt{H_{l}}}\left(W^{(l)}a^{(l-1)}+b^{(l)}\right)\right),\quad W^{(l)}\in\mathbb{R}^{H_{l}\times H_{l-1}}$ (2)
-where $W^{(l)}$ and $b^{(l)}$ are i.i.d.
-sampled from a random distribution and $\sigma:\mathbb{R}\to\mathbb{R}$ is a nonlinearity.
-The complete transform is then given by $\phi:=\phi^{(L)}\circ\dotsb\circ\phi^{(1)}$.
-[…]
-Deep neural network models.
-Additionally, we consider different deep neural network architectures used for vision (convolutional networks, ResNets, and Vision Transformers, Section 4.4) and ReLU networks trained through backpropagation (Section 4.5).
-Figure 1: a, Theoretical setup: we consider a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$.
-b, Training sets for symbolic addition.
+For $\phi$, we consider a neural network with random weights (in the infinite-width limit; see Section A.3).
+Our model setup includes as a special case the random feature model studied by Abbe et al.
+(2023) and further captures training via backpropagation in the kernel regime.
+Deep neural networks.
+Additionally, we consider a broad range of deep neural network architectures: convolutional networks, residual networks (He et al., 2016), Vision Transformers (ViTs; Dosovitskiy et al., 2020)), and ReLU networks; all trained through backpropagation.
+The code required to reproduce all experiments can be found at https://github.com/sflippl/compositional-generalization.
+Figure 1: a, Theoretical setup: a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$.
+b, Example training sets for symbolic addition.
 The grid represents nine components with associated values $-4,-3,\dotsc,4$.
-Each tile represents a data point in the training set.
+The training set consists of certain rows and columns of this grid.
 c, Context dependence.
@@
 A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes.
-3.3 Example tasks
-In addition to deriving a general theory, we consider a set of example tasks that are important building blocks for compositional reasoning in machine learning and cognitive science (Appendix D discusses additional tasks including logical operations and partial exposure):
-Symbolic addition.
-Many tasks involve inferring a magnitude associated with different underlying components (e.g.
-handwritten digits) [lorenzi_numerosities_2021, sheahan_neural_2021].
-We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(I_{1})$ and $v_{2}(I_{2})$.
-The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$.
-After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items [wiedemer_compositional_2023].
-In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values.
-This can require interpolation, extrapolation, or both (Fig.
-1b).
-Context dependence.
-The relevance of different stimuli often depends on the context we are in.
-Taking into account this context is a crucial aspect of cognition in humans and animals [bouton_stimulus_1999, taylor_context-dependent_2013, parker_cognitive_2014, ito_compositional_2022].
-We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$.
-The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response.
-Both features have six possible values which are split up in two categories.
-If the model has learned this context dependence, it should be able to generalize to novel feature combinations.
-We evaluate on the orthant for which $z_{f1}$ indicates Cat.
-2 and $z_{f2}$ indicates Cat.
-1.
-In the most extreme generalization test (CD-3), we leave out the entire orthant, in easier versions we leave out conjunctions of two or one of those features (CD-2 and CD-1) (Fig.
-1c).
-Transitive equivalence.
-Relational reasoning is an important instance of compositional generalization and often involves extending the learned relations to new item combinations [halford_relational_2010, battaglia_relational_2018].
-Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent.
-The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig.
-1d).
-This is an important instance of relational cognition (often studied as “associative inference” in cognitive science [schlichting_memory_2015, spalding_ventromedial_2018]).
-Note that prior work has found that kernel models often successfully generalize a transitive ordering relation [lippl_mathematical_2024].
-4 Results
-4.1 Kernel models with compositional structure are conjunction-wise additive
-Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs: specifically, they assign a value to each combination of components seen during training and generalize to test inputs by adding up the relevant combinations’ values.
-We call this motif “conjunction-wise additivity.” Below we formally state our finding and then explain its implications.
-We first note that the kernel of the disentangled representation, $K(z,z^{\prime})=z^{T}z^{\prime}$, only depends on the components for which the two inputs are overlapping, $O(z,z^{\prime}):=\left\{c|z_{c}=z_{c}^{\prime}\right\}$ where $c\in\{1,\dotsc,C\}$.
-This is necessitated by the fact that separate components are represented in orthogonal subspaces.
-We call any representation satisfying this criterion “compositionally structured.” Notably, the hidden representation $\phi(z)$ is also compositionally structured (in the infinite-width limit).
-This is because the kernel of the neural network’s representation, $K_{\phi}(z,z^{\prime}):=\phi(z)^{T}\phi(z)$, only depends on the inputs $z,z^{\prime}$ through their kernel $K(z,z^{\prime})$ and thus conserves this property [[]; see Appendix B]cho_kernel_2009,tsuchida_invariance_2018,tsuchida_richer_2019,han_fast_2022.
-Intriguingly, we find that any kernel model with a compositionally structured representation is constrained to be conjunction-wise additive.
-To state this finding, we define, for each $z\in Z$, the set of conjunctions for which $z$ overlaps with some element in the training set $z^{\text{tr}}\in Z^{\text{train}}$,
-$O(z|Z^{\text{train}}):=\left\{S\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}\text{ s.t. }S\subseteq O(z,z^{\text{tr}})\right\}.$ (3)
+4 Kernel models are conjunction-wise additive
+Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs.
+We find that even though these models can learn arbitrary training sets, their test set behavior is restricted to adding up values assigned to each conjunction (combination of components) seen during training.
+We call this motif “conjunction-wise additivity.” Below, we formally state our finding and explain its implications.
+4.1 Random networks yield compositionally structured representations
+We first note that a linear readout of the disentangled input representation effectively constrains the model to adding up a value for each component (“component-wise additivity”):
+$f(z)=w^{T}z=w^{T}\sum_{c=1}^{C}z_{c}=\sum_{c=1}^{C}f_{c}(z_{c}),\quad f_{c}(z_{c}):=w^{T}z_{c}.$ (2)
+Although models of this class perfectly generalize on component-wise additive tasks, such as symbolic addition, they are incapable of even learning the training data (much less generalizing) on tasks that are not component-wise additive.
+In particular, these models cannot learn context-dependent computations or equivalence relations — both fundamental instances of compositional reasoning.
+The nonlinear transform $\phi(z)$ can overcome this constraint; in particular, multi-layer nonlinear neural networks can learn arbitrary training data (in the infinite-width limit) (Hornik et al., 1989; Cybenko, 1989; Rigotti et al., 2013).
+In general, the wide range of possible neural network architectures and their nonlinearities makes it difficult to derive general statements about their representations, but in this case we can take advantage of a particularly useful property of disentangled categorical inputs.
+That is, its induced kernel, $K(z,z^{\prime})=z^{T}z^{\prime}$, depends only on the number of components that are identical between both inputs (specifically, $K(z,z^{\prime})=c$ if $c$ components are identical).
+We call any representation that satisfies this criterion “compositionally structured” (see Section A.2 for further motivation).
+Notably, in the infinite-width limit, the kernel of random weight neural networks, $K_{\phi}(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$, depends only on the input kernel $K(z,z^{\prime})$, making $K_{\phi}(z,z^{\prime})$ also compositionally structured (Cho & Saul, 2009, see Section A.3).
+This representational constraint allows us to derive strong constraints on the models’ generalization behavior.
+4.2 Compositionally structured kernel models are conjunction-wise additive
+We find that any kernel model with a compositionally structured representation is constrained to be conjunction-wise additive.
+To formally state this finding, we define, for each $z\in Z$, the set of conjunctions for which $z$ overlaps with some element in the training set $z^{\text{tr}}\in Z^{\text{train}}$,
+$\text{Conj}(z|Z^{\text{train}}):=\left\{J\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}:\forall c\in J:z_{c}=z^{\text{tr}}_{c}\right\}.$ (3)
 We then prove:
 Theorem 4.1.
-For $z\in Z$, any kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$ can be expressed as a sum of functions $f_{S}$ over all conjunctions $S\in O(z|Z^{\text{train}})$, where each $f_{S}$ associates a value with each unique combination of features $z_{c}$ for $c\in S$:
-$f(z)=\sum_{S\in\mathcal{O}(z|Z^{\text{train}})}f_{S}(z_{S}),\quad z_{S}:=(z_{c})_{c\in S}.$ (4)
-We prove the theorem in Appendix A.
-Intuitively, it holds because the model associates a distinct weight vector with each conjunction/combination of components.
-The weight vectors associated with conjunctions not seen during training remain at their initial value and can therefore not be used.
-To understand the theorem’s implications, we first consider a direct linear readout, $f(z)=w^{T}z$. This constrains the model to adding up a value for each component (“component-wise additivity”):
-$f(z)=w^{T}z=w^{T}\sum_{c=1}^{C}z_{c}=\sum_{c=1}^{C}f_{c}(z_{c}),\quad f_{c}(z_{c}):=w^{T}z_{c}.$ (5)
-Component-wise additivity is a strict subset of conjunction-wise additivity (each $f_{S}$ depending on more than one component is set to zero).
-It lets the model generalize perfectly on component-wise additive tasks like symbolic addition.
-However, it also prevents the model from learning any training set that is not component-wise additive, including context dependence and transitive equivalence.
-The nonlinear transform $\phi$ can overcome this constraint [hornik_multilayer_1989, cybenko_approximation_1989, leshno_multilayer_1993, rigotti_importance_2013].
-Indeed, conjunction-wise additive functions more broadly can learn arbitrary training data, as the full conjunction $S=\{1,\dotsc,C\}$ is in $O(z|Z^{\text{train}})$ for each $z\in Z^{\text{train}}$ and $f_{S}(z)$ can take on a distinct value for each $z$.
-However, for $z\in Z^{\text{test}}$ the full conjunction is not in $O(z|Z^{\text{train}})$.
-Thus, while conjunction-wise additivity (unlike component-wise additivity) does not impose a constraint on the kinds of training data the model can learn, it does impose constraints on the kinds of generalizations it can implement.
-In particular, for inputs with two components, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ for $z\in Z^{\text{train}}$.
-However, for $z\in Z^{\text{test}}$, $f_{12}(z)$ falls away and kernel models are constrained to an additive computation: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
-Thus, they cannot generalize on any non-additive task, including transitive equivalence: while they learn the training data (using $f_{12}(z)$), they cannot generalize to the test data.
-For inputs with more than two components, a conjunction-wise additive computation does not just encode each component and their full conjunction, but also partial conjunctions between components (e.g.
+Consider a kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$.
+For any such model, we can find conjunction-wise functions $f_{J}:\prod_{c\in J}Z_{c}\to\mathbb{R}$, where $J\subseteq\{1,\dotsc,C\}$, such that for any input $z\in Z$, the model response is given by
+$f(z)=\sum_{J\in\text{Conj}(z|Z^{\text{train}})}f_{J}(z_{J}),\quad z_{J}:=(z_{c})_{c\in J}.$ (4)
+We prove Theorem 4.1 in Section A.4, but to develop an intuition for why this theorem holds, we consider an example representation concatenating one-hot vectors representing each component and conjunction of components.
+For example, if $C=2$, this representation will concatenate one-hot representations of each component with a one-hot representation of the conjunction of both components.
+The readout weight for a conjunction $z_{J}$ seen during training will change (in Eq.
+4, this weight is given by $f_{J}(z_{J})$).
+In contrast, the weights associated with conjunctions not seen during training remain at their initial value of zero (which is why we only sum over $J\in\text{Conj}(z|Z^{\text{train}})$).
+4.3 Conjunction-wise additivity constrains the tasks kernel models can solve
+Theorem 4.1 immediately implies that compositionally structured kernel models cannot solve any task that cannot be expressed in conjunction-wise additive terms.
+This highlights fundamental computational restrictions on a broad range of models — including pre-trained models with disentangled representations that are finetuned in the kernel regime and infinite-width neural networks.
+Intriguingly, these restrictions are not caused by an architectural constraint, as the model can learn arbitrary training data.
+Rather, we highlight restrictions on how models – without any architectural constraints – can generalize (Zhang et al., 2021).
+Below we spell out the consequences of these restrictions.
+First, for inputs with two components, the model’s behavior on the training set can be expressed as $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$.
+However, on the testing set, $f_{12}(z)$ falls away and the model is constrained to be component-wise additive: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
+A striking consequence of this is that compositionally structured models – including infinite-width neural networks and regardless of implementational details – can only generalize on component-wise additive tasks.
+In particular, this implies that these models cannot generalize on transitive equivalence — a fundamental instance of relational reasoning.
+Intriguingly, transitive ordering – a superficially similar task – can be solved by a kernel model.
+This highlights the importance of a formal perspective on compositional tasks.
+For inputs with more than two components, a conjunction-wise additive computation additionally encodes partial conjunctions between components (e.g.
 12, 13, and 23 for three components).
-For test inputs, the model can still not rely on the full conjunction (123), but it can rely on any partial conjunction it may have seen before.
-To see whether such a model can, in principle, generalize on a given task, we must 1) identify the overlaps with the training set, $\mathcal{O}(z|Z^{\text{train}})$, and 2) determine whether the target can be written as a conjunction-wise sum.
-For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions (i.e.
-12 and 13), but not the feature-feature conjunction (i.e.
-23).
-As a result, model behavior on the test set can be written as $f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2})$.
-While the task is not component-wise additive, we can use $f_{12}$ and $f_{13}$ to encode the task: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the first feature $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
-Hence, conjunction-wise additivity not only tells us whether kernel models with compositionally structured representations can solve a certain task, but also how they solve it.
-Further, it tells us to make a task non-additive: in Section D.2.3, we describe a modification that makes context dependence non-additive by requiring generalizations to novel context-feature conjunctions.
-Importantly, conjunction-wise additivity only determines whether certain compositional generalizations are, in principle, feasible.
-Model generalization also depends on whether the model identifies the correct conjunction-wise function.
-We discuss this in the next two sections.
+For test inputs, the model cannot rely on the unseen full conjunction 123, but it can rely on any partial conjunction it has seen before.
+To see whether a kernel model can, in principle, generalize on such a task, we must 1) identify the conjunctions seen during training, $\text{Conj}(z|Z^{\text{train}})$, for each $z\in Z^{\text{test}}$, and 2) determine whether the target can be written as a conjunction-wise sum.
+For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions 12 and 13, but not the feature-feature conjunction 23.
+As a result, model behavior on the test set can be written as
+$f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$
+While context dependence is not component-wise additive, it is conjunction-wise additive.
+This is because we can encode it using $f_{12}$ and $f_{13}$: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
+Thus, the model can, in principle, identify a set of weights that would allow it to generalize correctly.
+This example illustrates that conjunction-wise additivity tells us both whether kernel models with compositionally structured representations can solve a certain task, and also how they solve it.
+As we noted, these insights immediately transfer to any model with compositionally structured representations that is trained or fine-tuned in the kernel regime.
+Because our theory determines the compositional computations emerging from dataset statistics, it is also likely to shed light on other learning models, including feature-learning deep neural networks (see Section 6).
+Lastly, our theory tells us how to make a task non-additive: in Section D.2.3, we describe such a modification to context dependence that tests generalization to novel context-feature conjunctions.
+By helping us design hard benchmark tasks that are not solvable by kernel models, our framework grounds research into learning mechanisms that can implement other compositional computations.
+5 Representational geometry still impacts generalization on conjunction-wise additive tasks in kernel models
+Conjunction-wise additivity only determines whether kernel models can, in principle, solve a certain task.
+However, model generalization also depends on whether the model identifies the correct conjunction-wise function.
+This depends on the model’s representational geometry and the dataset statistics.
+As noted in Section 2, the behavior of kernel models depends on their representation $\phi(z)$ only through the induced kernel $K_{\phi}(z,z^{\prime})$.
+In this section, we employ the modular approach suggested by this insight: We first investigate how different choices in network architecture influence its induced kernel (Section 5.1).
+Then, we characterize task performance on symbolic addition and context dependence across the full range of compositionally structured representations (Section 5.2).
 5.1 Overlap salience characterizes compositional representational geometry
-Model generalization is determined by how prominently different component conjunctions are represented.
-This is measured by the similarity $K(S)$ between inputs with overlap $S\subseteq\{1,\dotsc,C\}$.
-$K(S)$ depends on responses to that particular conjunction $S$ as well as all of its sub-conjunctions $S^{\prime}\subsetneq S$.
-We thus define the scalar salience of an overlap $S$ as its unique contribution.
-We then normalize this salience, so all saliences add up to one:
-Definition 4.2.
-For a representation with kernel $K$ and a conjunction $S\subseteq\{1,\dotsc,C\}$, we define
-$\overline{\text{Sal}}(\emptyset):=K(\emptyset),\quad\overline{\text{Sal}}(S):=K(S)-\sum_{S^{\prime}\subsetneq S}\overline{\text{Sal}}(S^{\prime}),\quad{\text{Sal}}(S)=\tfrac{\overline{\text{Sal}}(S)}{\sum_{\emptyset\neq S\subseteq\{1,\dotsc,C\}}\overline{\text{Sal}}(S)}.$ (6)
-Figure 2: Normalized salience (i.e. $\text{Sal}_{C}(k)$) for neural networks with random weights and different nonlinearities for inputs with three components.
-We now analyze how the representational salience evolves over different layers of a random network, considering an uncorrelated disentangled correlation with equal magnitudes (i.e.
-$\rho_{c}=0$ and $m_{c}=m_{c^{\prime}}$) (see Section B.2 for analysis beyond these assumptions).
-In that case, the salience only depends on the overlap size $k:=|S|$ and we denote it by $\text{Sal}_{C}(k)$.
-Notably, conjunctions are only encoded and can only be used if they have non-zero salience.
-In particular, the disentangled representation has zero salience for all conjunctions of more than one component.
-As the network gets deeper, the salience of these conjunctions increases.
-In fact, for (leaky) ReLU networks, the salience of the full conjunction converges to one:
+Figure 2: Representational salience (for three input components) in a random weight neural network with variable numbers of layers and nonlinearities.
+In Section 4, we found that a broad range of transformations yield compositionally structured representations, i.e.
+representations that have the same similarity $K(z,z^{\prime})$ for all inputs $z,z^{\prime}$ overlapping in the same number of components.
+To characterize representations in this space, we introduce a new metric: representational salience.
+We formally define this metric in Section B.1.
+Intuitively, for $c=1,\dotsc,C$, the salience $S(c;C)$ measures the unique contribution of the subpopulation representing a conjunction of $c$ components.
+Further, $S(c;C)$ is normalized so that the salience of all components and conjunctions of components sums up to one.
+For example, the input representation only encodes single components and therefore $S(1;C)=1/C$ and, for all $c>1$, $S(c;C)=0$.
+$S(c;C)$ is computed from the representational similarities $K(z,z^{\prime})$.
+It is useful because unlike those similarities, it makes immediately apparent how strongly different conjunctions are encoded (Section B.2).
+Further, it reduces the set of $C+1$ similarities characterizing the space of compositionally structured representations to $C-1$ free parameters: It removes the overall magnitude of the representation (by normalizing) and the baseline activity (by excluding the similarity between inputs with no overlap).
+This choice reflects the fact that both of these parameters often have negligible impact on model behavior (Section B.3).
+We now analyze how $S(c;C)$ evolves over different layers of a random network, using an input with three components as an example (Fig.
+2; see Section B.4 for other values of $C$).
+At the input stage, only individual components are represented and so $S(2;3)$ and $S(3;3)$ are zero.
+As the network gets deeper, these saliences increase, whereas $S(1;3)$ decreases.
+Eventually, the salience of the intermediate conjunction $S(2;3)$ decreases again, whereas $S(3;3)$ continues to increase.
+Indeed, for ReLU networks, we prove the following statement:
 Proposition 5.1.
-For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $\text{Sal}_{C}(k)\to 0$ for $k<C$ and $\text{Sal}_{C}(C)\to 1$.
-We prove this statement in Section B.1.
-Empirically, we find that $\text{Sal}_{C}(C)$ indeed increases with depth, whereas $\text{Sal}_{C}(1)$ decreases (Fig.
-2).
-The salience of intermediate conjunctions (e.g.
-$\text{Sal}_{3}(2)$) first increases and then decreases, its trajectory depending on the nonlinearity.
-4.3 Kernel models suffer from two failure modes: memorization leak and shortcut bias
-In Section 4.1, we found that kernel models sum up values assigned to different component conjunctions.
-This leaves ambiguous, however, whether they generalize correctly, as they can learn the training data using different combinations of conjunctions.
-Notably, kernel models trained with gradient descent or ridge regression generally learn the readout weights with minimal $\ell_{2}$-norm [[, Section 2;]]soudry_implicit_2018,gunasekar_characterizing_2018,ji_gradient_2020.
-This inductive bias gives rise to weights that are distributed across different predictive features.
-For in-distribution generalization, this is often useful, as it allows us to integrate different kinds of evidence.
-However, for compositional generalization, we often need to rely on a small group of conjunctions and the tendency towards distributed weights causes generalization failures.
-Here we highlight two particular failure modes arising from this.
-First, the full conjunctions always serve as predictive features (“memorization leak”).
-Second, the model can often use statistical shortcuts to predict parts of the training set, using the full conjunction to learn the rest (“shortcut bias”).
-Below we explain how these failure modes impact symbolic addition and context dependence.
+For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $S(k;C)\to 0$ for $k<C$ and $S(C;C)\to 1$.
+We prove the proposition in Section B.5. It implies that in very deep networks, the full conjunction dominates the representation, making compositional generalization to overlapping inputs harder.
+5.2 Kernel models suffer from memorization leak and shortcut bias
+While there are many readout weights that can lead to minimal error, kernel models trained with gradient descent or ridge regression learn the weights with minimal $\ell_{2}$-norm (see Section 2).
+We now characterize how this inductive bias influences compositional generalization across different representational geometries.
+To characterize the full range of compositionally structured representations, we compute the behavior of the model directly using the kernel (see Appendix C).
+Importantly, as implied by our modular approach, our analysis clarifies the behavior of all random weight neural networks, including those covered in the previous section.
 Symbolic addition suffers from a memorization leak.
-The functional form $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ implies that using the full conjunction $f_{12}(z)$ to learn the training set will necessarily distort the inferred values $f_{1}(z_{1}),f_{2}(z_{2})$.
-This, in turn, negatively impacts generalization on the test set.
-To analyze the impact of such a memorization leak in detail, we characterize model behavior on symbolic addition analytically for training sets that are balanced around zero:
+We first consider an example case ($S(1;2)=0.4,\mathcal{W}=\{0\}$).
+We find that though the model perfectly learns the training cases, it underestimates the test cases by a proportional factor (Fig.
+3a).
+To understand why, we consider the model’s functional form on the training cases: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ (see Section 4).
+Intuitively, $\ell_{2}$-norm minimization tends to yield distributed weights.
+As a result, unless the salience of the conjunction $S(2;2)$ is zero, the kernel model will associate some non-zero weight with it (i.e.
+$f_{12}(z)\neq 0$).
+This, in turn, necessarily distorts the generalization on the test set, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
+We call this tendency to use the full conjunction (and its effect on generalization) “memorization leak.” To characterize it more systematically (and add mathematical rigor to the intuition above), we analytically characterize model behavior on a general family of symbolic addition tasks:
 Proposition 5.2.
-Consider input elements in $\{[v]\}_{v\in\mathcal{V}},\mathcal{V}\subset\mathbb{R}$ with associated values $v$.
-We assume that the training set contains all pairs such that at least one component is $z_{c}\in\{[w]\}_{w\in\mathcal{W}},\mathcal{W}\subset\mathbb{R}$ and that the average value in both $\mathcal{V}$ and $\mathcal{W}$ is zero.
+Consider inputs $z_{1},z_{2}\in\{[v]\}_{v\in\mathcal{V}},\mathcal{V}\subset\mathbb{R}$, with associated values $v$.
+We assume that the training set contains all pairs such that at least one component is $z_{c}\in\{[w]\}_{w\in\mathcal{W}},\mathcal{W}\subset\mathcal{V}$ and that the average value in both $\mathcal{V}$ and $\mathcal{W}$ is zero.
 Then, model behavior on the test set is given by
-$f([i],[j])=m(i+j),\quad m:=\tfrac{p\cdot\text{Sal}_{2}(1)}{1+(p-2)\text{Sal}_{2}(1)},\quad p:=|\mathcal{W}|$ (7)
-We prove the proposition in Section D.1.1.
-It implies that for any representation that encodes the conjunction (i.e.
-has $\text{Sal}_{2}(2)>0$ and therefore $\text{Sal}_{2}(1)<\tfrac{1}{2}$), the model underestimates the inferred values by a constant factor $m<1$.
-Further, $m$ decreases for smaller $\text{Sal}_{2}(1)$ and increases for larger training set size (Fig.
-3a,b).
-Perhaps surprisingly, these are the only two factors influencing $m$.
-In particular, while we often assume that interpolation is easier than extrapolation, these changes in training set do not impact model behavior here.
-We found qualitatively similar behavior for randomly generated, dispersed training sets (Section D.1.2, Fig.
-7).
+$f([v_{1}],[v_{2}])=m(v_{1}+v_{2}),\quad m:=\tfrac{p\cdot S(1;2)}{1+(p-2)S(1;2)},\quad p:=|\mathcal{W}|$ (5)
+We prove the proposition in Section D.1.
+It implies that for this entire task family, model generalization is distorted by a proportional factor $m$ (see Fig.
+3b).
+The formula implies that $m<1$ as long as $S(1;2)<\tfrac{1}{2}$ (i.e.
+as long as $S(2;2)>0$).
+Further, $m$ is smaller for lower $S(1;2)$ and smaller training set size $p$.
+(Note that $\mathcal{W}$ corresponds to the rows/columns making up the example training sets in Fig.
+1b.) Interestingly, these are the only two factors that influence $m$.
+In particular, while interpolation is often seen as easier than extrapolation, this does not impact model behavior here.
+More broadly, the memorization leak is a ubiquitous issue for statistical learning models trained on compositional tasks: $\ell_{2}$-norm minimization tends to yield distributed weights and so if the conjunctive population is represented, the model will generally end up partially relying on this conjunction.
+As shown above, this necessarily distorts generalization (see also Section D.3 for a minimal example).
 Figure 3: Kernel models’ behavior on (a,b) symbolic addition and (c,d) context dependence.
-a, Factor $m$ distorting the inferred values as a function of $\text{Sal}_{2}(1)$ and the training set size $p=|\mathcal{W}|$.
-b, Inferred values on extrapolation task.
-c, Variants of context dependence for which different representational saliences yield successful generalization.
+a, Model predictions on training and test set plotted against ground truth for an example case ($S(1;2)=0.4$, $\mathcal{W}=\{0\}$).
+b, The slope of the test set as a function of $S(1;2)$ and the training set size $p=|\mathcal{W}|$.
+c, Generalization on context dependence as a function of representational salience.
 Trajectories of networks with different nonlinearities are highlighted (color scale see Fig.
 2).
-d, Inferred coefficients of the different conjunction types for two example representations: a network with three layers and a rectified linear or quadratic nonlinearity.
-Notably, because the full conjunction can always be used to learn the training set, the memorization leak impacts a broad range of compositional tasks, including partial exposure (Section D.4).
+d, Coefficients of the different conjunction types for two example networks with three layers and different nonlinearities.
 Context dependence suffers from a shortcut bias.
@@
 We found that on a given task, each model either generalizes with 100% or 0% accuracy.
-For CD-3, the model only generalizes when $\text{Sal}_{3}(2)$ is high relative to $\text{Sal}_{3}(1)$ (Fig.
-3c).
+For CD-3 (the task version leaving out the largest subset of feature combinations), the model only generalizes when $S(2;3)$ is high relative to $S(1;3)$ (Fig.
+3c, dark green area).
 As a result, whether the network generalizes is highly sensitive to the nonlinearity and depth of the network.
 In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
-To better understand this, we analyzed the total magnitude of model weights associated with the different conjunctions.
-Notably, there are many possible conjunctions the model could rely on (see Section D.2.4).
-In particular, context $1$ is more likely to have the target $y=1$ and context $2$ is more likely to have target $y=-1$.
-The model may thus rely on this statistical shortcut, using the full conjunction to memorize the remaining training data.
-Indeed, we found that representations that generalized unsuccessfully on CD-3 consistently had a high weight associated with the context component and the full conjunction (Fig.
+To understand this phenomenon, we determined the total magnitude of model weights associated with the different conjunctions (Fig.
 3d).
-Notably, for CD-3, the context shortcut yields an accuracy of $\tfrac{2}{3}$.
-This explains why when $\text{Sal}_{3}(1)$ and $\text{Sal}_{3}(3)$ are high relative to $\text{Sal}_{3}(2)$, the model uses the context-driven shortcut together with the full conjunction, resulting in a failure to generalize.
-For CD-2, the context-driven shortcut only results in an accuracy of $\tfrac{9}{16}$, making it much less useful.
-This explains why only very low $\text{Sal}_{3}(2)$ yields failure to generalize on CD-2 or CD-1.
-Our analysis illustrates how conjunction-wise additivity lets us identify component conjunctions that yield a shortcut.
-The representational salience of these conjunctions and the extent to which they can explain the training set then helps us understand when this shortcut impacts generalization.
-4.4 Conjunction-wise additivity can describe generalization behavior in deep neural networks
-So far, we have analyzed kernel models, for their analytical tractability.
-We conjecture that large-scale neural network architectures with more complex, interrelated inputs might also be prone to implementing a conjunction-wise additive computation.
-In that case, our analyis of the kernel models would shed light on when the statistics of the training data enable or prevent compositional generalization in deep neural networks as well.
-To test our conjecture, we trained convolutional networks (ConvNets, [lecun_backpropagation_1989]) on a version of the tasks with concatenated MNIST digits [lecun_gradient-based_1998] and residual networks (ResNets, [he_deep_2016]) and Vision Transformers (ViTs, [dosovitskiy_image_2020]) on a version of the tasks with concatenated CIFAR-10 images [krizhevsky_learning_2009].
-We considered the same tasks except that each component was now given by images of a particular category, rather than being a single instance.
+We found that unsuccessful models (e.g.
+blue color in Fig.
+3d) had much larger magnitudes associated with the context component and the full conjunction.
+Notably, on CD-3, context is highly correlated with the target and can predict 2/3 of the training data.
+Models with high $S(1;3)$ (context) and $S(3;3)$ (full conjunction) exploit the context shortcut and use the full conjunction to learn the remaining training data.
+This strategy explains why these models fail to generalize to the test set.
+For CD-2, context is much less predictive on the training data (accuracy of $\tfrac{9}{16}$).
+This explains why only very low $S(2;3)$ yields failure to generalize on CD-2 or CD-1.
+Our analysis shows how model and task structure interact to either give rise to a generalizable rule or a statistical shortcut (see Section D.3 for a minimal example).
+This highlights that for compositional tasks with strong spurious correlations, model behavior will be highly sensitive to architectural details affecting the representational geometry such as depth and nonlinearity.
+6 Our theory describes the behavior of deep networks on conjunction-wise additive tasks
+Figure 4: Testing our theory in deep networks trained on MNIST and CIFAR versions of compositional tasks.
+Ranges indicate mean $\pm$ one std.
+error (often too small to be visible).
+a, $S(1;2)$ in an intermediate ConvNet layer for different distances between digits.
+Lower distance yields a more conjunctive representation.
+b, Average model prediction for each combination of components plotted against the ground truth (MNIST, distance of zero, $\mathcal{W}=\{0\}$).
+Generalization on the compositional split is distorted by a proportional factor.
+c, d, Slope of this linear relationship across all datasets as a function of c, distance (each line corresponding to a particular $\mathcal{W}$) and d, training set (for a distance of zero).
+For MSE instead of slope, see Fig.
+9.
+e, Accuracy on all variants of context dependence.
+So far, our analyses have been limited to simple kernel models, for their analytical tractability.
+We next test whether the insights developed from these simple models extend to a broader class of models: large-scale neural networks.
+We considered symbolic addition and context dependence on inputs created by concatenations of images from MNIST (Lecun et al., 1998) or CIFAR-10 (Krizhevsky et al., 2009).
+Each component corresponds to random samples from a particular category rather than the one-hot input considered so far and we study both in-distribution and compositional generalization.
 Notably, different image categories are not necessarily equally correlated with each other.
 To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment.
-To compare the network behaviors to our theoretical predictions, we fit a conjunction-wise additive function to the network output using linear regression (“additivity analysis”; see Section C.3).
-Symbolic addition.
-We trained the ConvNets on a total of 20,000 randomly generated MNIST samples for 100 epochs and trained the ResNets and ViTs on a total of 40,000 CIFAR-10 samples for 100 and 200 epochs, respectively.
-Remarkably, we found that the kernel theory matches network behavior almost perfectly: the networks’ average predictions (across the different possible permutations of the digit categories) were well predicted by an additive structure and their inferred values generally underestimated the ground truth (Fig.
-4a, Figs.
-9 and 10).
-This suggests that the kernel theory can capture the behavior of relevant neural network architectures trained on natural image inputs.
-Next, to investigate the effect of increased conjunctivity, we varied the distance between the two MNIST digits.
+We considered several relevant vision architectures: convolutional networks (ConvNets, LeCun et al., 1989) (trained on MNIST) and residual networks (ResNets, He et al., 2016) and Vision Transformers (ViTs, Dosovitskiy et al., 2020) (trained on CIFAR-10).
+Spatial distance of components impact salience in internal representations.
+First, we examined how changes in the input structure impact the networks’ representational geometry.
 We hypothesized that the ConvNets’ local weight structure should produce a more conjunctive representation for digits that are closer together.
-To test this, we determined $\text{Sal}_{2}(1)$ in an intermediate layer of the network, averaging over different instances of all digits.
-We found that $\text{Sal}_{2}(1)$ was indeed smaller for lower distances (Fig.
+To test this, we determined $S(1;2)$ in an intermediate layer of the network, averaging over different instances of all digits.
+We found that $S(1;2)$ was indeed smaller for lower distances (Fig.
+4a), indicating a more conjunctive population.
+This suggests that varying the distance between two digits provides a practical way of manipulating $S(1;2)$; below we use this insight to test predictions about how conjunctivity influences generalization behavior.
+Deep networks implement a conjunction-wise additive computation.
+Conjunction-wise additivity imposes a substantial computational restriction: on test set inputs, models can only add up values assigned to each conjunction of components seen during training.
+We therefore investigated whether the deep networks’ computations could be captured in these terms.
+Remarkably, we found that such a model was highly predictive of the model responses (see Section C.2).
+This indicates that large-scale neural networks tend to implement conjunction-wise additive computations — at least when trained on conjunction-wise additive tasks.
+Our theory suggests that they do so because the statistical structure of the dataset makes such a computation natural.
+Deep networks are impacted by a memorization leak on symbolic addition.
+Having found that our theory captures the networks’ general computational structure, we investigated whether it was also able to explain their specific performance.
+We first considered symbolic addition, training the ConvNets on seven different training sets with 20,000 samples and the ResNets and ViTs on three training sets with 40,000 samples.
+We varied both the size of $\mathcal{W}$ and whether generalization required interpolation or also extrapolation (Section D.1.3).
+We tested three theoretical predictions made by Proposition 5.2.
+First, our theory predicts that the model’s test set response should be distorted by a proportional factor.
+To test this, we plotted the average model prediction for each combination of categories against the ground truth.
+We found that model predictions were indeed compressed relative to the ground truth, but still exhibited a strong linear relationship (Fig.
 4b).
-Further, the inferred values were more compressed for smaller distances, confirming that more conjunctive inputs exacerbate the memorization leak in ConvNets as well (Fig.
-4c and Fig.
-9).
-Figure 4: Behavior of networks trained on MNIST and CIFAR versions of compositional tasks.
-Shaded regions indicate mean $\pm$ two standard errors.
-a, Inferred values on symbolic addition (extrapolation, distance of zero between digits).
-b, $\text{Sal}_{2}(1)$ in an intermediate layer of the ConvNet, for different distances between digits.
-c, Inferred values for the extrapolation task and different distances between digits.
-d, For the networks trained on variants of context dependence, the inferred coefficients for different conjunction types (distance of 20px).
-Context dependence.
-We then trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples.
-Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but failing to generalize on CD-3.
-The additivity analysis revealed that this was due to a context-driven shortcut (Fig.
+This was also the case for ResNets and ViTs, which exhibited a slightly noisier linear relationship (Fig.
+8).
+The memorization leak therefore affects generalization in large-scale networks as well.
+This is especially significant as our theoretical argument suggests that memorization leaks likely arise for a broad range of compositional tasks.
+We then estimated the slope of this linear dependency for each dataset.
+Our theory predicts that the compositional generalization slope should increase with increasing $S(1;2)$ (i.e.
+increasing distance between components, Fig.
+4a).
+On ConvNets, we found that higher distance indeed increases the slope (Fig.
+4c).
+This was not due to an increase in task difficulty, as in-distribution generalization is not systematically affected by distance.
+On ResNets and ViTs, higher distance also yielded a higher slope on the compositional generalization set (though the effect was much subtler for ViTs).
+(Notably, these networks did not perfectly predict the training split, which may affect the results.)
+Finally, our theory predicts that larger $\mathcal{W}$ should increase the compositional generalization slope.
+Our experiments confirmed this prediction (Fig.
 4d).
-4.5 Feature learning can overcome the limitations of kernel models
-Figure 5: Feature learning enables generalization on transitive equivalence.
-a, Learning curves in the lazy and rich regime.
-b, The weights of the network in the subspace corresponding to one underlying XOR-task.
-c, Weights for the same unit are plotted against each other and colored by whether they correspond to equivalent items (purple) or non-equivalent items (green).
-d, Accuracy of a ConvNet trained on an MNIST version of transitive equivalence.
-We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence.
-To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input.
-By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime.
-Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs.
-8 and 11).
-On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig.
-5a).
-To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm ([lyu_gradient_2020, chizat_implicit_2020]; cf.
-[vardi_implicit_2021]).
-In particular, a one-hidden-layer ReLU network tends to learn a sparse set of features [savarese_how_2019, chizat_implicit_2020].
-Transitive equivalence consists of multiple overlapping equality relations (e.g.
-$A=B\neq D=E$).
-Notably, ReLU networks such an XOR-type problem by specializing one unit to each conjunction ([brutzkus_why_2019, saxe_neural_2022, xu_benign_2023]; Fig.
-5b).
-Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g.
-$(A,B),(A,C)$, and $(B,C)$).
-This causes the unit to generalize to unseen item combinations (e.g.
-$(A,C)$), enabling the network to generalize.
-Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig.
-5c).
-Thus, rich networks’ capacity for abstraction gives rise to an additional compositional motif, allowing them to generalize on transitive equivalence.
-In particular, our findings highlight that transitive equivalence and transitive ordering are solved by fundamentally different network motifs.
-To see whether large-scale neural networks can also benefit from this feature-learning mechanism, we trained ConvNets on an MNIST version of transitive equivalence.
-The networks were trained for 150 epochs on 20,000 samples.
-We found that if the digits were presented with a distance of zero, the network did not generalize compositionally at all.
-However, with increasing distance, the network started to improve its compositional generalization (Fig.
-5e), demonstrating that a convolutional network can benefit from this rich compositional motif.
+Further, whether the training set required interpolation or also extrapolation did not systematically affect the resulting slope, as predicted by our theory (though interestingly, it did affect in-distribution generalization).
+However, there was one exception to this: for $p=4$, the extrapolation dataset had a smaller slope than the interpolation data set.
+Our theory can therefore explain much but not all of the deep network behavior.
+Deep networks are impacted by a shortcut bias on context dependence.
+Lastly, we trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples.
+Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but having worse-than-chance accuracy on CD-3 (Fig.
+4e).
+This confirms that the deep networks also suffered from a context-driven shortcut on CD-3.
+7 Outlook: feature learning enables non-additive compositional generalization
+In the previous section, we found that neural networks trained on symbolic addition or context dependence implement a conjunction-wise additive computation.
+However, conjunction-wise additivity is not the only mechanism by which deep neural networks generalize compositionally.
+To demonstrate this, we show in Appendix E that feature-learning networks can generalize on transitive equivalence (a non-additive task) due to their tendency to learn low-dimensional internal representations.
+This highlights that neural networks can implement non-additive compositional computations.
+Future work should explore more systematically how feature learning impacts compositional generalization.
 8 Discussion
-Humans often generalize to new situations by stitching together concepts and knowledge from prior experience.
-Despite the broad importance of this ability (both for cognitive science and machine learning), it has remained unclear under what conditions compositionally structured representations give rise to compositional generalization.
-Here we have taken a step toward formalizing this relationship by clarifying the full range of compositional motifs implemented by kernel models with categorical compositionally structured representations (“conjunction-wise additivity”).
-For conjunction-wise additive tasks, we analyzed how representational geometry and training data impact successful generalization on important compositional building blocks, validating our analysis for deep neural networks trained on natural image data.
-For non-additive tasks, our results immediately imply that kernel models are unable to generalize.
-We then analyzed how feature-learning mechanisms can overcome these limitations and demonstrated that deep networks benefit from the same mechanisms.
-Taken together, this suggests that conjunction-wise additivity provides a useful generalization class for understanding compositional generalization in neural networks and perhaps even humans.
-Limitations.
-We considered tasks with a fixed number of input components and fixed output dimensionality and assumed that all input components are represented orthogonally, excluding continuously varying components such as position.
-Further, our theoretical analysis focused on kernel models with fixed representations (though we considered deep neural networks through empirical simulations).
-Future work could extend the presented theory beyond these assumptions, and could further explore the range of compositional motifs in other learning models (e.g.
-rich neural networks and modular networks) more systematically.
-Nevertheless, we believe that this work is an important step towards a comprehensive theory of compositional generalization, covering a broad range of tasks and highlighting several phenomena that might help us understand more complex models as well.
-Acknowledgments
-[…]
-Appendix A Proof of Theorem 4.1
+Humans often generalize to new situations by stitching together concepts and knowledge from prior experience in new ways.
+Despite the broad importance of this ability (both for cognitive science and machine learning), a general theory of when and how neural networks accomplish compositional generalization has remained elusive.
+Here we have taken a step towards formalizing this relationship by characterizing kernel models with compositionally structured representations, a framework that captures neural network training in the kernel regime and more broadly lets us understand the impact of dataset statistics on generalization.
+We found that they implement a specific compositional computation (“conjunction-wise additivity”).
+We then investigated how dataset statistics and representational geometry impact successful generalization, highlighting two failure modes arising from the inductive bias of gradient descent (memorization leak and shortcut bias).
+Finally, we validated our theory in deep neural networks trained on natural image data.
+Our results show how simple statistical models can implement (or at least approximate) abstract rules like context dependence.
+However, they also highlight that building in compositional structure by itself is often insufficient for compositional generalization.
+Indeed, contextual generalization is highly sensitive to the specific representational geometry and training data.
+This may explain why investigations into compositional generalization and the benefits of disentangled representation have yielded such inconsistent results.
+Overall, our insights highlight the utility of kernel models in systematically investigating model generalization (or at least generating an initial hypothesis as to the important factors).
+We here used these models to characterize deep neural networks, but they could be equally useful for better understanding human compositional generalization.
+While our work covers a broad range of different tasks, a number of limitations remain.
+Though out of scope here, it would be interesting to test this theory in extremely large-scale models, e.g.
+large language models.
+Future work should also consider inputs and outputs with broader ranges of formats (e.g.
+sequences with variable length) as well as representations that are not compositionally structured.
+Most importantly, our theory is limited to a particular learning mechanism (fixed features with linear readout); other learning mechanisms could either implement a conjunction-wise additive model that overcomes the failure modes highlighted here or implement novel compositional computations beyond conjunction-wise additivity (we gave an initial example of this in Section 7).
+By clarifying the general relationship between dataset statistics and compositional generalization, our theory provides an important foundation for such advances.
+Acknowledgements
+We are grateful to the members of the Center for Theoretical Neuroscience for helpful comments and discussions.
+We thank Elom Amematsro, Ching Fang, and Matteo Alleman for detailed feedback.
+The work was supported by NSF 1707398 (Neuronex), Gatsby Charitable Foundation GAT3708, and the NSF AI Institute for Artificial and Natural Intelligence (ARNI).
+Appendix A Mathematical analysis
+A.1 Generalized compositionally structured representations
+In the main text, we considered “compositionally structured” representations.
+This is the class of representations $\phi(z)$ whose kernel $K_{\phi}(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ only depends on the number of components $z$ and $z^{\prime}$ have in common.
+Here we define a slightly more general class of representations whose kernel depends on the identity of the components they have in common, i.e.
+$O(z,z^{\prime}):=\{c=1,\dotsc,C|z_{c}=z_{c}^{\prime}\}.$ (6)
+Thus, pairs of representations overlapping in the first component may have a different similarity than pairs of representations overlapping in the second component, but any pair of representations overlapping in the first component still need to have the same similarity.
+This case captures, for example, cases where one feature is more salient than another and therefore influences the representational similarity more strongly, or where conjunctions between certain components are more saliently represented than conjunctions between other components.
+For example, perhaps our input consists of two objects with different shapes and colors.
+In that case, we may wish to represent the conjunction between shape and color of the first object more strongly than the conjunction between the shape of the first object and the color of the second.
+Note that we consider the constrained definition of compositionally structured representations in the main text purely for didactical purposes.
+In particular, our observations in Sections A.3 and A.4 extend to all generalized compositionally structured representations.
+A.2 Why are we considering compositionally structured representations?
+So far, we have motivated the concept of compositionally structured representations in terms of the fact that disentangled representations as well as many nonlinear transforms of disentangled representations are compositionally structured.
+Here we discuss a more conceptual motivation for them: compositionally structured representations guarantee that the only basis for generalization is the compositional nature of the data.
+This is because we ensure that the model has no knowledge about certain components that may be more similar to each other.
+Otherwise, the component-wise similarity could serve as a basis for generalization.
+For example, in transitive equivalence (Fig.
+1d), if items that are in the same equivalence class are represented as more similar to each other, that would serve as a basis for generalization on the task that is not compositional in nature.
+For this reason, we are particularly interested in compositionally structured representations, as they ensure that kernel models are only able to generalize compositionally.
+Characterizing compositional generalization in non-compositionally structured representations requires us to meaningfully distinguish between compositional and non-compositional aspects of their generalization.
+We consider this question out of scope for this manuscript, but believe it would be interesting future work.
+A.3 Nonlinear transformations and compositional structure
+A broad range of transforms $\phi$ induces a kernel $K_{\phi}(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ that only depends on the input similarity $K(z,z^{\prime})=z^{T}z^{\prime}$.
+In particular (as noted in the main text), this condition is satisfied by the hidden layers and neural tangent kernel of randomly initialized neural networks (in the infinite-width limit) (Han et al., 2022):
+Definition A.1.
+Given an input $z\in\mathbb{R}^{d}$, a network depth $L\geq 2$, and a set of widths $H_{1},\dotsc,H_{L}\in\mathbb{N}$, we define a neural network by recursively defining the operation $\phi^{(l)}$ of the $l^{\text{th}}$ layer as
+$a^{(0)}:=z,\quad a^{(l)}:=\phi^{(l)}(a^{(l-1)}):=\sigma\left(\tfrac{1}{\sqrt{H_{l}}}\left(W^{(l)}a^{(l-1)}+b^{(l)}\right)\right),\quad W^{(l)}\in\mathbb{R}^{H_{l}\times H_{l-1}}$ (7)
+where $W^{(l)}$ and $b^{(l)}$ are i.i.d.
+sampled from a random distribution and $\sigma:\mathbb{R}\to\mathbb{R}$ is a nonlinearity.
+The complete transform is then given by $\phi:=\phi^{(L)}\circ\dotsb\circ\phi^{(1)}$.
+The infinite-width limit is characterized by $H_{1},\dotsc,H_{L-1}\to\infty$ (the order of these limits does not matter).
+This setup includes, in particular, the random feature model investigated by Abbe et al.
+(2023).
+Further, the condition is met by most commonly used kernel functions including any radial basis function that depends on the Euclidean $\ell_{2}$-norm (e.g.
+the Gaussian kernel).
+Any nonlinear transform that meets this condition also conserves compositional structure, i.e.
+if $z$ is compositionally structured then so is $\phi(z)$.
+This allows us to derive computational restrictions on this broad range of models.
+A.4 Proof of Theorem 4.1
 See 4.1
 Proof.
-Because $K$ is compositionally structured, its similarity $K(I,I^{\prime})$ only depends on the overlap $O(I,I^{\prime})\subseteq\{1,\dotsc,C\}$.
-We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define the overlap in the training dataset as
-$\mathcal{D}_{S}(I):=\left\{I^{\prime}\in\mathcal{D}^{(\text{train})}|\forall_{c\in S}I_{c}=I^{\prime}_{c}\right\}.$ (8)
-The key idea is to decompose $\mathcal{D}^{(\text{train})}$ into these different overlaps in order to separate the sum into its components.
-However, by our definition, the datasets $\mathcal{D}_{S}(I)$ are not disjoint.
-Indeed, $S\subseteq S^{\prime}$ implies $\mathcal{D}_{S^{\prime}}(I)\subseteq\mathcal{D}_{S}(I)$ and in particular $\mathcal{D}_{\emptyset}(I)=\mathcal{D}^{(\text{train})}$.
-To adjust for this, we define $\delta_{S}$ as the similarity added by $\kappa_{S}$ to the similarity between conjunctions with one component fewer, recursively defining
-$\delta_{\emptyset}=\kappa_{\emptyset},\quad\delta_{S}=\kappa_{S}-\sum_{S^{\prime}\subsetneq S}\delta_{S^{\prime}}.$ (9)
+Because the kernel $K$ is compositionally structured, its similarity $K(z,z^{\prime})$ only depends on the overlap $O(z,z^{\prime})\subseteq\{1,\dotsc,C\}$ (see definition in Eq.
+6).
+We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define set of training items overlapping with $z\in Z^{\text{test}}$ in $S$ as
+$Z^{\text{train}}(z,J):=\left\{z^{\text{tr}}\in Z^{\text{train}}|\forall_{c\in J}z_{c}=z^{\text{tr}}_{c}\right\}.$ (8)
+The key idea is to decompose $Z^{\text{train}}$ into these different overlaps in order to separate the sum into its components.
+However, by our definition, the datasets $Z^{\text{train}}(z,J)$ are not disjoint.
+Indeed, $S\subseteq S^{\prime}$ implies $Z^{\text{train}}(z,J)\subseteq Z^{\text{train}}(z,J^{\prime})$ and in particular $Z^{\text{train}}(z,\emptyset)=Z^{\text{train}}$.
+To adjust for this, we define $\delta_{J}$ as the similarity added by $\kappa_{S}$ to the similarity between conjunctions with one component fewer, recursively defining
+$\delta_{\emptyset}=\kappa_{\emptyset},\quad\delta_{J}=\kappa_{J}-\sum_{J^{\prime}\subsetneq J}\delta_{J^{\prime}}.$ (9)
 We then decompose
-$f(I)=\sum_{I^{\prime}\in\mathcal{D}^{(\text{train})}}a_{I^{\prime}}K(I,I^{\prime})=\sum_{S\subseteq\{1,\dotsc,C\}}\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}}.$ (10)
-This equality obtains because for each $I^{\prime}\in\mathcal{D}^{(\text{train})}$,
-$\sum_{S:I^{\prime}\in\mathcal{D}_{S}(I)}\delta_{S}=\delta_{O(I,I^{\prime})}+\sum_{S\subsetneq O(I,I^{\prime})}\delta_{S^{\prime}}=\kappa_{O(I,I^{\prime})}=K(I,I^{\prime}),$ (11)
-which is true by definition. We note that for $S\notin\mathcal{O}(I)$, $\mathcal{D}_{S}(I)=\emptyset$. Defining
-$f_{S}(I):=\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}},$ (12)
+$f(z)=\sum_{z^{\text{tr}}\in Z^{\text{train}}}a_{z^{\text{tr}}}K(z,z^{\text{tr}})=\sum_{J\subseteq\{1,\dotsc,C\}}\delta_{S}\sum_{z^{\text{tr}}\in Z^{\text{train}}(z,J)}a_{z^{\text{tr}}}.$ (10)
+This equality obtains because for each $z\in Z,z^{\text{tr}}\in Z^{\text{train}}$,
+$\sum_{J:z^{\text{tr}}\in Z^{\text{train}}(z,J)}\delta_{J}=\delta_{O(z,z^{\text{tr}})}+\sum_{J^{\prime}\subsetneq O(z,z^{\text{tr}})}\delta_{J^{\prime}}=\kappa_{O(z,z^{\text{tr}})}=K(z,z^{\text{tr}}),$ (11)
+which is true by definition. We note that for $J\notin\text{Conj}(z|Z^{\text{train}})$, $Z^{\text{train}}(z,J)=\emptyset$. Defining
+$f_{J}(z):=\delta_{J}\sum_{z^{\text{tr}}\in Z^{\text{train}}(z,J)}a_{z^{\text{tr}}},$ (12)
 proves the proposition. ∎
-Appendix B Representational geometries across different deep neural network architectures
-We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022].
-B.1 Proof of Proposition 4.3
-See 4.3
+Appendix B Representation salience: a metric for compositionally structured representations
+B.1 Definition
+Below we formally define representational salience.
+To do so, we denote the similarity between two trials $z,z^{\prime}$ by $\text{Sim}(O(z,z^{\prime})):=K(z,z^{\prime})$.
+Note that this is well-defined for compositionally structured representations, as $K(z,z^{\prime})$ is identical for all pairs of trials with the same $O(z,z^{\prime})$.
+We then define:
+Definition B.1.
+For a generalized compositionally structured representation with kernel $K$ and a conjunction $J\subseteq\{1,\dotsc,C\}$, we recursively define
+$\overline{S}(\emptyset):=K(\emptyset),\quad\overline{S}(J):=\text{Sim}(J)-\sum_{J^{\prime}\subsetneq J}\overline{S}(J^{\prime}),\quad S(J)=\tfrac{\overline{S}(J)}{\sum_{\emptyset\neq J\subseteq\{1,\dotsc,C\}}\overline{S}(J)}.$ (13)
+When emphasizing the total number of components, we denote $S(J;C)=S(J)$.
+Further, for compositionally structured representation, the similarity only depends on the total number of overlaps $|J|$.
+We therefore write $S(c;C)=S(J)$, where $c:=|J|$.
+B.2 Why is representational salience a useful metric?
+Intuitively, the representational salience captures the unique contribution of a population representing a particular conjunction $J$ (as in our informal definition in the main text).
+As we noted in the main text, distinguishing these unique contributions from the similarities directly would be more difficult.
+For example, changes in the similarity between inputs having two components in common could arise from changes in how saliently single components or conjunctions of two components are represented.
+To distinguish between these cases, we’d have to additionally look at the similarity between inputs having a single component in common.
+Our definition of representational salience avoids this issue; we therefore believe that this perspective could more broadly be useful for analyzing representations of compositional data.
+B.3 Why does representational salience leave out magnitude and baseline activity?
+Notably, many learning models are largely invariant to a constant rescaling of their representation.
+For limit behavior (e.g.
+in the limit of infinite training in gradient descent), this scale has no impact.
+We see this, for instance, in Proposition 5.2, where the scaling turns out to be entirely irrelevant.
+For regularized models, on the other hand, the scale of the representation is confounded with the strength of the regularization and so we suggest that it is again best seen as separate from the representational geometry.
+The baseline activity, on the other hand, may determine how easily an intercept is learned.
+However, many linear readout models (e.g.
+support vector machines) do not regularize their intercept, again rendering this parameter entirely irrelevant.
+Notably, in the case of gradient descent, the magnitude of the baseline activity does influence how easily an intercept is learned.
+However, the impact of this is often negligible; for example, in Proposition 5.2, we again find that the baseline activity is entirely irrelevant.
+B.4 Extended representational analysis
+Figure 5: Extended analysis of overlap salience in random neural networks.
+a, Salience of the first component (1), the second component (2), and their conjunction (12), where we vary the two components’ saliences in the input.
+b, Salience for inputs with four components.
+We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022).
+In Fig.
+5b, we plot the different saliences for an input with four components.
+Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig.
+2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions.
+Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two.
+This highlights that for larger numbers of components, the dependence of generalization behavior on specific architectural choices will likely be even stronger.
+Finally, we consider an example of a generalized compositionally structured representation.
+Here we assume that the different input components have different magnitudes.
+In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig.
+5a).
+As the random weight neural network becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation.
+While the salience of the full conjunction still eventually converges to one for most nonlinearities, it takes longer to do so for a less balanced input representation.
+Further, for a Gaussian nonlinearity, an imbalanced representation actually decreases the limit salience the representation appears to be converging to for the full conjunction.
+This highlights that for disentangled representations that do not represent all components with equal magnitude, compositional generalization behavior may vary strongly with different neural network architectures.
+B.5 Proof of Proposition 5.1
+See 5.1
 Proof.
-Note that the proof is a minor extension of Lemma S1.3 in [lippl_mathematical_2024]. We present it here in a self-contained manner. We consider a nonlinearity
+Note that the proof is a minor extension of Lemma S1.3 in Lippl et al. (2024). We present it here in a self-contained manner. We consider a nonlinearity
 $\sigma(u):=A\min(u,0)+\max(u,0),\quad A\in[0,1).$ (14)
-By prior work, [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022],
+By prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022),
 $\mathbb{E}_{w}\left[\phi({w}^{T}{h}^{(l)}(z))\phi({w}^{T}{h}^{(l)}(z^{\prime}))\right]=\sigma^{2}\|h^{(l)}(z)\|_{2}\|h^{(l)}(z^{\prime})\|_{2}k\left(\hat{h}^{(l)}(z)^{T}\hat{h}^{(l)}(z^{\prime})\right),$ (15)
@@
 where we infer the latter inequality by visual inspection of the plot of this function. ∎
-B.2 Extended analysis of representational salience
-Figure 6: Extended analysis of overlap salience in random neural networks.
-a, Salience of the first component (1), the second component (2), and their conjunction (12), where we vary the two components’ saliences in the input.
-b, Salience for inputs with four components.
-To complement Section 4.2, we analyze the impact of different magnitudes for different components.
-In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig.
-6a).
-As becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation.
-While the salience of the full conjunction still eventually converges to one for most nonlinearities, it takes longer to do so for a less balanced input representation.
-Further, for a Gaussian nonlinearity, an imbalanced representation actually decreases the limit salience the representation appears to be converging to for the full conjunction.
-In Fig.
-6, we plot the different saliences for an input with four components.
-Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig.
-2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions.
-Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two.
 Appendix C Detailed methods
+The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization.
 C.1 Models
 Kernel model.
-We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn [pedregosa_scikit-learn_2011].
+We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn Pedregosa et al.
+(2011).
+Note that this is equivalent to using a feature basis with the same resulting similarities.
+However, hand-specifying these similarities enabled us to easily explore the full range of possible representations.
 Rich and lazy ReLU networks.
-All networks were trained with Pytorch and Pytorch Lightning [paszke_pytorch_2019].
+All networks were trained with Pytorch and Pytorch Lightning Paszke et al.
+(2019).
 We consider ReLU networks with one hidden layer and $H=1000$ units.
@@
 Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation.
-All weights are initialized with He initialization [he_delving_2015].
+All weights are initialized with He initialization (He et al., 2015).
 Residual neural networks.
@@
 We did not use data augmentation for MNIST. For CIFAR-10, we used a random flip and a random crop.
-C.2 Reproducibility
-Computational simulations.
-We ran all computational simulations either on a CPU or on a single GPU.
-Each kernel experiment (total number: 53 experiments) ran within fifteen minutes, each rich network (total number: 450 experiments) and ConvNet experiment (total number: 600 experiments) took between one and two hours, and each ViT or ResNet experiment (total number: 140 experiments) took between five and ten hours.
-We used a computational cluster to run many of the experiments in parallel.
-As a rough estimate, the total experiments therefore required 900 hours of CPU computations and 2,600 hours of GPU computations.
-Code repository.
-The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization.
-We additionally uploaded the repository including all simulated data to Zenodo (https://doi.org/10.5281/zenodo.11308156).
-C.3 Additivity analysis
+C.2 Additivity analysis
 To analyze how well a conjunction-wise additive computation can describe network behavior, we considered as the set of possible features a concatenation of one-hot vectors coding for each possible conjunction.
@@
 $R^{2}=1$ indicates full conjunction-wise additivity).
-Furthermore, we can use the inferred values assigned to these different conjunctions to compare kernel models, rich and lazy networks, and convolutional networks.
-Note that for the convolutional networks, we first average the model predictions across all different images instantiating a given compositional input.
-Appendix D Compositional tasks
+Furthermore, we can use the inferred values assigned to these different conjunctions to compare kernel models, feature-learning networks, and vision networks.
+Note that for the vision networks, we first average the model predictions across all different images instantiating a given compositional input.
+Notably, our theory predicts that the constrained conjunction-wise additive function can predict behavior on the test set, but not necessarily on the training set.
+To test this prediction in vision network, we compared the additivity of the test set alone to the additivity of the entire dataset (Fig.
+6).
+First, we found that the predictivity on the test set was indeed generally very high.
+Second, we found that it was substantially higher than on the training and test set taken together.
+Figure 6: $R^{2}$ of the conjunction-wise additive predictor on the test set as well as on both the training and test set.
+Appendix D Compositional task space
 D.1 Symbolic addition
 D.1.1 Proof of Proposition 5.2
-See 4.4
+See 5.2
 Proof.
@@
 Let $p:=|\mathcal{W}|$ and $q:=|\mathcal{V}|-|\mathcal{W}|$.
+We denote by $\kappa_{c}$ the similarity between inputs overlapping in $c$ components.
 Then, setting $\delta_{2}:=\kappa_{2}-\kappa_{0}$ and $\delta_{1}:=\kappa_{1}-\kappa_{0}$, the set of dual equations is given by
@@
 Finally, $\overline{w}=\overline{v}=0$ immediately implies that $b=c=0$ and therefore $d=0$. ∎
-D.1.2 Behavior of kernel models on other tasks
-Figure 7: Kernel models on dispersed training sets.
-a, Four example training sets containing two random instances containing each item.
-b, Generalization loss as a function of the representation’s overlap salience.
-c,d, Inferred values for c) four example tasks and d) the average across fifty randomly drawn training sets.
-To test whether the observed behavior was specific to the kinds of training sets investigated in Proposition 4.4 and the main text, we additionally considered dispersed, randomly generated training sets with two randomly drawn trials containing each item (Fig.
-7a).
-We found that the generalized loss (averaged across all cases) increased roughly linearly with salience (Fig.
-7b).
-For individual training sets, the inferred values were distorted in a complex, irregular manner (Fig.
-7c).
-Further, because the training sets were no longer symmetric, the items took on different values when presented as the first or second component.
-On average, however, the inferred values were roughly proportional to their actual value and the factor of proportionality still decreased with salience (Fig.
-7d).
-D.1.3 Behavior of rich networks
-Figure 8: Behavior of rich networks on symbolic addition.
-a, Additivity as a function of initialization scale in networks with one hidden layer.
-b, Similarity between distinct, identical, and overlapping trials, before and after training in the lazy or rich regime.
-c, Inferred values in the lazy and rich regime for networks with one hidden layer.
-d, Zoomed-in inferred values for interpolation part of the task involving interpolation and extrapolation.
-e, Additivity as a function of the number of hidden layers.
-f, Inferred values for networks with different numbers of layers.
-We first trained rich networks with one hidden layer on the different variants of symbolic addition.
-We found that they generally were highly additive ($R^{2}>0.992$, Fig.
-8a).
-Fig.
-8b confirms that the ReLU networks indeed change their representation in the rich regime: at initialization, the similarity between different trials is approximately clustered by whether those trials are distinct, overlapping, or identical.
-After lazy training, this remains unchanged, whereas after rich training, the similarities look entirely different.
-Finally, the inferred values look similar on extrapolation, but do not exhibit a memorization leak on interpolation (Fig.
-8c).
-On the task involving both interpolation and extrapolation, the networks suffer from a memorization leak on the extrapolaton region but not the interpolation region (Fig.
-8d).
-This indicates that rich networks tend to perform better at interpolation, unlike lazy networks which are not sensitive to this difference.
-D.1.4 Behavior of deep networks
-Next, we trained networks with one to four hidden layers, using $\sigma=0.1$.
-We found that the networks were generally highly additive ($R^{2}>0.985$, Fig.
-8e) and inferred highly similar values.
-In particular, deeper networks also suffered from a memorization leak on extrapolation but not interpolation.
-D.1.5 Behavior of vision models trained on MNIST and CIFAR-10
-We now discuss in more detail the behavior of the different vision models.
-First, we trained ConvNets on an MNIST version of symbolic addition for the different training sets.
-For all tasks, we found that compositional generalization was worse than in-distribution generalization (Fig.
-9a) and become worse with smaller distance between the digits (i.e.
-higher conjunctivity).
-Further all networks were highly additive ($R^{2}>0.98$).
-Finally, the values inferred by the network were well consistent with the kernel theory: First, they were an affine function of the true item value for all training sets.
-Second, extrapolation was worse than interpolation and the task involving extrapolation and interpolation.
-Notably, interpolation and the task involving both yielded equally good values indicating that these networks may also not behave worse just based on extrapolation vs.
-interpolation.
-Further, smaller distance between digits (i.e.
-more conjunctive inputs) generally exacerbated the value misestimation.
-Finally, we also considered an asymmetric extrapolation task (“Asymmetric”), which trained on all trials containing an item $[-4]$.
-Consistent with our theory, we found that this resulted in an affine distortion in values.
-Figure 9: Behavior of ConvNets trained on MNIST versions of symbolic addition.
-a, Mean squared error for different training sets and different evaluation sets.
-b, Additivity on the different tasks.
-c, Inferred values.
-We then trained a ResNet and a ViT on a CIFAR-10 version of the task.
-We found that the error on the compositional test dataset was consistently higher than on in-distribution generalization (Fig.
-10a).
-Further, similarly to the ConvNets trained on MNIST, it was particularly high on the asymmetric training set.
-Notably, the additivity of all of these models was reasonably high ($R^{2}>0.93$, Fig.
-10b).
-Finally, we found that the models all inferred values that were distorted in a roughly affine manner (Fig.
-10c).
-Figure 10: Behavior of ResNets and ViTs on CIFAR versions of symbolic addition. a, Mean squared error over training. b, Additivity. c, Inferred values for the different tasks.
+D.1.2 Behavior of deep ReLU networks
+Next, we examined the behavior of deep ReLU networks with varying depth (Fig.
+7).
+On the extrapolation task ($\mathcal{W}=\{0\}$), these networks also had compressed model predictions on the compositional split that were approximately linearly distorted (though for deeper networks, this relationship seemed to have a slight S-shape).
+Intriguingly, these networks perfectly generalized on the interpolation task ($\mathcal{W}=\{-4,4\}$).
+For $\mathcal{W}=\{-2,2\}$, they perfectly generalized on the interpolation portion, but not the extrapolation portion.
+Figure 7: Model prediction plotted against ground truth for different training sets and depths.
+D.1.3 Behavior of vision models
+We trained the Convnets on MNIST using seven different training sets $\mathcal{W}$: $\{[0]\}$, $\{[-4],[4]\}$, $\{[-2],[2]\}$, $\{[-4],[0],[4]\}$, $\{[-1],[0],[1]\}$, $\{[-4],[-3],[3],[4]\}$, $\{[-2],[-1],[1],[2]\}$.
+This allowed us to vary both the size of the training set and whether it involved only interpolation or also extrapolation.
+We trained these networks for 100 epochs on 20,000 samples.
+We trained the ResNets and ViTs on CIFAR-10 using the first three training sets listed above.
+We trained the ResNets for 100 epochs and the ViTs for 200 epochs, training all networks on 40,000 samples.
+Our findings on these experiments are summarized in Section 6.
+Further, Fig.
+8 depicts, for each training set, the average prediction for each combination of components.
+We can see that the ConvNet predictions are highly linearly correlated with the ground truth.
+The ResNet and ViT predictions are also linearly correlated, but exhibit a slightly noisier relationship.
+Figure 8: Average prediction for each combination of components plotted against the ground truth for a, ConvNets trained on MNIST, b, ResNets trained on CIFAR-10, and c, ViTs trained on CIFAR-10.
+Additionally, we also plot the mean squared error of the networks on the different training sets (Fig.
+9).
+This provides a more conventional (but harder to interpret) measure of performance compared to the slope depicted in Fig.
+4.
+We again see that the mean squared error on the compositional generalization set is decreasing with increasing distance (Fig.
+9a).
+Further, it is decreasing with increasing training set size (Fig.
+9), though we note that the differences in the scales of the generalization set for different sets $\mathcal{W}$ renders it harder to directly to compare these values.
+In contrast, the slopes described in the main text are more easily comparable.
+Figure 9: a, Mean squared error as a function of distance. b, Mean squared error as a function of training set size $p$ (different dots corresponding to interpolation versus extrapolation.
 D.2 Context dependence
@@
 We then compute the average absolute magnitude within each of these groups in order to determine their overall relevance to model behavior.
-D.2.5 Rich networks
-We find that rich networks generalize consistently on CD-1 and CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 11).
+D.2.5 Feature-learning ReLU networks
+We find that feature-learning ReLU networks generalize consistently on CD-1 anREGULARIZATIONd CD-2 but not CD-3.
+They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig.
+10).
 Figure 10: Inferred values for different conjunctive groups on context dependence.
-D.2.6 Deep neural networks trained on vision data
+D.2.6 Vision networks
+We additionally analyzed ConvNets for different distances between the digits.
 We found that convolutional neural networks trained on MNIST successfully generalized on CD-1 and CD-2, but not CD-3 (Fig.
-12).
+11).
 For smaller distances between digits, the models tended to generalize worse on CD-1 and CD-2 and gradually reverted to chance accuracy (i.e.
@@
 Further, the networks were generally highly additive ($R^{2}>0.95$), but became worse for lower distance (Fig.
-12b).
+11b).
 Finally, across all distances, they had a high magnitude associated with the context cue, though this magnitude decreased for small distances — consistent with the accuracy of the network increasing from below chance to chance level (Fig.
-12c).
+11c).
 Figure 11: Performance of convolutional neural networks trained on an MNIST version of context dependence.
 For different distances and training sets, we plot a, the accuracy on different splits, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
-Finally, we considered ResNets and ViTs trained on a CIFAR version of the task.
-We found that they generally performed above chance for CD-2 and CD-1, but below chance for CD-3 (Fig.
-13a).
-Further, their additivity was generally high for ResNets ($R^{2}>0.95$) and slightly lower for ViTs ($R^{2}>0.9$) (Fig.
-13b) and both networks had a large magnitude associated with the context cue for CD-3, but not CD-2 or CD-1 (Fig.
-13c).
-Figure 13: Performance of ResNets and ViTs trained on a CIFAR version of context dependence.
-We plot a, their accuracy on different splits over training, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
-D.3 Transitive equivalence
-We additionally trained ReLU networks of various depth on the transitive equivalence (using $\sigma=0.1$ as initialization magnitude).
-We consistently found that they were able to generalize to the test set (Fig.
-14).
-Figure 14: Performance of networks with 100 hidden units per layer and varying depth on transitive equivalence. We used an initialization constant $\sigma=0.1$.
-D.4 Invariance and partial exposure
-We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut distortion.
+D.3 Invariance and partial exposure
+We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut bias.
 In both cases, the input consists of two components and the mapping only depends on the first (Fig.
-15a).
+12a).
 In the invariance case, we don’t see the second component vary at all, in the partial exposure case, we see one instance of the second component.
-Note that the partial exposure task has previously been studied in the context of network generalization [dasgupta_distinguishing_2022, chan_transformers_2022].
+Note that the partial exposure task has previously been studied in the context of network generalization (Dasgupta et al., 2022; Chan et al., 2022).
 To understand the impact of different representational saliences, we consider the generalization margin $m:=y\hat{y}$ on the test set, where $y$ is the ground-truth label and $\hat{y}$ is the model’s estimate.
 Because we consider support vector machines, the margin on the training set is one; a smaller margin on the test set indicates worse performance.
-We determined a mathematical formula for the margins as a function of $\text{Sal}_{2}(1)$.
+We determined a mathematical formula for the margins as a function of $S(1;2)$.
 Below we first describe its implications and then how we derived this.
 Invariance suffers from a memorization leak.
-On invariance, we find that the model’s test margin is expressed as $m=\tfrac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}$.
-This means that for a fully compositional representation ($\hat{\kappa}_{1}=0.5$), its training and test margins are both one.
-However, as $\hat{\kappa}_{1}$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig.
-15b).
+On invariance, we find that the model’s test margin is expressed as $m=\tfrac{S(1;2)}{1-S(1;2)}$.
+This means that for a fully compositional representation ($S(1;2)=0.5$), its training and test margins are both one.
+However, as $S(1;2)$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig.
+12b).
 Shortcut distortion.
@@
 This is a statistical shortcut and we find that norm minimization (just as for context dependence) ends up partially relying on this strategy as this decreases the $\ell_{2}$-norm of the readout weights.
-As a result, the test margin for the partial exposure task decreases even more strongly as a function of $\text{Sal}_{2}(1)$: $m=\tfrac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig.
-15b).
+As a result, the test margin for the partial exposure task decreases even more strongly as a function of $S(1;2)$: $m=\tfrac{2S(1;2)^{2}}{1-2S(1;2)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig.
+12b).
 Figure 12: a, Schema for invariance and partial exposure. b, Generalization margin for the two tasks.
@@
 As the ground truth labels are $y=\{1,-1\}$, the margin $m=y\hat{y}$ is identical for both test set points:
-$m=\frac{(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})}{\kappa_{2}^{2}-\kappa_{1}^{2}}=\frac{\kappa_{1}-\kappa_{0}}{\kappa_{2}-\kappa_{1}}=\frac{(\kappa_{2}-\kappa_{0})\text{Sal}_{2}(1)}{(\kappa_{2}-\kappa_{0})-(\kappa_{1}-\kappa_{0})}=\frac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}.$ (50)
+$m=\frac{(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})}{\kappa_{2}^{2}-\kappa_{1}^{2}}=\frac{\kappa_{1}-\kappa_{0}}{\kappa_{2}-\kappa_{1}}=\frac{(\kappa_{2}-\kappa_{0})S(1;2)}{(\kappa_{2}-\kappa_{0})-(\kappa_{1}-\kappa_{0})}=\frac{S(1;2)}{1-S(1;2)}.$ (51)
 For partial exposure, the training set is given by $\{(-1,-1),(-1,1),(1,-1)\}$ and its kernel is therefore
@@
 We solve this equation for the special case where $\kappa_{0}=0$ and $\kappa_{1}=1$ using Mathematica and find that
-$m=\frac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}.$ (54)
-D.5 Other mathematical operations
-We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}[I_{1}]$ and $v_{2}[I_{2}]$ together with some composition function $C(v_{1}[I_{1}],v_{2}[I_{2}])$.
+$m=\frac{2S(1;2)^{2}}{1-2S(1;2)^{2}}.$ (55)
+D.4 Other mathematical operations
+We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$ together with some composition function $C(v_{1}(z_{1}),v_{2}(z_{2}))$.
 This task will only be additive if the composition function is additive (e.g.
@@
 multiplication, division, or exponentiation, the task will be non-additive.
-D.6 Logical operations
+D.5 Logical operations
 […]
-D.7 Transitive ordering
-Transitive ordering is a popular task in cognitive science (often called transitive inference, [mcgonigle_are_1977]).
-Here the subject is presented with two items $I_{1}$, $I_{2}$ drawn from an unobserved hierarchy $>$.
-It should then categorize whether $I_{1}>I_{2}$ or $I_{2}>I_{1}$.
-Crucially, this task can be solved by assigning a rank $r(I_{c})$ to each item and computing the response as $f(I)=r(I_{1})-r(I_{2})$ [lippl_mathematical_2024].
+D.6 Transitive ordering
+Transitive ordering is a popular task in cognitive science (often called transitive inference, (McGonigle & Chalmers, 1977)).
+Here the subject is presented with two items $z_{1}$, $z_{2}$ drawn from an unobserved hierarchy $>$.
+It should then categorize whether $z_{1}>z_{2}$ or $z_{2}>z_{1}$.
+Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ (Lippl et al., 2024).
 It is therefore additive, in contrast to transitive equivalence.
@@
 In this case, the model would generalize to comparisons between these different hierarchies as well.
+Appendix E A feature-learning mechanism for non-additive compositional generalization
+Figure 13: Feature learning enables generalization on transitive equivalence.
+a, Learning curves in the lazy and rich regime.
+b, The weights of the network in the subspace corresponding to one underlying XOR-task.
+c, Weights for the same unit are plotted against each other and colored by whether they correspond to equivalent items (purple) or non-equivalent items (green).
+d, Accuracy of a ConvNet trained on an MNIST version of transitive equivalence.
+We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence.
+To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input.
+By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime.
+Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs.
+7 and 10).
+On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig.
+13a).
+To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm (Lyu & Li (2020); Chizat & Bach (2020); cf.
+Vardi & Shamir (2021)).
+In particular, a one-hidden-layer ReLU network tends to learn a sparse set of features Savarese et al.
+(2019); Chizat & Bach (2020).
+Transitive equivalence consists of multiple overlapping equality relations (e.g.
+$A=B\neq D=E$).
+Notably, ReLU networks such an XOR-type problem by specializing one unit to each conjunction (Brutzkus & Globerson (2019); Saxe et al.
+(2022); Xu et al.
+(2023); Fig.
+13b).
+Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g.
+$(A,B),(A,C)$, and $(B,C)$).
+This causes the unit to generalize to unseen item combinations (e.g.
+$(A,C)$), enabling the network to generalize.
+Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig.
+13c).
+Thus, rich networks’ capacity for abstraction gives rise to an additional compositional motif, allowing them to generalize on transitive equivalence.
+In particular, our findings highlight that transitive equivalence and transitive ordering are solved by fundamentally different network motifs.
+To see whether large-scale neural networks can also benefit from this feature-learning mechanism, we trained ConvNets on an MNIST version of transitive equivalence.
+The networks were trained for 150 epochs on 20,000 samples.
+We found that if the digits were presented with a distance of zero, the network did not generalize compositionally at all.
+However, with increasing distance, the network started to improve its compositional generalization (Fig.
+13d), demonstrating that a convolutional network can benefit from this rich compositional motif.

### v3

@@
-
+Samuel Lippl Center for Theoretical Neuroscience
 Abstract
@@
 Compositionally structured (e.g.
-disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear.
-To address this gap, we present a general theory of compositional generalization in kernel models with fixed representations, a tractable framework for characterizing the impact of dataset statistics on generalization.
-We find that kernel models are constrained to adding up values assigned to each combination of components seen during training (“conjunction-wise additivity”).
-This imposes fundamental restrictions on the set of tasks these models can learn, in particular preventing them from transitively generalizing equivalence relations.
-Even for compositional tasks that kernel models can in principle learn, we identify novel failure modes in compositional generalization that arise from biases in the training data and affect important compositional building blocks such as symbolic addition and context dependence (memorization leak and shortcut bias).
+disentangled) representations support this ability; however, the conditions under which they are sufficient for the emergence of compositional generalization remain unclear.
+To address this gap, we present a theory of compositional generalization in kernel models with fixed, compositionally structured representations.
+This provides a tractable framework for characterizing the impact of training data statistics on generalization.
+We find that these models are limited to functions that assign values to each combination of components seen during training, and then sum up these values (“conjunction-wise additivity”).
+This imposes fundamental restrictions on the set of tasks compositionally structured kernel models can learn, in particular preventing them from transitively generalizing equivalence relations.
+Even for compositional tasks that they can learn in principle, we identify novel failure modes in compositional generalization (memorization leak and shortcut bias) that arise from biases in the training data.
 Finally, we empirically validate our theory, showing that it captures the behavior of deep neural networks (convolutional networks, residual networks, and Vision Transformers) trained on a set of compositional tasks with similarly structured data.
-Ultimately, this work provides a theoretical perspective on how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
+Ultimately, this work examines how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
 1 Introduction
 Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant.
-Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Lake et al., 2017; Frankland & Greene, 2020).
-Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018).
+Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Fig.
+1a) (Lake et al., 2017; Frankland & Greene, 2020).
+Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018; Lepori et al., 2023).
 While a wide range of studies have investigated the conditions under which compositionally structured (i.e.
@@
 Some work suggests that disentangled representations improve compositional generalization (Esmaeili et al., 2019; van Steenkiste et al., 2019).
-However, others challenge this view (Locatello et al., 2019; Schott et al., 2022).
-In general, it remains unclear why insights from some compositional tasks fail to generalize to others, and a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
-To make progress on this question, we focus on standard statistical learning, a fundamental basis for compositional generalization that affects almost any machine learning model.
+Other studies challenge this view (Locatello et al., 2019; Schott et al., 2022).
+In general, a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
+Figure 1: Overview of main theoretical findings (Section 4).
+a, We consider inputs with several categorical components and study compositional generalization to novel combinations of components.
+b, We assume compositionally structured representations for which trials with the same number of overlaps have identical similarities.
+c, We find that a random weights neural network conserves compositional structure: if its input is compositionally structured then so is its output.
+d, We find that compositionally structured kernel models are constrained to adding up values for each conjunction seen during training (“conjunction-wise additivity”).
+To make progress on this question, we focus on standard statistical learning, a fundamental basis for generalization that affects almost any machine learning model.
 Generalization in statistical learning depends on the similarities between different inputs.
-Notably, in compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square).
-As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the generalization behavior of the model.
+In compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square).
+As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the model’s generalization behavior.
 To understand when and how statistical learning leads to successful compositional behavior, we developed a theory for the behavior of kernel models on compositional tasks.
-Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008, 2009).
+Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008; 2009).
 For example, kernel models accurately describe the behavior of learning and fine-tuning in neural networks under certain conditions (Jacot et al., 2018; Malladi et al., 2023).
-More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a, b).
+More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a; b).
 Despite their broad relevance and relative simplicity, it has remained unclear how even kernel models generalize on compositional tasks (though see Abbe et al., 2023; Lippl et al., 2024, see Section 2).
 To address this gap, we present a theory of compositional generalization in kernel models.
-We define a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory), theoretically characterize the compositional behavior of kernel models trained on such tasks, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures.
+We define a general class of “compositionally structured representations” (Fig.
+1b) and a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory).
+We then theoretically characterize the compositional behavior of compositionally structured kernel models, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures.
 Our specific contributions are as follows:
-  • In Section 4, we show that kernel models with disentangled inputs are constrained to summing up values implicitly assigned to each component or combination of components seen during training.
+  • In Section 4, we show that compositionally structured kernel models are constrained to summing up values implicitly assigned to each component or combination of components seen during training (Fig.
+1).
 We call such computations “conjunction-wise additive.”
   • In Section 5, we then characterize how representational geometry determines whether kernel models will generalize successfully on conjunction-wise additive compositional tasks, highlighting two important failure modes (memorization leak and shortcut bias).
-This highlights that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
+This demonstrates that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
   • Finally, in Section 6, we validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks.
-Overall, we take a step towards a general theory of compositional generalization by characterizing the impact of dataset statistics on compositional generalization for a broad range of tasks.
-Our theory systematically clarifies the compositional generalization behavior of deep neural networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
+Overall, we take a step towards a general theory of compositional generalization.
+Our theory systematically clarifies the compositional generalization behavior of deep networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
 2 Related work
 Compositional generalization.
-Compositionality is an important theme in both human and machine reasoning problems, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023).
-Recent breakthroughs have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023).
-Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017), in the hopes that different modules will specialize for different components (Schug et al., 2024).
-Constraining a network to be linear (or additive) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a, b).
+Compositionality is an important theme in both human and machine cognition, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023).
+While recent breakthroughs have led to massive improvements in models’ compositional capacities, these models can still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023).
+Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017).
+Constraining a network’s compositional function (e.g.
+to adding up a value for each component) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a; b; Schug et al., 2024).
 However, these networks are extremely limited in the tasks they can learn (as we will show below) and end-to-end training of modular architectures without such constraints often does not result in the desired specialization (Bahdanau et al., 2019; Mittal et al., 2022; Jarvis et al., 2023).
-Conjunctions.
-Henceforth, we will call a combination of components a conjunction.
-Conjunctive codes have a long history in neuroscience (Alvarado & Rudy, 1992; Baker et al., 2002), and are theoretically linked to the formation of highly specific episodic memories.
 Kernel regime.
-When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(z)=w^{T}\phi(z)$, it can be shown that the resulting behavior of that model can be described in terms of the kernel $K(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ induced by the representation $\phi(z)$ (Schölkopf, 2000).
-We will refer to such models as kernel models.
-Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, a kernel model can be written in its “dual form” as
-$f_{a}(z)=\sum_{i=1}^{n}a_{i}K(z,z_{i}),\mbox{ for dual coefficients}a\in\mathbb{R}^{n}.$ (1)
-Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by gradient descent on a model with a fixed representation (a kernel model) (Jacot et al., 2018).
-This is called the “kernel regime,” in contrast to the feature-learning regime, which is brought forth, for example, by small initial weights or small width (Chizat et al., 2019).
+When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(x)=w^{T}\phi(x)$, the behavior of that model can be described in terms of the kernel $K_{\phi}(x,x^{\prime})=\phi(x)^{T}\phi(x^{\prime})$ induced by $\phi(x)$ (Schölkopf, 2000).
+Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, such “kernel models” can be written in dual form as $f_{a}(x)=\sum_{i=1}^{n}a_{i}K_{\phi}(x,x_{i})$, for inferred dual coefficients $a\in\mathbb{R}^{n}$.
+Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by a kernel model (“kernel regime”) (Jacot et al., 2018).
+In contrast, the “feature-learning regime” (which can be brought forth, for example, by small initial weights or small width) yields more substantial changes in neural networks’ internal representations (Chizat et al., 2019).
 Norm minimization.
 Gradient descent, ridge regression, and neural networks in the kernel regime learn the readout weights that describe the training data with minimal $\ell_{2}$-norm (Soudry et al., 2018; Gunasekar et al., 2018; Ji et al., 2020).
-Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021b, 2023) and we here apply it to the compositional task setting.
+Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021a; b; 2023) and we here apply it to the compositional task setting.
 This is similar to Lippl et al.
@@
 Shortcut learning refers to models exploiting spurious correlations between certain features of the input data and the target (Shah et al., 2020; Nagarajan et al., 2021).
-This substantially impacts their performance out of distribution (where those correlations may not hold) (Geirhos et al., 2020).
+This substantially impacts their performance out of distribution, where those correlations may not hold (Geirhos et al., 2020).
 We theoretically analyze how shortcut biases impact compositional generalization.
@@
 3.1 Task space
-We assume that our model extracts from its input a set of disentangled components $(z_{c})_{c=1}^{C}$, e.g.
-color, shape, and background (Fig.
-1a).
-For instance, the model could have been pre-trained to extract these components from an image or an input sentence.
-Each component $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components represented by $d$-dimensional real vectors and because we consider categorical components, we assume that all components have unit magnitude and are orthogonal to each other (these assumptions can be relaxed, Section A.1).
-We then assume that the model represents these components by adding them together: $z:=\sum_{c=1}^{C}z_{c}$.
-A simple example of this is given by a concatenation of one-hot representations of all components (Fig.
-1a).
+We consider an input $x$ representing a set of underlying components $z=(z_{c})_{c=1}^{C}$, where each component $z_{c}\in Z_{c}$ is drawn from a discrete, finite set of possible components.
+For example, $x$ could be a “multi-hot” concatenation of one-hot representations of all components (Fig.
+1c).
+We make a specific assumption about how the trial-by-trial similarity of input representations $x$ relates to the underlying components $z$ (Fig.
+1b):
+Definition 3.1.
+A representation $x$ is “compositionally structured” iff its kernel $K(x,x^{\prime})=x^{T}x^{\prime}$ only depends on the number of components that are identical between $z$ and $z^{\prime}$, where $z$ and $z^{\prime}$ are the components represented by $x$ and $x^{\prime}$ respectively.
+In particular, the multi-hot representation described above is compositionally structured, but we will see below that this concept captures a richer set of representations as well.
 The target, $y\in\mathbb{R}$, is given by an arbitrary function of $z$, ensuring that our framework is agnostic to the underlying compositional structure.
 After training models on certain combinations of components $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$.
-Our theory applies to any task in this family; that is, any task with a fixed number of categorical input components and a scalar output, regardless of the input-output mapping.
-Below we describe three example tasks within this general framework that represent important building blocks for compositional reasoning in machine learning and cognitive science.
+Our theory characterizes constraints on model behavior for arbitrary tasks within this family, regardless of the ground-truth input-output mapping.
+Below we describe three example tasks that represent important building blocks for compositional reasoning in machine learning and cognitive science.
 We will focus on these tasks in the main text, to illustrate our general theory and analyze its consequences in more detail.
-In Appendix D, we describe additional tasks captured by our theory.
+In Appendix E, we describe additional tasks captured by our theory (Fig.
+2d).
+Figure 2: The compositional task space.
+a, Example training sets for symbolic addition.
+The grid represents pairs of components with associated values $-4,\dotsc,4$.
+The training set consists of certain rows and columns of this grid.
+b, Context dependence.
+In context 1, feat.
+1 determines the category; in context 2, feat.
+2 determines the category.
+The training sets leave out subsets of the lower right orthant.
+c, Transitive equivalence: six items are split up into two equivalence classes (e.g.
+A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes (e.g.
+$A=B$ and $B=C$ implies $A=C$).
+d, Our theory partitions the task space into conjunction-wise additive (which can be solved by kernel models) and non-additive tasks (which cannot).
 Symbolic addition.
@@
 adding handwritten digits) (Lorenzi et al., 2021; Sheahan et al., 2021).
-We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$.
+We consider two components $(z_{1},z_{2})$ with hidden values $v_{1}(z_{1})$ and $v_{2}(z_{2})$.
 The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$.
-After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items.
-In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values (given by the rows and columns in the training set in Fig.
-1b).
-We vary the size of this subset (i.e.
-the number of rows/columns) as well as whether the task requires only interpolation or also extrapolation.
+After sufficient exposure to individual items in different combinations, an additive model could in principle generalize to novel combinations.
+We consider training sets consisting of entire rows and columns of the compositional dataset (Fig.
+2a), varying the number of rows and columns as well as whether generalization involves interpolation or also extrapolation.
 Context dependence.
-The relevance of different stimuli often depends on our current context.
-Taking this into account is a crucial aspect of cognition in humans and animals (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022).
-We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$.
+The relevance of different stimuli often depends crucially on our current context (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022).
+We consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$.
 The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response.
-Both features have six possible values which are split up in two categories.
-If the model has learned this context dependence, it should be able to generalize to novel feature combinations.
+Features have six possible values which are divided into two classes.
+If the model has learned this context dependence, it should generalize to novel feature combinations.
 We evaluate on the subset of data for which $z_{f1}$ indicates Cat.
@@
 1.
-In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of two or one of those features (CD-2 and CD-1) (Fig.
-1c).
+In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of one or two features (CD-1, CD-2) (Fig.
+2b).
 Transitive equivalence.
-Relational reasoning is an important example of compositional generalization and often involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018).
-Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent.
+Relational reasoning involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018).
+Given an unobserved (and arbitrary) equivalence relation, the task here is to determine whether two presented items $(z_{1},z_{2})$ are equivalent.
 The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig.
-1d).
+2c).
 This is an important instance of relational cognition (often studied as “associative inference” in cognitive science (Schlichting & Preston, 2015; Spalding et al., 2018)).
@@
 3.2 Models
-Kernel models.
-Our theory characterizes kernel models with a disentangled input $z\in\mathbb{R}^{d}$ that apply a transform $\phi(z)\in\mathbb{R}^{h}$ and learn a linear readout, $f_{w}(x):=w^{T}\phi(z)$, using gradient descent or ridge regression (Fig.
-1a).
+Our theory characterizes kernel models with a compositionally structured input $x\in\mathbb{R}^{d}$ that apply a transform $\phi(x)\in\mathbb{R}^{h}$ and learn a linear readout, $f_{w}(x):=w^{T}\phi(x)$, using gradient descent with initial weights $w_{0}=0$.
 For $\phi$, we consider a neural network with random weights (in the infinite-width limit; see Section A.3).
-Our model setup includes as a special case the random feature model studied by Abbe et al.
-(2023) and further captures training via backpropagation in the kernel regime.
-Deep neural networks.
-Additionally, we consider a broad range of deep neural network architectures: convolutional networks, residual networks (He et al., 2016), Vision Transformers (ViTs; Dosovitskiy et al., 2020)), and ReLU networks; all trained through backpropagation.
-The code required to reproduce all experiments can be found at https://github.com/sflippl/compositional-generalization.
-Figure 1: a, Theoretical setup: a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$.
-b, Example training sets for symbolic addition.
-The grid represents nine components with associated values $-4,-3,\dotsc,4$.
-The training set consists of certain rows and columns of this grid.
-c, Context dependence.
-In context 1, feat.
-1 determines the category; in context 2, feat.
-2 determines the category.
-The training sets leave out different subsets of the lower right orthant.
-d, Transitive equivalence: six items are split up into two arbitrary equivalence classes (e.g.
-A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes.
+Our setup captures, for example, the random feature model studied by Abbe et al.
+(2023) and training via backpropagation in the kernel regime.
 4 Kernel models are conjunction-wise additive
-Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs.
-We find that even though these models can learn arbitrary training sets, their test set behavior is restricted to adding up values assigned to each conjunction (combination of components) seen during training.
+Our primary theoretical contribution is to characterize the full range of compositional computations that can be implemented by kernel models with compositionally structured inputs.
+We find that even though these models can learn arbitrary training sets, their test set behavior is restricted to adding up values for each conjunction (combination of components) seen during training.
 We call this motif “conjunction-wise additivity.” Below, we formally state our finding and explain its implications.
 4.1 Random networks yield compositionally structured representations
-We first note that a linear readout of the disentangled input representation effectively constrains the model to adding up a value for each component (“component-wise additivity”):
-$f(z)=w^{T}z=w^{T}\sum_{c=1}^{C}z_{c}=\sum_{c=1}^{C}f_{c}(z_{c}),\quad f_{c}(z_{c}):=w^{T}z_{c}.$ (2)
+We first note that a linear readout of the multi-hot input constrains the model to adding up a value for each component (“component-wise additivity”): $f(x)=\sum_{c=1}^{C}f_{c}(z_{c})$.
 Although models of this class perfectly generalize on component-wise additive tasks, such as symbolic addition, they are incapable of even learning the training data (much less generalizing) on tasks that are not component-wise additive.
 In particular, these models cannot learn context-dependent computations or equivalence relations — both fundamental instances of compositional reasoning.
-The nonlinear transform $\phi(z)$ can overcome this constraint; in particular, multi-layer nonlinear neural networks can learn arbitrary training data (in the infinite-width limit) (Hornik et al., 1989; Cybenko, 1989; Rigotti et al., 2013).
-In general, the wide range of possible neural network architectures and their nonlinearities makes it difficult to derive general statements about their representations, but in this case we can take advantage of a particularly useful property of disentangled categorical inputs.
-That is, its induced kernel, $K(z,z^{\prime})=z^{T}z^{\prime}$, depends only on the number of components that are identical between both inputs (specifically, $K(z,z^{\prime})=c$ if $c$ components are identical).
-We call any representation that satisfies this criterion “compositionally structured” (see Section A.2 for further motivation).
-Notably, in the infinite-width limit, the kernel of random weight neural networks, $K_{\phi}(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$, depends only on the input kernel $K(z,z^{\prime})$, making $K_{\phi}(z,z^{\prime})$ also compositionally structured (Cho & Saul, 2009, see Section A.3).
-This representational constraint allows us to derive strong constraints on the models’ generalization behavior.
+The nonlinear transform $\phi(x)$ can overcome this constraint; in particular, multi-layer nonlinear neural networks can learn arbitrary training data (in the infinite-width limit) (Hornik et al., 1989; Cybenko, 1989; Rigotti et al., 2013).
+Normally, the wide range of possible network architectures makes it difficult to derive general statements about their representations, but in this case we can take advantage of the fact that $\phi(x)$ will also be compositionally structured (Fig.
+1c):
+Proposition 4.1.
+For a random weights neural network $\phi$ with a compositionally structured input $x$, in the infinite-width limit, $\phi(x)$ is also compositionally structured.
+Proposition 4.1 holds because in the infinite-width limit, the kernel of random weight neural networks depends only on the input kernel (Cho & Saul, 2009, see Section A.3).
+We now consider the constraints that compositional structure imposes on the models’ generalization behavior.
 4.2 Compositionally structured kernel models are conjunction-wise additive
@@
 To formally state this finding, we define, for each $z\in Z$, the set of conjunctions for which $z$ overlaps with some element in the training set $z^{\text{tr}}\in Z^{\text{train}}$,
-$\text{Conj}(z|Z^{\text{train}}):=\left\{J\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}:\forall c\in J:z_{c}=z^{\text{tr}}_{c}\right\}.$ (3)
-We then prove:
+$\text{Conj}(z|Z^{\text{train}}):=\left\{J\subseteq\{1,\dotsc,C\}|\mbox{there is }z^{\text{tr}}\in Z^{\text{train}}\mbox{ such that }z_{c}=z_{c}^{\text{tr}}\mbox{ for all }c\in J\right\}.$ (1)
 Theorem 4.2.
-Consider a kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$.
-For any such model, we can find conjunction-wise functions $f_{J}:\prod_{c\in J}Z_{c}\to\mathbb{R}$, where $J\subseteq\{1,\dotsc,C\}$, such that for any input $z\in Z$, the model response is given by
-$f(z)=\sum_{J\in\text{Conj}(z|Z^{\text{train}})}f_{J}(z_{J}),\quad z_{J}:=(z_{c})_{c\in J}.$ (4)
-We prove Theorem 4.1 in Section A.4, but to develop an intuition for why this theorem holds, we consider an example representation concatenating one-hot vectors representing each component and conjunction of components.
-For example, if $C=2$, this representation will concatenate one-hot representations of each component with a one-hot representation of the conjunction of both components.
-The readout weight for a conjunction $z_{J}$ seen during training will change (in Eq.
-4, this weight is given by $f_{J}(z_{J})$).
-In contrast, the weights associated with conjunctions not seen during training remain at their initial value of zero (which is why we only sum over $J\in\text{Conj}(z|Z^{\text{train}})$).
+For any kernel model $f$ with a compositionally structured representation, we can find conjunction-wise functions $f_{J}:\prod_{c\in J}Z_{c}\to\mathbb{R}$, where $J\subseteq\{1,\dotsc,C\}$, such that for any input $x\in\mathbb{R}^{d}$ representing components $z\in Z$, the model response is given by
+$f(x)=\sum_{J\in\text{Conj}(z|Z^{\text{train}})}f_{J}(z_{J}),\quad z_{J}:=(z_{c})_{c\in J}.$ (2)
+We prove the theorem in Section A.4.
+It implies that for a given test input, kernel models add up a value for each component and partial conjunction seen during training (Fig.
+1d).
+We call this computation “conjunction-wise additive.” In particular, for inputs with two components, the model’s behavior on the training set can be expressed as $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z_{1},z_{2})$.
+Thus, $f_{12}(z_{1},z_{2})$ enables the model to learn arbitrary training data.
+However, if $x$ represents $z\in Z^{\text{test}}$, the training set does not contain any input with the conjunction $(z_{1},z_{2})$ and the model is component-wise additive: $f(x)=f_{1}(z_{1})+f_{2}(z_{2})$.
+For inputs with more than two components, a conjunction-wise additive computation additionally encodes partial conjunctions seen during training, e.g., if $Z^{\text{train}}$ contains the conjunction $(z_{1},z_{2})$, $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{3}(z_{3})+f_{12}(z_{1},z_{2})$.
 4.3 Conjunction-wise additivity constrains the tasks kernel models can solve
-Theorem 4.1 immediately implies that compositionally structured kernel models cannot solve any task that cannot be expressed in conjunction-wise additive terms.
-This highlights fundamental computational restrictions on a broad range of models — including pre-trained models with disentangled representations that are finetuned in the kernel regime and infinite-width neural networks.
-Intriguingly, these restrictions are not caused by an architectural constraint, as the model can learn arbitrary training data.
+Theorem 4.2 implies that compositionally structured kernel models can only solve tasks that can be expressed in conjunction-wise additive terms.
+This highlights a fundamental computational restriction, partitioning the compositional task space into tasks that can and cannot be solved by kernel models (Fig.
+2d).
+Intriguingly, these restrictions are not caused by an architectural constraint, as the model can learn arbitrary training data (at least with a nonlinear transformation $\phi$, see Section 5.1).
 Rather, we highlight restrictions on how models – without any architectural constraints – can generalize (Zhang et al., 2021).
-Below we spell out the consequences of these restrictions.
-First, for inputs with two components, the model’s behavior on the training set can be expressed as $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$.
-However, on the testing set, $f_{12}(z)$ falls away and the model is constrained to be component-wise additive: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
-A striking consequence of this is that compositionally structured models – including infinite-width neural networks and regardless of implementational details – can only generalize on component-wise additive tasks.
+We now spell out the consequences of the highlighted restrictions.
+First, for inputs with two components, we noted above that the model’s behavior on the test set is component-wise additive.
+A consequence of this is that compositionally structured models can only generalize on component-wise additive tasks.
 In particular, this implies that these models cannot generalize on transitive equivalence — a fundamental instance of relational reasoning.
@@
 This highlights the importance of a formal perspective on compositional tasks.
-For inputs with more than two components, a conjunction-wise additive computation additionally encodes partial conjunctions between components (e.g.
-12, 13, and 23 for three components).
-For test inputs, the model cannot rely on the unseen full conjunction 123, but it can rely on any partial conjunction it has seen before.
-To see whether a kernel model can, in principle, generalize on such a task, we must 1) identify the conjunctions seen during training, $\text{Conj}(z|Z^{\text{train}})$, for each $z\in Z^{\text{test}}$, and 2) determine whether the target can be written as a conjunction-wise sum.
-For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions 12 and 13, but not the feature-feature conjunction 23.
-As a result, model behavior on the test set can be written as
-$f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$
-While context dependence is not component-wise additive, it is conjunction-wise additive.
-This is because we can encode it using $f_{12}$ and $f_{13}$: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
+More generally, to see whether a kernel model can, in principle, generalize on a task with more than two input components, we must 1) identify, for each $z\in Z^{\text{test}}$, the conjunctions seen during training, $\text{Conj}(z|Z^{\text{train}})$, and 2) determine whether the target can be written as a conjunction-wise sum (Eq.
+2).
+For context dependence for example, for all $z\in Z^{\text{test}}$, we have seen the context-feature conjunctions 12 and 13, but not the feature-feature conjunction 23.
+As a result, model behavior on a test trial $x$ representing the components $z_{\text{co}},z_{f1},z_{f2}$ can be written as $f(x)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$ Importantly, this conjunction-wise additive function can encode context dependence, specifically with the following schema: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
 Thus, the model can, in principle, identify a set of weights that would allow it to generalize correctly.
 This example illustrates that conjunction-wise additivity tells us both whether kernel models with compositionally structured representations can solve a certain task, and also how they solve it.
-As we noted, these insights immediately transfer to any model with compositionally structured representations that is trained or fine-tuned in the kernel regime.
-Because our theory determines the compositional computations emerging from dataset statistics, it is also likely to shed light on other learning models, including feature-learning deep neural networks (see Section 6).
-Lastly, our theory tells us how to make a task non-additive: in Section D.2.3, we describe such a modification to context dependence that tests generalization to novel context-feature conjunctions.
+It even tells us how to make a task non-additive: in Section E.2.3, we describe such a modification to context dependence that tests generalization to novel context-feature conjunctions.
 By helping us design hard benchmark tasks that are not solvable by kernel models, our framework grounds research into learning mechanisms that can implement other compositional computations.
+What do our findings imply about compositional generalization in more general representations?
+In this section, we analyzed compositionally structured inputs, as we consider them the best-case scenario for compositional generalization.
+Accordingly, we expect that the limitations revealed by our theory are also relevant for kernel models with non-compositionally structured representations.
+We present two pieces of preliminary evidence for this conjecture in the appendix: First, we theoretically prove that randomly sampled representations (that are only compositionally structured in expectation) are also conjunction-wise additive in expectation (Section A.5).
+Second, we empirically investigate compositional generalization in disentangled representation learning models trained on the DSprites dataset (Locatello et al., 2019), finding that, when averaged across randomly sampled task instances, these models are also limited to conjunction-wise additive computations (Appendix C).
+These findings suggest that the generalization class we have highlighted in this section captures fundamental limitations of kernel models beyond our theoretical setting.
 5 Representational geometry still impacts generalization on conjunction-wise additive tasks in kernel models
@@
 However, model generalization also depends on whether the model identifies the correct conjunction-wise function.
-This depends on the model’s representational geometry and the dataset statistics.
-As noted in Section 2, the behavior of kernel models depends on their representation $\phi(z)$ only through the induced kernel $K_{\phi}(z,z^{\prime})$.
-In this section, we employ the modular approach suggested by this insight: We first investigate how different choices in network architecture influence its induced kernel (Section 5.1).
-Then, we characterize task performance on symbolic addition and context dependence across the full range of compositionally structured representations (Section 5.2).
+This depends on the model’s representational geometry and the training data statistics.
+As noted in Section 2, the behavior of kernel models depends on their representation $\phi(x)$ only through the induced kernel $K_{\phi}(x,x^{\prime})$.
+In this section, we first investigate how different choices in network architecture influence its induced kernel (Section 5.1).
+Then, we characterize how this kernel impacts task performance on symbolic addition and context dependence (Section 5.2).
 5.1 Overlap salience characterizes compositional representational geometry
+To characterize compositionally structured representations, we introduce a new metric: representational salience.
+We formally define this metric in Section B.1.
+Intuitively, for $k=1,\dotsc,C$, the salience $S(k;C)$ measures the unique contribution of the subpopulation representing a conjunction of $c$ components.
+Further, $S(k;C)$ is normalized so that all saliences sum up to one.
+For example, the multi-hot input only encodes single components and therefore $S(1;C)=1/C$ and, for all $k>1$, $S(k;C)=0$.
+Note that the model can only use conjunctions whose salience is nonzero.
+For example, the multi-hot input is constrained to a component-wise sum (as all other saliences are zero).
+$S(k;C)$ is computed from the representational similarities $K_{\phi}(x,x^{\prime})=\phi(x)^{T}\phi(x^{\prime})$.
+This metric is useful because unlike $K_{\phi}(x,x^{\prime})$, it makes immediately apparent how strongly different conjunctions are encoded (Section B.2).
+Furthermore, it reduces the $C+1$ similarities characterizing the space of compositionally structured representations to $C-1$ free parameters, removing the overall magnitude of the representation (by normalizing) and the baseline activity (by not considering the similarity between inputs with no overlap).
+This is because these parameters often have negligible impact on model behavior (Section B.3).
 Figure 3: Representational salience (for three input components) in a random weight neural network with variable numbers of layers and nonlinearities.
-In Section 4, we found that a broad range of transformations yield compositionally structured representations, i.e.
-representations that have the same similarity $K(z,z^{\prime})$ for all inputs $z,z^{\prime}$ overlapping in the same number of components.
-To characterize representations in this space, we introduce a new metric: representational salience.
-We formally define this metric in Section B.1.
-Intuitively, for $c=1,\dotsc,C$, the salience $S(c;C)$ measures the unique contribution of the subpopulation representing a conjunction of $c$ components.
-Further, $S(c;C)$ is normalized so that the salience of all components and conjunctions of components sums up to one.
-For example, the input representation only encodes single components and therefore $S(1;C)=1/C$ and, for all $c>1$, $S(c;C)=0$.
-$S(c;C)$ is computed from the representational similarities $K(z,z^{\prime})$.
-It is useful because unlike those similarities, it makes immediately apparent how strongly different conjunctions are encoded (Section B.2).
-Further, it reduces the set of $C+1$ similarities characterizing the space of compositionally structured representations to $C-1$ free parameters: It removes the overall magnitude of the representation (by normalizing) and the baseline activity (by excluding the similarity between inputs with no overlap).
-This choice reflects the fact that both of these parameters often have negligible impact on model behavior (Section B.3).
-We now analyze how $S(c;C)$ evolves over different layers of a random network, using an input with three components as an example (Fig.
-2; see Section B.4 for other values of $C$).
-At the input stage, only individual components are represented and so $S(2;3)$ and $S(3;3)$ are zero.
-As the network gets deeper, these saliences increase, whereas $S(1;3)$ decreases.
-Eventually, the salience of the intermediate conjunction $S(2;3)$ decreases again, whereas $S(3;3)$ continues to increase.
-Indeed, for ReLU networks, we prove the following statement:
-Proposition 5.1.
-For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $S(k;C)\to 0$ for $k<C$ and $S(C;C)\to 1$.
-We prove the proposition in Section B.5. It implies that in very deep networks, the full conjunction dominates the representation, making compositional generalization to overlapping inputs harder.
+We now analyze how $S(k;C)$ evolves over different layers of a random network, using three components as an example (Fig.
+3; see Section B.4 for other values of $C$).
+The input only represents individual components and so $S(2;3)$ and $S(3;3)$ are zero.
+As the network gets deeper, these saliences increase and $S(1;3)$ decreases.
+In particular, because $S(3;3)>0$ for all networks with at least one layer, the resulting representations can learn arbitrary training data (using the full conjunction).
+As depth increases, $S(2;3)$ eventually decreases again, whereas $S(3;3)$ continues to increase.
+In fact, we prove that for ReLU networks in the limit of infinite depth, $S(C;C)\to 1$ (Proposition B.2).
+Put differently, very deep networks exclusively encode the full conjunction.
+As a result, these models effectively implement a look-up table, memorizing the training data without generalizing to novel combinations of components.
+In the next section, we find that even a moderately conjunctive representation presents significant challenges to compositional generalization.
 5.2 Kernel models suffer from memorization leak and shortcut bias
-While there are many readout weights that can lead to minimal error, kernel models trained with gradient descent or ridge regression learn the weights with minimal $\ell_{2}$-norm (see Section 2).
-We now characterize how this inductive bias influences compositional generalization across different representational geometries.
-To characterize the full range of compositionally structured representations, we compute the behavior of the model directly using the kernel (see Appendix C).
-Importantly, as implied by our modular approach, our analysis clarifies the behavior of all random weight neural networks, including those covered in the previous section.
+While there are many readout weights that can lead to minimal error, kernel models trained with gradient descent learn the weights with minimal $\ell_{2}$-norm (see Section 2).
+We now characterize how this inductive bias influences compositional generalization.
+To characterize the full range of compositionally structured representations, we compute the behavior of the model directly using the kernel (see Appendix D).
+Importantly, our analysis clarifies the behavior of all random weight neural networks, including those covered in the previous section.
 Symbolic addition suffers from a memorization leak.
-We first consider an example case ($S(1;2)=0.4,\mathcal{W}=\{0\}$).
+We first consider an example task ($z_{1},z_{2}\in\{[-4],\dotsc,[4]\}$; the training set consists of all trials containing a $[0]$, Fig.
+2a) and an example representation ($S(1;2)=0.4$).
 We find that though the model perfectly learns the training cases, it underestimates the test cases by a proportional factor (Fig.
-3a).
-To understand why, we consider the model’s functional form on the training cases: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ (see Section 4).
+4a).
+To understand why, we consider the model’s functional form on the training cases: $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ (see Section 4).
 Intuitively, $\ell_{2}$-norm minimization tends to yield distributed weights.
 As a result, unless the salience of the conjunction $S(2;2)$ is zero, the kernel model will associate some non-zero weight with it (i.e.
-$f_{12}(z)\neq 0$).
-This, in turn, necessarily distorts the generalization on the test set, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
-We call this tendency to use the full conjunction (and its effect on generalization) “memorization leak.” To characterize it more systematically (and add mathematical rigor to the intuition above), we analytically characterize model behavior on a general family of symbolic addition tasks:
+$f_{12}(z_{1},z_{2})\neq 0$).
+This distorts the test set generalization, $f(x)=f_{1}(z_{1})+f_{2}(z_{2})$.
+We call the tendency to use the full conjunction (and its detrimental effect on generalization) “memorization leak.” To characterize it more systematically (and add mathematical rigor to the intuition above), we analytically characterize model behavior on a general family of symbolic addition tasks:
 Proposition 5.1.
@@
 $f([v_{1}],[v_{2}])=m(v_{1}+v_{2}),\quad m:=\tfrac{p\cdot S(1;2)}{1+(p-2)S(1;2)},\quad p:=|\mathcal{W}|$ (3)
-We prove the proposition in Section D.1.
-It implies that for this entire task family, model generalization is distorted by a proportional factor $m$ (see Fig.
-3b).
+We prove the proposition in Section E.1.
+It implies that for this entire task family, model generalization is distorted by a proportional factor $m$ (Fig.
+4b).
 The formula implies that $m<1$ as long as $S(1;2)<\tfrac{1}{2}$ (i.e.
@@
 (Note that $\mathcal{W}$ corresponds to the rows/columns making up the example training sets in Fig.
-1b.) Interestingly, these are the only two factors that influence $m$.
-In particular, while interpolation is often seen as easier than extrapolation, this does not impact model behavior here.
+2a.) Interestingly, these are the only two factors that influence $m$.
+In particular, even though interpolation is often seen as easier than extrapolation, this does not impact model behavior here.
 More broadly, the memorization leak is a ubiquitous issue for statistical learning models trained on compositional tasks: $\ell_{2}$-norm minimization tends to yield distributed weights and so if the conjunctive population is represented, the model will generally end up partially relying on this conjunction.
-As shown above, this necessarily distorts generalization (see also Section D.3 for a minimal example).
+As shown above, this necessarily distorts generalization.
+This issue also arises in tasks that involve directly decoding specific components, a popular task for evaluating disentangled representations (Locatello et al., 2019; Schott et al., 2022) (see Section E.3 for a minimal example).
 Figure 4: Kernel models’ behavior on (a,b) symbolic addition and (c,d) context dependence.
@@
 Trajectories of networks with different nonlinearities are highlighted (color scale see Fig.
-2).
+3).
 d, Coefficients of the different conjunction types for two example networks with three layers and different nonlinearities.
@@
 For CD-3 (the task version leaving out the largest subset of feature combinations), the model only generalizes when $S(2;3)$ is high relative to $S(1;3)$ (Fig.
-3c, dark green area).
-As a result, whether the network generalizes is highly sensitive to the nonlinearity and depth of the network.
+4c, dark green area).
+As a result, compositional generalization is highly sensitive to the nonlinearity and depth of the network.
 In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
 To understand this phenomenon, we determined the total magnitude of model weights associated with the different conjunctions (Fig.
-3d).
+4d).
 We found that unsuccessful models (e.g.
-blue color in Fig.
-3d) had much larger magnitudes associated with the context component and the full conjunction.
-Notably, on CD-3, context is highly correlated with the target and can predict 2/3 of the training data.
-Models with high $S(1;3)$ (context) and $S(3;3)$ (full conjunction) exploit the context shortcut and use the full conjunction to learn the remaining training data.
+the blue color in Fig.
+4d) had much larger magnitudes associated with the context component and the full conjunction.
+Notably, on CD-3, context is highly correlated with the target and can predict $\tfrac{2}{3}$ of the training data.
+Models with high $S(1;3)$ (context) and $S(3;3)$ (full conjunction) exploit this context shortcut and use the full conjunction to learn the remaining training data.
 This strategy explains why these models fail to generalize to the test set.
@@
 This explains why only very low $S(2;3)$ yields failure to generalize on CD-2 or CD-1.
-Our analysis shows how model and task structure interact to either give rise to a generalizable rule or a statistical shortcut (see Section D.3 for a minimal example).
-This highlights that for compositional tasks with strong spurious correlations, model behavior will be highly sensitive to architectural details affecting the representational geometry such as depth and nonlinearity.
-6 Our theory describes the behavior of deep networks on conjunction-wise additive tasks
+Our analysis shows how model and task structure interact to either give rise to a generalizable rule or a statistical shortcut.
+This highlights that for compositional tasks with strong spurious correlations, model behavior will be highly sensitive to architectural hyperparameters affecting the representational geometry such as depth and nonlinearity.
+Indeed, this sensitivity to minor experimental choices could explain why the literature has been so divided on the usefulness of disentangled representations.
+6 Our theory can describe the behavior of deep networks on conjunction-wise additive tasks
 Figure 5: Testing our theory in deep networks trained on MNIST and CIFAR versions of compositional tasks.
@@
 For MSE instead of slope, see Fig.
-9.
+12.
 e, Accuracy on all variants of context dependence.
 So far, our analyses have been limited to simple kernel models, for their analytical tractability.
-We next test whether the insights developed from these simple models extend to a broader class of models: large-scale neural networks.
-We considered symbolic addition and context dependence on inputs created by concatenations of images from MNIST (Lecun et al., 1998) or CIFAR-10 (Krizhevsky et al., 2009).
+We next tested whether the insights developed from these simple models extend to a broader class of models: large-scale neural networks.
+We considered symbolic addition and context dependence on inputs created by concatenations of images from MNIST (Lecun et al., 1998) or CIFAR-10 (Krizhevsky et al., 2009) (see Appendix D).
 Each component corresponds to random samples from a particular category rather than the one-hot input considered so far and we study both in-distribution and compositional generalization.
@@
 To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment.
-We considered several relevant vision architectures: convolutional networks (ConvNets, LeCun et al., 1989) (trained on MNIST) and residual networks (ResNets, He et al., 2016) and Vision Transformers (ViTs, Dosovitskiy et al., 2020) (trained on CIFAR-10).
-Spatial distance of components impact salience in internal representations.
+We considered several relevant vision architectures, which we all trained with backpropagation: convolutional networks (ConvNets, LeCun et al., 1989) (trained on MNIST) and residual networks (ResNets, He et al., 2016) and Vision Transformers (ViTs, Dosovitskiy et al., 2020) (trained on CIFAR-10).
+Spatial distance of components impacts salience in internal representations.
 First, we examined how changes in the input structure impact the networks’ representational geometry.
@@
 We found that $S(1;2)$ was indeed smaller for lower distances (Fig.
-4a), indicating a more conjunctive population.
+5a), indicating a more conjunctive population.
 This suggests that varying the distance between two digits provides a practical way of manipulating $S(1;2)$; below we use this insight to test predictions about how conjunctivity influences generalization behavior.
-Deep networks implement a conjunction-wise additive computation.
+Deep networks tend to implement conjunction-wise additive computations.
 Conjunction-wise additivity imposes a substantial computational restriction: on test set inputs, models can only add up values assigned to each conjunction of components seen during training.
-We therefore investigated whether the deep networks’ computations could be captured in these terms.
-Remarkably, we found that such a model was highly predictive of the model responses (see Section C.2).
-This indicates that large-scale neural networks tend to implement conjunction-wise additive computations — at least when trained on conjunction-wise additive tasks.
+We therefore investigated whether the deep networks’ computations could be captured in these terms — a nontrivial prediction as these networks can in principle express arbitrary compositional computations.
+Remarkably, we found that a conjunction-wise additive model was highly predictive of the model responses (see Section D.3).
+This indicates that large-scale neural networks tend to implement conjunction-wise additive computations, at least when trained on conjunction-wise additive tasks.
 Our theory suggests that they do so because the statistical structure of the dataset makes such a computation natural.
@@
 Having found that our theory captures the networks’ general computational structure, we investigated whether it was also able to explain their specific performance.
-We first considered symbolic addition, training the ConvNets on seven different training sets with 20,000 samples and the ResNets and ViTs on three training sets with 40,000 samples.
-We varied both the size of $\mathcal{W}$ and whether generalization required interpolation or also extrapolation (Section D.1.3).
+We first considered symbolic addition, varying the size of the training set and whether generalization required interpolation or also extrapolation.
+We trained the ConvNets on 20,000 samples and the ResNets and ViTs on 40,000 samples (Section E.1.3).
 We tested three theoretical predictions made by Proposition 5.1.
@@
 We found that model predictions were indeed compressed relative to the ground truth, but still exhibited a strong linear relationship (Fig.
-4b).
+5b).
 This was also the case for ResNets and ViTs, which exhibited a slightly noisier linear relationship (Fig.
-8).
+11).
 The memorization leak therefore affects generalization in large-scale networks as well.
@@
 increasing distance between components, Fig.
-4a).
+5a).
 On ConvNets, we found that higher distance indeed increases the slope (Fig.
-4c).
+5c).
 This was not due to an increase in task difficulty, as in-distribution generalization is not systematically affected by distance.
@@
 (Notably, these networks did not perfectly predict the training split, which may affect the results.)
-Finally, our theory predicts that larger $\mathcal{W}$ should increase the compositional generalization slope.
+Finally, our theory predicts that larger training sets should increase the compositional generalization slope.
 Our experiments confirmed this prediction (Fig.
-4d).
+5d).
 Further, whether the training set required interpolation or also extrapolation did not systematically affect the resulting slope, as predicted by our theory (though interestingly, it did affect in-distribution generalization).
-However, there was one exception to this: for $p=4$, the extrapolation dataset had a smaller slope than the interpolation data set.
+Notably, there was one exception to the latter finding: for $p=4$, the extrapolation dataset had a smaller slope than the interpolation data set.
 Our theory can therefore explain much but not all of the deep network behavior.
@@
 Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but having worse-than-chance accuracy on CD-3 (Fig.
-4e).
-This confirms that the deep networks also suffered from a context-driven shortcut on CD-3.
-7 Outlook: feature learning enables non-additive compositional generalization
-In the previous section, we found that neural networks trained on symbolic addition or context dependence implement a conjunction-wise additive computation.
-However, conjunction-wise additivity is not the only mechanism by which deep neural networks generalize compositionally.
-To demonstrate this, we show in Appendix E that feature-learning networks can generalize on transitive equivalence (a non-additive task) due to their tendency to learn low-dimensional internal representations.
-This highlights that neural networks can implement non-additive compositional computations.
-Future work should explore more systematically how feature learning impacts compositional generalization.
+5e).
+This confirms that the deep networks also suffered from a context-driven shortcut on CD-3 (see also Fig.
+16).
 7 Discussion
-Humans often generalize to new situations by stitching together concepts and knowledge from prior experience in new ways.
+Humans often generalize to new situations by stitching together concepts and knowledge from prior experience.
 Despite the broad importance of this ability (both for cognitive science and machine learning), a general theory of when and how neural networks accomplish compositional generalization has remained elusive.
-Here we have taken a step towards formalizing this relationship by characterizing kernel models with compositionally structured representations, a framework that captures neural network training in the kernel regime and more broadly lets us understand the impact of dataset statistics on generalization.
+Here we have taken a step towards formalizing this relationship by characterizing kernel models with compositionally structured representations, a framework that captures neural network training or finetuning in the kernel regime and more broadly lets us understand the impact of representational geometry on generalization.
 We found that they implement a specific compositional computation (“conjunction-wise additivity”).
@@
 Our results show how simple statistical models can implement (or at least approximate) abstract rules like context dependence.
-However, they also highlight that building in compositional structure by itself is often insufficient for compositional generalization.
+However, they also highlight that building in compositional structure is often insufficient for compositional generalization.
 Indeed, contextual generalization is highly sensitive to the specific representational geometry and training data.
-This may explain why investigations into compositional generalization and the benefits of disentangled representation have yielded such inconsistent results.
-Overall, our insights highlight the utility of kernel models in systematically investigating model generalization (or at least generating an initial hypothesis as to the important factors).
+This may explain why investigations into compositional generalization and the benefits of disentangled representations have yielded such inconsistent results.
+Overall, our insights highlight the utility of kernel models in systematically investigating model generalization.
 We here used these models to characterize deep neural networks, but they could be equally useful for better understanding human compositional generalization.
 While our work covers a broad range of different tasks, a number of limitations remain.
-Though out of scope here, it would be interesting to test this theory in extremely large-scale models, e.g.
+We demonstrate that our theory captures many qualitative phenomena in deep neural networks, but do not provide any quantitative bounds.
+Further, though out of scope here, it would be interesting to test this theory in extremely large-scale models, e.g.
 large language models.
-Future work should also consider inputs and outputs with broader ranges of formats (e.g.
-sequences with variable length) as well as representations that are not compositionally structured.
-Most importantly, our theory is limited to a particular learning mechanism (fixed features with linear readout); other learning mechanisms could either implement a conjunction-wise additive model that overcomes the failure modes highlighted here or implement novel compositional computations beyond conjunction-wise additivity (we gave an initial example of this in Section 7).
-By clarifying the general relationship between dataset statistics and compositional generalization, our theory provides an important foundation for such advances.
-Acknowledgements
+Future work should also consider other input/output formats (e.g.
+sequences with variable length) and non-compositionally structured representations.
+Most importantly, our theory is limited to a particular learning mechanism (kernel models).
+Other learning mechanisms could overcome the limitations highlighted in Sections 4 and 5 (we give an initial example in Appendix F).
+By helping us design new benchmark tasks not solvable by kernel models, our theory provides an important foundation for such advances.
+Acknowledgments
 We are grateful to the members of the Center for Theoretical Neuroscience for helpful comments and discussions.
-We thank Elom Amematsro, Ching Fang, and Matteo Alleman for detailed feedback.
+We thank Elom Amematsro, Ching Fang, and Matteo Alleman for detailed feedback, and Zeb Kurth-Nelson for comments on the manuscript.
 The work was supported by NSF 1707398 (Neuronex), Gatsby Charitable Foundation GAT3708, and the NSF AI Institute for Artificial and Natural Intelligence (ARNI).
@@
 For example, in transitive equivalence (Fig.
-1d), if items that are in the same equivalence class are represented as more similar to each other, that would serve as a basis for generalization on the task that is not compositional in nature.
+2d), if items that are in the same equivalence class are represented as more similar to each other, that would serve as a basis for generalization on the task that is not compositional in nature.
 For this reason, we are particularly interested in compositionally structured representations, as they ensure that kernel models are only able to generalize compositionally.
 Characterizing compositional generalization in non-compositionally structured representations requires us to meaningfully distinguish between compositional and non-compositional aspects of their generalization.
-We consider this question out of scope for this manuscript, but believe it would be interesting future work.
+We present an example of such an analysis in Section A.5, proving that if the non-compositional representational components are random, the models still generalize in a conjunction-wise additive manner in expectation.
 A.3 Nonlinear transformations and compositional structure
@@
 Any nonlinear transform that meets this condition also conserves compositional structure, i.e.
-if $z$ is compositionally structured then so is $\phi(z)$.
+if $x$ is compositionally structured then so is $\phi(x)$.
 This allows us to derive computational restrictions on this broad range of models.
 A.4 Proof of Theorem 4.2
-See 4.1
+See 4.2
 Proof.
 Because the kernel $K$ is compositionally structured, its similarity $K(z,z^{\prime})$ only depends on the overlap $O(z,z^{\prime})\subseteq\{1,\dotsc,C\}$ (see definition in Eq.
-6).
+4).
 We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define set of training items overlapping with $z\in Z^{\text{test}}$ in $S$ as
@@
 proves the proposition. ∎
-Appendix B Representation salience: a metric for compositionally structured representations
+A.5 Non-compositionally structured representations
+So far, we have considered representations that are compositionally structured.
+In practice, however, representations will almost never be exactly compositionally structured.
+Even when there is no specific similarity structure within components, there will be random noise in the representations.
+Here we characterize one such scenario, demonstrating that in expectation, these representations still yield conjunction-wise additive computations.
+Proposition A.2.
+Consider an input with $C$ components.
+We consider a representation that represents each component and conjunction of components $z_{J}$, $J\subseteq\{1,\dotsc,C\}$ by a random vector $x_{J}[z_{J}]\in\mathbb{R}^{d}$, $x_{J}\sim\mathcal{N}(0,\sigma_{k}^{2})$, $\sigma_{k}^{2}>0$, where $k=|J|$.
+The representation itself is given by the sum of all these vectors: $x=\sum_{J\subseteq\{1,\dotsc,C\}}x_{J}[z_{J}]$.
+All $x_{J}[z_{J}]$ are sampled independently from each other.
+This means that the similarity between different components as well as the entanglement of different components varies randomly.
+As a result, the kernel regression estimator $f(z)$ is not conjunction-wise additive.
+However, its expectation, $\mathbb{E}[f(z)]$, is conjunction-wise additive, indicating that all deviations from conjunction-wise additivity arise from random noise and cannot be used for systematic generalization.
+Proof.
+We consider the training representation $X=(x[z_{J}])_{z\in L^{\text{train}}}$, $X\in\mathbb{R}^{N_{\text{train}}\times d}$ and the test representation $\tilde{X}=(x[z_{J}])_{z\in L^{\text{test}}}$, $\tilde{X}\in\mathbb{R}^{N_{\text{test}}\times d}$.
+This gives rise to the training kernel $K=XX^{T}$ and the train-test kernel, $\tilde{K}=\tilde{X}X^{T}$.
+The dual coefficients are given by $a=K^{-1}y$.
+We now consider a test set input $\tilde{x}=\tilde{x}[\tilde{z}]$ representing the underlying components $\tilde{z}$.
+Denoting the similarity to the training set by $k(\tilde{x},X)=X\tilde{x}\in\mathbb{R}^{N_{\text{train}}}$, the test set behavior is given by
+$f(\tilde{x})=a^{T}k(\tilde{x},X)=\sum_{J\in\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]+\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]=:f_{1}(\tilde{x})+f_{2}(\tilde{x}),$ (11)
+where we’ve simply split up the sum into the conjunction-wise additive part and the remainder.
+Clearly, this remainder, $f_{2}(\tilde{x}):=\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]$, will generally not be zero.
+However, we will now prove that $\mathbb{E}[f_{2}(\tilde{x})]=0$.
+As $f_{1}(\tilde{x})$ is conjunction-wise additive, this will prove the proposition.
+To do so, we note that we can assume that we have first sampled all Gaussian vectors relevant for the training set (we will denote this set of random variables by $\mathcal{X}^{\text{train}}$) and subsequently sample the set of Gaussian vectors only relevant for the test trial, i.e.
+$(\tilde{x}_{J})_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}$ (we will denote this set of random variables by $\mathcal{X}^{\text{test}}$).
+Then,
+$\mathbb{E}[f_{2}(\tilde{x})]=\mathbb{E}_{\mathcal{X}^{\text{train}}}\left[\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[f_{2}(\tilde{x})|\mathcal{X}^{\text{train}}\right]\right]=\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}\mathbb{E}_{\mathcal{X}^{\text{train}}}\left[a^{T}X\mathbb{E}_{\mathcal{Z}^{\text{test}}}\left[x_{J}[\tilde{z}_{J}]|\mathcal{Z}^{\text{train}}\right]\right].$ (12)
+This is zero, as $\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[x_{J}[z_{J}]\right]=\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[x_{J}[z_{J}]|\mathcal{X}^{\text{train}}\right]=0$, which follows from the fact that $x_{J}[z_{J}]$ for $J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})$ is sampled independently from $\mathcal{X}^{\text{train}}$.
+∎
+Figure 6: Analysis of randomly sampled representations as considered in Proposition A.2.
+In all cases we sample 100-dimensional representations.
+For addition, we consider representations with weights that explore $S(1;2)\in[0.1,1]$ and for context dependence, we consider representations where $(\sigma_{1},\sigma_{2},\sigma_{3})\in\{(0.5,1,0.1),(1,0.5,0.1)\}$, where $\sigma_{k}$ indicates the standard deviation of the Gaussian vector representing conjunctions of $k$ components.
+a, We compare the additivity of individual models to the additivity of the model behavior averaged across 100 randomly sampled representations.
+While individual models may be severely non-additive, their average behavior is consistently highly additive, empirically confirming our proposition.
+b, As a result, when these models are trained on transitive equivalence, they also systematically exhibit chance performance, as they are still, on average, constrained to be conjunction-wise additive.
+c, We then estimated the slope of the model trained on symbolic addition for each random instance of a representation against corresponding representation’s $S(1;2)$.
+The individual slope estimates are depicted by the small translucent dots, whereas the average across all random instances is depicted by the larger points.
+Finally, our theoretical predictions in the exactly compositionally structured case are depicted by the lines.
+The average model behaviors are well described by our theory.
+d, We plot the distribution of generalization accuracies on context dependence across different model seeds, the two considered representations, and for CD-3 and CD-2.
+On CD-3, the representation with a higher salience for individual components (rep.
+1, $(\sigma_{1},\sigma_{2},\sigma_{3})=(0.5,1,0.1)$ performs systematically below chance, whereas the representation with a higher salience for conjunctions of two components (rep.
+2, $(\sigma_{1},\sigma_{2},\sigma_{3})=(1,0.5,0.1)$) generalizes above chance.
+In contrast, on CD-2, both representation exhibit better-than-chance-accuracy.
+To test our theory empirically, we sample a range of Gaussian representations and train them on symbolic addition, transitive equivalence, and context dependence.
+For symbolic addition and transitive equivalence, we consider representations with expected $S(1;2)\in[0.1,1]$ using ten equally spaced values (i.e.
+$0.1,0.2,\dotsc$).
+For context dependence, we consider two types of representation: rep.
+1, where $\sigma_{1}=1,\sigma_{2}=0.5,\sigma_{3}=0.1$, (i.e.
+single components are most saliently represented), and rep.
+2, where $\sigma_{1}=0.5,\sigma_{2}=1,\sigma_{3}=0.1$ (i.e.
+conjunctions of two components are most saliently represented).
+We consider representations with $d=100$ and sample 100 instances of each representations.
+We then fit each representation to these three tasks.
+We first estimate these models’ additivity on symbolic addition and context dependence.
+We find that while many model instance are highly additive, some can be highly non-additive (Fig.
+6a).
+However, when averaging across the model behavior 100 randomly sampled representations, this average behavior is perfectly described by a conjunction-wise additive function.
+This empirically confirms our proposition.
+Our proposition implies that these non-compositionally structured representations can still only systematically generalize on conjunction-wise additive tasks.
+To confirm this insight, we plotted the distribution of accuracies on the transitive equivalence problem across all random model instances.
+We found that while the models consistently learned the training set, they were indeed unable to generalize to the test set (Fig.
+6b).
+Finally, we investigated whether our analysis in Section 5 predicted the behavior of randomly sampled representations.
+Importantly, we have no theoretical guarantees for this scenario.
+Interestingly, however, the model behaviors were still well predicted by our theory.
+Specifically, we estimated the slopes that best described the compositional generalization behavior on symbolic addition for different training sets (as in Fig.
+5).
+We found that across different model instances, these slopes were highly varied (Fig.
+6c).
+However, the average slope across all random samples was well predicted by our analytical theory.
+Similarly, our theoretical insights on context dependence provided meaningful insight into the generalization behavior of these randomly sampled models.
+Specifically, on CD-3, the model more saliently representing the single components failed to generalize, whereas the model more saliently representing conjunctions of two components generalized above chance.
+In contrast, on CD-2, both models generalized above chance.
+This indicates that our insights on how prevalent shortcut biases are for different training datasets extends to randomly sampled representations.
+Appendix B Representational salience: a metric for compositionally structured representations
 B.1 Definition
@@
 Further, for compositionally structured representation, the similarity only depends on the total number of overlaps $|J|$.
-We therefore write $S(c;C)=S(J)$, where $c:=|J|$.
+We therefore write $S(k;C)=S(J)$, where $k:=|J|$.
+To illustrate how salience would be computed in practice, we consider an example representation $X\in\mathbb{R}^{n\times d}$.
+This representation gives rise to a kernel $K=XX^{T}\in\mathbb{R}^{n\times n}$.
+We can also understand the trial-by-trial similarity in terms of the components it is overlapping in — denoting the set of all subsets of $\{1,\dotsc,C\}$ by $\mathcal{J}$, we denote this by $O\in\mathcal{J}^{n\times n}$.
+In particular, $O_{ii}=\{1,\dotsc,C\}$.
+For example, if data points $1$ and $2$ overlap in the third component, $O_{12}=\{3\}$.
+$\text{Sim}(J)$ is then defined as the entries $K_{ij}$ where $O_{ij}=J$.
+Note that in a compositionally structured representation, whenever $O_{ij}=J$, $K_{ij}$ takes on the same value, but more generally, we can define $\text{Sim}(J)$ by taking the average similarity.
+We can then define the unnormalized salience $\overline{S}$ by recursively computing it from $\text{Sim}(J)$ using Eq.
+13.
+We then normalize the salience to get our final estimate.
 B.2 Why is representational salience a useful metric?
@@
 b, Salience for inputs with four components.
-We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022).
+We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018; 2019; Han et al., 2022).
 In Fig.
-5b, we plot the different saliences for an input with four components.
+7b, we plot the different saliences for an input with four components.
 Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig.
-2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions.
+3), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions.
 Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two.
@@
 In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig.
-5a).
+7a).
 As the random weight neural network becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation.
@@
 This highlights that for disentangled representations that do not represent all components with equal magnitude, compositional generalization behavior may vary strongly with different neural network architectures.
-B.5 Proof of Proposition 5.1
-See 5.1
+B.5 Salience in very deep ReLU networks
+Proposition B.2.
+For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $S(k;C)\to 0$ for $k<C$ and $S(C;C)\to 1$.
 Proof.
@@
 $\sigma(u):=A\min(u,0)+\max(u,0),\quad A\in[0,1).$ (14)
-By prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022),
+By prior work (Cho & Saul, 2009; Tsuchida et al., 2018; 2019; Han et al., 2022),
 $\mathbb{E}_{w}\left[\phi({w}^{T}{h}^{(l)}(z))\phi({w}^{T}{h}^{(l)}(z^{\prime}))\right]=\sigma^{2}\|h^{(l)}(z)\|_{2}\|h^{(l)}(z^{\prime})\|_{2}k\left(\hat{h}^{(l)}(z)^{T}\hat{h}^{(l)}(z^{\prime})\right),$ (15)
@@
 where we infer the latter inequality by visual inspection of the plot of this function. ∎
-Appendix C Detailed methods
-The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization.
-C.1 Models
+Appendix C DSprites representations
+To investigate whether our theory can describe compositional generalization in practically used disentangled models, we considered six different model architectures considered in Locatello et al.
+(2019) and trained on the DSprites dataset, an important benchmark for disentangled representation learning.
+This paper considers fifty random seeds for each model and further considers six different possible hyperparameters per architectures, resulting in a total of 1,800 models.
+In our analysis, we only analyze differences between architectures, pooling across all hyperparameter choices.
+C.1 Task setup
+DSprites consists of small black-and-white shapes and has five different underlying components: shape (three categories: hearts, ovals, and squares), size, x-position, y-position, and rotation.
+We consistently consider the largest possible size and hold the rotation fixed at zero.
+While x- and y-position can each take on 32 possible values, we only consider four different categories for each.
+Importantly, this is still a highly non-compositionally structured representation: not only do the disentangled representation learning methods often fail to discover the underlying factors of variation; the model also likely represents x- and y-positions that are closer to each other, as more similar, violating our compositionally structured assumption.
+These representations therefore present a particularly challenging test of our framework.
+For symbolic addition and transitive equivalence, we consider x- and y-position as the two components.
+For each task instance, we randomly determine which position takes on which component’s role.
+In total, we consider 50 randomly sampled task instance for each of the 1,800 models.
+For context dependence, we subsample two shapes and assume that these shapes provide the context cue.
+We then consider x- and y-position as feature 1 and feature 2.
+While the context dependence considered in the main text had six possible feature values for each feature, we now only consider four.
+As a result, we only consider two different datasets: CD-2, which leaves out all trials where feature 1 indicates category 2 and feature 2 indicates category 1; and CD-1 which only leaves out one conjunction of features.
+Put differently, on a given trial, the task could for example look as follows: if we see a heart shape, the model should output whether this object is in a certain x-position; if we see a square shape, the model should output whether this object is in a certain y-position.
+While the model has seen all x- and y-positions individually, it has not seen each combination of x- and y-positions.
+Finally, we consider either a direct linear readout from the disentangled representation or an initial transformation by a one-hidden-layer ReLU neural network with random weights.
+Figure 8: Conjunction-wise additivity in empirical disentangled models.
+a, Additivity across the different model architectures and four different task settings.
+All of the model behaviors are generally highly additive.
+b, Distribution of accuracies on transitive equivalence when averaged across different component roles.
+c, Variance ratio, capturing the ratio between the deviation between the representational similarity matrix and its compositionally structured approximation and the general variance of the representational similarity matrix.
+The numbers indicate that all disentangled models discover some compositional structure in their representation, but still deviate substantially from a compositionally structured representation.
+d, Generalization error on symbolic addition plotted against the estimated $S(1;2)$.
+e, Average accuracy on two versions of context dependence with a total of four features: CD-2 which now leaves out an entire orthant, and CD-1, which only leaves out a single data point.
+Again, models generally perform below chance on CD-2 but above chance on CD-1, though performance is generally really low.
+C.2 The models are well described by a conjunction-wise additive computation
+We first investigated whether the model predictions on the test set were well characterized by a conjunction-wise additive computation (see Section D.3), averaging this model behavior across the fifty random task instances.
+We found that they were generally well captured, though they were certainly not perfectly conjunction-wise additive (Fig.
+8a).
+Further, none of these models were able to systematically generalize on transitive equivalence (Fig.
+8b).
+This indicates that conjunction-wise additivity may characterize the generalization class of these highly non-compositionally structured representations as well — at least when averaged across task instances.
+C.3 The representations are partially compositionally structured
+We next investigated whether these representations exhibit compositional structure according to our definition.
+To determine this, we computed their representational similarity matrix and computed the average similarity for each set of overlaps.
+This instantiates the compositionally structured representational similarity matrix that most closely described the empirical representational similarity.
+We then determined the average squared deviation from this matrix, $\sigma_{cs}^{2}$, comparing it to the overall variance of this matrix, $\sigma_{var}^{2}$.
+Overall, the variance ratio, $\sigma_{cs}^{2}/\sigma_{var}^{2}$, characterizes the degree to which the given representation is compositionally structured.
+In a fully non-compositionally structured representation, the variance ratio will be roughly one, whereas in a fully compositionally structured representation, it will be zero.
+For the disentangled representation learning models, we found that this variance ratio took on value between 0.5 and 0.9 (Fig.
+8c).
+Thus, the models’ representation partially captured compositional structure, but also contained a lot of non-compositional structure.
+C.4 The non-compositional representational structure substantially impacts model behavior
+Finally, we tested whether our representational geometry analysis in Section 5 extended to the DSprites representations.
+We first estimated their salience $S(1;2)$ by computing the average representational similarity for each overlap and then computing the salience using Eq.
+13.
+Proposition 5.1 would predict that the generalization error should decrease with increasing $S(1;2)$.
+However, we do not observe such a trend (Fig.
+8d).
+This indicates that our representational geometry analysis does not apply to the DSprites representations, due to the way in which they violate the compositional structure assumption.
+Finally, we determined generalization accuracy on CD-2 and CD-1.
+We found that most models performed quite poorly, perhaps owing to the fact that shape is barely represented in these models.
+Nevertheless, we found that most models performed systematically below chance for CD-2 but systematically above chance for CD-1.
+Our analysis of context dependence in compositionally structured representations suggests that this may be because these models exploit a context-driven shortcut on CD-2 but not CD-1.
+Overall, our analysis therefore paints a nuanced picture of the applicability of our theory to practical disentangled representations.
+Our findings suggest that these models may still be well approximated by conjunction-wise additive computations and that this generalization class may therefore shed light on fundamental limits to the generalization of linear readout models.
+At the same time, to understand how these models generalize on conjunction-wise additive tasks, taking into account the specific ways in which they violate the compositional structure assumption is likely important.
+Appendix D Detailed methods
+D.1 Models
 Kernel model.
-We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn Pedregosa et al.
-(2011).
+We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn (Pedregosa et al., 2011).
 Note that this is equivalent to using a feature basis with the same resulting similarities.
@@
 When reporting results on lazy network, we assume $\sigma=1$.
+Compositional MNIST/CIFAR-10.
+We create a compositional version of MNIST and CIFAR-10 by concatenating multiple images along different channels.
+For example, the input to the MNIST version of the symbolic addition task had two channels, each containing one digit whereas the CIFAR-10 version of the symbolic addition task had six channels, three for each digit.
+We further varied the distance between the concatenated images either presenting them all on top of each other or presenting them offset by a certain number of pixels.
+Each image category corresponded to a certain component, where its role as ranodmly sampled for each task instance.
+The output was still given by a single scalar: the total magnitude for symbolic addition and the two categories for context dependence and transitive equivalence.
 Convolutional neural networks.
@@
 All weights are initialized with He initialization (He et al., 2015).
+We trained these networks with SGD using a learning rate of $10^{-4}$ and momentum of $0.9$.
 Residual neural networks.
@@
 We did not use data augmentation for MNIST. For CIFAR-10, we used a random flip and a random crop.
-C.2 Additivity analysis
+D.2 Reproducibility
+The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization.
+D.3 Additivity analysis
 To analyze how well a conjunction-wise additive computation can describe network behavior, we considered as the set of possible features a concatenation of one-hot vectors coding for each possible conjunction.
@@
 To test this prediction in vision network, we compared the additivity of the test set alone to the additivity of the entire dataset (Fig.
-6).
+9).
 First, we found that the predictivity on the test set was indeed generally very high.
@@
 Figure 9: $R^{2}$ of the conjunction-wise additive predictor on the test set as well as on both the training and test set.
-Appendix D Compositional task space
-D.1 Symbolic addition
-D.1.1 Proof of Proposition 5.2
-See 5.2
+Appendix E Compositional task space
+E.1 Symbolic addition
+E.1.1 Proof of Proposition 5.1
+See 5.1
 Proof.
@@
 Finally, $\overline{w}=\overline{v}=0$ immediately implies that $b=c=0$ and therefore $d=0$. ∎
-D.1.2 Behavior of deep ReLU networks
+E.1.2 Behavior of deep ReLU networks
 Next, we examined the behavior of deep ReLU networks with varying depth (Fig.
-7).
+10).
 On the extrapolation task ($\mathcal{W}=\{0\}$), these networks also had compressed model predictions on the compositional split that were approximately linearly distorted (though for deeper networks, this relationship seemed to have a slight S-shape).
@@
 Figure 10: Model prediction plotted against ground truth for different training sets and depths.
-D.1.3 Behavior of vision models
+E.1.3 Behavior of vision models
 We trained the Convnets on MNIST using seven different training sets $\mathcal{W}$: $\{[0]\}$, $\{[-4],[4]\}$, $\{[-2],[2]\}$, $\{[-4],[0],[4]\}$, $\{[-1],[0],[1]\}$, $\{[-4],[-3],[3],[4]\}$, $\{[-2],[-1],[1],[2]\}$.
@@
 Additionally, we also plot the mean squared error of the networks on the different training sets (Fig.
-9).
+12).
 This provides a more conventional (but harder to interpret) measure of performance compared to the slope depicted in Fig.
-4.
+5.
 We again see that the mean squared error on the compositional generalization set is decreasing with increasing distance (Fig.
-9a).
+12a).
 Further, it is decreasing with increasing training set size (Fig.
-9), though we note that the differences in the scales of the generalization set for different sets $\mathcal{W}$ renders it harder to directly to compare these values.
+12), though we note that the differences in the scales of the generalization set for different sets $\mathcal{W}$ renders it harder to directly to compare these values.
 In contrast, the slopes described in the main text are more easily comparable.
 Figure 12: a, Mean squared error as a function of distance. b, Mean squared error as a function of training set size $p$ (different dots corresponding to interpolation versus extrapolation.
-D.2 Context dependence
-D.2.1 General task definition
+Overall, our analysis indicates that our theory provides several valuable qualitative insights into the behavior of deep neural network models.
+Importantly, however, we are not able to provide quantitative bounds at the moment.
+To demonstrate this, we computed the salience $S(1;2)$ of the neural tangent kernel of the convolutional neural networks before and after being trained on the MNIST version of symbolic addition with $\mathcal{W}=\{0\}$ (Fig.
+13a).
+Importantly, $S(1;2)$ changed substantially during training, indicating that these networks are not trained in the kernel regime.
+As such, our theory does not apply exactly.
+To see whether it provided appropriate quantitative bounds, we then estimated the slope of each model instance on the compositional generalization dataset and plotted it against $S(1;2)$ of the neural tangent kernel at the beginning of training.
+We found that higher $S(1;2)$ generally resulted in a larger slope, as predicted by our theory.
+However, the relationship between $S(1;2)$ and slope did not adhere exactly to the quantitative predictions made by our theory.
+This indicates that while our theory can shed light on certain qualitative behaviors, it still leaves certain generalization behaviors in deep neural networks unclear.
+Figure 13: a, $S(1;2)$ in the convolutional neural network’s neural tangent kernel, before and after being trained on symbolic addition with $\mathcal{W}=\{0\}$.
+b, The estimated slope for each model instance plotted against $S(1;2)$.
+While increased $S(1;2)$ generally yields a larger slope (as predicted by our theory), the relationship between $S(1;2)$ and the slope does not follow the exact quantitative relationship predicted by our theory.
+Finally, our analysis in Section 5.1 suggests that deeper models should yield more conjunctive representations, and therefore exhibit worse generalization on symbolic addition.
+To test this prediction, we varied the number of fully connected layers in the convolutional neural network and trained these networks on the symbolic addition task with the training set $\mathcal{W}=\{0\}$.
+We found that deeper networks indeed generalized worse (Fig.
+14).
+Figure 14: Mean squared error on the compositional generalization set for different depths of the fully connected network at the end of the convolutional neural network architecture.
+E.2 Context dependence
+E.2.1 General task definition
 We consider inputs with three components, $(z_{\text{co}},z_{f1},z_{f2})$.
@@
 Note that in the main text, we consider $C_{1}=\{1\}$, $C_{2}=\{2\}$, and six possible values for $z_{f1},z_{f2}$, where the decision function maps three onto $1$ and three onto $-1$.
-D.2.2 Novel stimulus compositions are conjunction-wise additive
+E.2.2 Novel stimulus compositions are conjunction-wise additive
 If the test set consists in novel combinations of stimuli, this is a conjunction-wise additive computation.
@@
 $f(z_{\text{co}},z_{f1},z_{f2})$ $=f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$ (46)
-D.2.3 Novel rule compositions are not conjunction-wise additive
+E.2.3 Novel rule compositions are not conjunction-wise additive
 We could also imagine an alternative generalization rule in a task where there are multiple components indicating the same context: $C_{1}=\{\text{co}_{1},\text{co}_{2}\}$ and $C_{2}=\{\text{co}_{3},\text{co}_{4}\}$.
@@
 However, the conjunction-wise additive mapping depends on having observed each context in conjunction with each feature and this task is therefore non-additive.
-D.2.4 Coefficient groups
+E.2.4 Coefficient groups
 In Fig. 4d, we grouped the inferred coefficients into categories. We here explain these categories:
@@
 We then compute the average absolute magnitude within each of these groups in order to determine their overall relevance to model behavior.
-D.2.5 Feature-learning ReLU networks
-We find that feature-learning ReLU networks generalize consistently on CD-1 anREGULARIZATIONd CD-2 but not CD-3.
-They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig.
-10).
+E.2.5 Feature-learning ReLU networks
+We find that feature-learning ReLU networks generalize consistently on CD-1 and CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 15).
 Figure 15: Inferred values for different conjunctive groups on context dependence.
-D.2.6 Vision networks
+E.2.6 Vision networks
 We additionally analyzed ConvNets for different distances between the digits.
 We found that convolutional neural networks trained on MNIST successfully generalized on CD-1 and CD-2, but not CD-3 (Fig.
-11).
+16).
 For smaller distances between digits, the models tended to generalize worse on CD-1 and CD-2 and gradually reverted to chance accuracy (i.e.
@@
 Further, the networks were generally highly additive ($R^{2}>0.95$), but became worse for lower distance (Fig.
-11b).
+16b).
 Finally, across all distances, they had a high magnitude associated with the context cue, though this magnitude decreased for small distances — consistent with the accuracy of the network increasing from below chance to chance level (Fig.
-11c).
+16c).
 Figure 16: Performance of convolutional neural networks trained on an MNIST version of context dependence.
 For different distances and training sets, we plot a, the accuracy on different splits, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
-D.3 Invariance and partial exposure
+E.3 Invariance and partial exposure
 We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut bias.
 In both cases, the input consists of two components and the mapping only depends on the first (Fig.
-12a).
+17a).
 In the invariance case, we don’t see the second component vary at all, in the partial exposure case, we see one instance of the second component.
@@
 However, as $S(1;2)$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig.
-12b).
+17b).
 Shortcut distortion.
@@
 As a result, the test margin for the partial exposure task decreases even more strongly as a function of $S(1;2)$: $m=\tfrac{2S(1;2)^{2}}{1-2S(1;2)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig.
-12b).
+17b).
 Figure 17: a, Schema for invariance and partial exposure. b, Generalization margin for the two tasks.
@@
 $m=\frac{2S(1;2)^{2}}{1-2S(1;2)^{2}}.$ (55)
-D.4 Other mathematical operations
+E.4 Other mathematical operations
 We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$ together with some composition function $C(v_{1}(z_{1}),v_{2}(z_{2}))$.
@@
 multiplication, division, or exponentiation, the task will be non-additive.
-D.5 Logical operations
+E.5 Logical operations
 […]
-D.6 Transitive ordering
-Transitive ordering is a popular task in cognitive science (often called transitive inference, (McGonigle & Chalmers, 1977)).
+E.6 Transitive ordering
+Transitive ordering is a popular task in cognitive science (often called transitive inference, McGonigle & Chalmers (1977)).
 Here the subject is presented with two items $z_{1}$, $z_{2}$ drawn from an unobserved hierarchy $>$.
 It should then categorize whether $z_{1}>z_{2}$ or $z_{2}>z_{1}$.
-Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ (Lippl et al., 2024).
+Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ Lippl et al.
+(2024).
 It is therefore additive, in contrast to transitive equivalence.
@@
 In this case, the model would generalize to comparisons between these different hierarchies as well.
-Appendix E A feature-learning mechanism for non-additive compositional generalization
+Appendix F A feature-learning mechanism for non-additive compositional generalization
 Figure 18: Feature learning enables generalization on transitive equivalence.
@@
 Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs.
-7 and 10).
+10 and 15).
 On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig.
-13a).
+18a).
 To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm (Lyu & Li (2020); Chizat & Bach (2020); cf.
@@
 (2023); Fig.
-13b).
+18b).
 Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g.
@@
 Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig.
-13c).
+18c).
 Thus, rich networks’ capacity for abstraction gives rise to an additional compositional motif, allowing them to generalize on transitive equivalence.
@@
 However, with increasing distance, the network started to improve its compositional generalization (Fig.
-13d), demonstrating that a convolutional network can benefit from this rich compositional motif.
+18d), demonstrating that a convolutional network can benefit from this rich compositional motif.
+Figure 19: This figure composes the relevant theoretical insights and empirical experiments to illustrate how they speak to each other.